### v1

# VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap

Abstract
Recent interest in Large Vision-Language Models (LVLMs) for practical applications is moderated by the significant challenge of hallucination or the inconsistency between the factual information and the generated text. In this paper, we first perform an in-depth analysis of hallucinations and discover several novel insights about how and when LVLMs hallucinate. From our analysis, we show that: (1) The community’s efforts have been primarily targeted towards reducing hallucinations related to visual recognition (VR) prompts (e.g., prompts that only require describing the image), thereby ignoring hallucinations for cognitive prompts (e.g., prompts that require additional skills like reasoning on contents of the image). (2) LVLMs lack visual perception, i.e., they can see but not necessarily understand or perceive the input image. We analyze responses to cognitive prompts and show that LVLMs hallucinate due to a perception gap: although LVLMs accurately recognize visual elements in the input image and possess sufficient cognitive skills, they struggle to respond accurately and hallucinate. To overcome this shortcoming, we propose Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method for alleviating hallucinations. Specifically, we first describe the image and add it as a prefix to the instruction. Next, during auto-regressive decoding, we sample from the plausible candidates according to their KL-Divergence (KLD) to the description, where lower KLD is given higher preference. Experimental results on several benchmarks and LVLMs show that VDGD improves significantly over other baselines in reducing hallucinations. We also propose VaLLu, a benchmark for the comprehensive evaluation of the cognitive capabilities of LVLMs.
1 Introduction
Large Language Models (LLMs) pre-trained at on web-scale text data with the next token prediction objective implicitly compress world knowledge in their parameters zhao2023survey . These models learn general-purpose representations, which can then be aligned with the desired response characteristics zhang2023instruction . This step generally involves fine-tuning on instruction-response pairs, also known as instruction tuning (IT) wei2022finetuned , and is followed by an optional step of reinforcement learning from human feedback (RLHF) bai2022training . Such aligned LLMs can be engaged through text-based inputs, or prompts, directing them to perform various tasks, including those based on information retrieval and reasoning.
Recent advancements in the research community have demonstrated success in multi-modal alignment, whereby fine-tuning LLMs with multi-modal instruction-response pairs enables them to execute tasks that conform to the text prompt as well as additional multi-modal inputs yin2023survey . However, similar to text-only LLMs, Multi-modal Large Language Models (MLLMs) are also prone to hallucination liu2024survey , or the misalignment between factual visual content and corresponding textual responses generated. This is a prominent concern that impacts the reliability and applicability of MLLMs across domains. In this paper, we investigate hallucinations in Large Vision Language Models (LVLMs) to explore the underlying causes and mechanisms.
The research community has made significant attempts to reduce LVLM hallucinations liu2024survey . These attempts include better architectures liu2023improved ; wang2023cogvlm , better alignment techniques chen2024far , and explicit hallucination mitigation techniques yin2023woodpecker ; zhang2024debiasing ; wang2024mitigating . Each approach has demonstrated superior performance over others on benchmark datasets. In this paper, we conduct a comprehensive analysis of LVLM hallucinations to explore the mechanisms and conditions under which LVLMs hallucinate and to measure the community’s progress in addressing these issues. Our results reveal the following:
  1. 1.
    Prior work has majorly focused on alleviating hallucinations for visual recognition (VR) prompts (e.g., object hallucinations in prompts that demand image descriptions). While this has led to significant improvements in VR benchmarks (like AMBER), improvements in cognitive prompts that require additional skills, like reasoning and information seeking, have remained stagnant. Additionally, we perform an in-depth analysis to show that there has been relatively less progress in alleviating hallucinations for certain aspects (like descriptions of non-real world scenes) and categories (described in Section 4.2) of VR hallucinations.
  2. 2.
    For cognitive prompts we perform controlled experiments to show that LVLMs suffer from a visual perception gap: although they can accurately recognize visual elements in the input image and possess sufficient knowledge and reasoning skills to provide a factual response, they fail to fully perceive the image to generate a factual response.
  3. 3.
    Inspired by our findings, we propose a Visual Description Grounding Decoding (VDGD), a simple and novel training-free technique to alleviate LVLM hallucinations. Specifically, we first generate a description of the input image. Next, at every decoding step during auto-regressive response generation, we calculate the Kullback–Leibler divergence (KLD) between the top-k tokens in the current logit and all the logits of the image description. Finally, to predict the next word, we sample from the plausible candidates according to their KLD to the description, where lower KLD is given higher preference. VDGD improves performance on standard benchmarks over other mitigation techniques from literature.
2 Experimental Setup
LVLMs. For the scope of this work, we experiment with 6 different types of LVLMs, namely LLaVA-v1 liu2023visual , LLaVA-v1.5 liu2023improved , LLaVA-v1.6 liu2024llavanext , mPLUG-Owl2 ye2023mplugowl2 , InternLM-X zhang2023internlm and CogVLM wang2023cogvlm , each utilizing a 7B parameter language model, though the number of vision parameters varies. To accurately evaluate their capabilities in visual recognition and perception, we ensure that all models use the same foundational language model. More details can be found in Appendix F.
Evaluation Setup. For evaluating LVLMs, we perform a mix of expert human evaluation and automatic evaluation using GPT-4 Turbo (gpt-4-turbo-2024-04-09). Recent research indicates that using ChatGPT and GPT-4 to score and assess outputs from LLMs aligns well with human evaluations, offering the added benefit of cost reduction zhou2023lima ; ghosh2024closer ; lin2023unlocking ; chan2024chateval ; xu-etal-2023-instructscore ; zheng2023judging . We devise an explainable and multi-aspect evaluation prompt lin2023unlocking ; ghosh2024closer , which prompts the judge (GPT-4) to assign a score between 1 and 5 to an LVLM’s response to an instruction on five aspects: helpfulness, clarity, factuality, depth, and engagement. It also provides an explanation for each score and extracts hallucinated phrases. Prompt is in Appendix D. With hallucinations as our primary focus, all results shown in the main paper are for the factuality metric. The metric is designed to penalize hallucinations and evaluate correctness in open-ended generations. Scores for all metrics and all graphs are in Appendix I.
Evaluation Benchmarks. We employ a range of benchmarks for LVLM evaluation that assess the diverse capabilities of a model. These include AMBER wang2023llm (visual recognition), SynthDoG kim2022donut (OCR), Oven hu2023open (information seeking), MMMU yue2023mmmu (general and holistic reasoning), MMC liu2023mmc , MathVista lu2023mathvista and MATH-Vision wang2024measuring (mathematical reasoning), HallusionBench guan2023hallusionbench , and MME fu2023mme (chart understanding). We only use the sub-sets of these datasets that demand open-ended generation for response, as LLMs have been shown not to be robust to MCQ zheng2024large . Additionally, since automatic evaluation on all these datasets can be expensive, we propose VaLLu, detailed in Section 6.2.
3 Current Progress on Solving LVLM Hallucinations
As LVLMs advance and become more accessible, their application extends beyond merely generating image descriptions. Thus, to better understand hallucinations, we break down information processing by an LVLM, when prompted with an input image and a text instruction, into its constituent steps:
{mdframed}
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=-3pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt]
[…]
As depicted in Fig. 1, diverse prompts applied to a single image may evaluate the LVLM for various skills. Visual Recognition (VR) involves recognizing visual elements within the image (like objects, backgrounds, etc.), their interrelations, and their unique characteristics.
Figure 1: Depending on the text instruction, an LVLM might be assessed on one or more different capabilities.
A prompt that just seeks a description of the image or details about a specific element explicitly evaluates only visual recognition capabilities. Visual Perception (VP), which follows and moves beyond VR, entails interpreting and contextualizing the observed elements in the context of the entire image fu2024blink . For tasks that extend beyond basic image description, accurate visual perception is crucial, surpassing the mere recognition of individual elements. If the prompt seeks further knowledge-specific information related to the elements in the image, then the next step involves extracting relevant information (IE) from the knowledge learned during pre-training. Finally, some prompts may necessitate the LVLM to reason over the visual data, textual prompt, and any implicitly extracted information to deduce a response. While VR relies on the performance of the vision encoder, the abilities to extract accurate information and reason are based on the strength of the foundational language model used in alignment tuning (as alignment does not teach a model new knowledge or skills ghosh2024closer ; zhou2023lima ). VP, which is done implicitly by the model, is a skill taught during alignment tuning and serves as a bridge between VR and other cognitive abilities. barrow1978recovering ; torralba2002depth ; hartley2003multiple
A closer look at the progress in mitigating hallucinations. Fig. 2 (left) compares
Figure 2: (Left) Performance comparison of different LVLMs on various benchmarks. (Right) Performance comparison of different hallucination mitigation techniques applied to LLaVA-1.5.
the performance of LVLMs across 5 benchmarks. As we clearly see, while newer models that were trained using improved architectures and alignment training data boosted performance on the AMBER benchmark (for recognition), performance on other reasoning and information-seeking benchmarks remained stagnant.
Fig. 2 (right) compares the performance of LLaVA-1.5 across the same 5 benchmarks when employed with explicit hallucination mitigation techniques, namely, Visual Contrastive Decoding (VCD) damonlpsg2023vcd , OPERA huang2023opera , Woodpecker yin2023woodpecker and LURE zhou2023analyzing . As we clearly see, while all these methods boost performance on AMBER, similar to the previous experiments, performance on other reasoning and information-seeking benchmarks remains stagnant.
{mdframed}
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Current advancements in reducing hallucinations via improved model architectures, training data, and explicit mitigation strategies are limited to improving the visual recognition capabilities of LVLMs. Progress in addressing hallucinations beyond those related to visual recognition has been relatively quite less.
4 A Closer Look at Visual Recognition Hallucinations
As we have seen in the earlier section, advancements in LVLMs and hallucination mitigation techniques have led to considerable improvements in visual recognition capabilities by these models. In the next few subsections, we take a closer look at visual recognition hallucinations to show that the models and methods discussed above only alleviate hallucinations in some specific cases.
4.1 Moving Beyond Real-World Scenes
As LVLMs are now being employed on a wide range of tasks, considerable research efforts are being made to evaluate the efficacy of LVLMs to reason on images beyond real world
Figure 3: (Left) Performance comparison of different LVLMs on various benchmarks when prompted to only describe the image. (Right) Performance comparison of different hallucination mitigation techniques applied to LLaVA-1.5.
scenes, i.e., charts liu2023mmc or math problems lu2024mathvista . However, how often do LVLMs hallucinate while describing images beyond real-world scenes? Fig 3 (left) compares the performance of various LVLMs across 5 benchmarks. While MMMU and MathVista include images with mathematical figures, MMC includes images with charts from diverse sources, and SynthDoG prompts the model for OCR. For AMBER, MMMU, MathVista, and MMC, we prompt the LVLM with the prompt “Describe the image.” and for SynthDoG we prompt the LVLM with the prompt Extract text from scanned PDFs, photos, and even handwriting.. As we clearly see, compared to AMBER, models hallucinate more often when describing visuals beyond real-world scenes. Fig 3 (right) further shows that hallucination mitigation techniques improve performance on AMBER but rarely other benchmarks.
4.2 Causal Categorization of Visual Recognition Hallucinations
Overview. In our detailed analysis of model responses on the AMBER dataset, we identified intriguing patterns in which hallucinations occur. Although the community has primarily focused on the broader topic of object hallucinations liu2024survey , our findings reveal that visual recognition hallucinations can be classified into several subcategories based on their specific occurrence patterns.
Algorithm 1 Categorizing Visual Hallucinations
Given: Model Response $R$, IT dataset $D_{\text{IT}}$, Aligned LVLM $M_{\text{aligned}}$, base LLM $M_{\text{base}}$,
Obtain Top-k similar instances $S_{R}$ to $R$ from $D_{\text{IT}}$ using CLIP
Extract visual elements $V$ from $R$.
Extract hallucinated phrases $P$ from $R$ using LLM judge divided into Object, Relation, and Action hallucinations
procedure TokenAnalysis($R,M_{\text{aligned}},M_{\text{base}}$)
   for each word $p$ in $P$ do
     if $p$ in $V$ then $\triangleright$ Consider only 1st token of $p$
       Calculate Base-Rank BR of $p$triangleright Defined in Section 4.2
       if $\text{{BR}}>0$ then $\triangleright$ marginal or shifted token
         if $p$ in $S_{R}$ then
            $p$ is an IT Hallucination
         else
            if $p$ == Relation & $p$ != Object then
              $p$ is a Style Hallucination
            else
              $p$ is a Vision Hallucination
            end if
         end if
       $\triangleright$ unshifted token
         $p$ is a Language Hallucination
       end if
     end if
   end for
end procedure
[…]
(1) Language Hallucinations: These are hallucinations that are caused by language priors and are unshifted tokens (tokens that would have been output by the non-aligned LLM knowledge through prior context without an image). During auto-regressive decoding, the model might over-rely on the prior tokens in the response rather than the input image. Thus, this may leads the model to mention details in the response from its pre-trained knowledge that is not actually faithful to the image.
Figure 4: Types of Visual Recognition Hallucinations. We define Algo. 1 to divide VR hallucinations into 4 different categories automatically: Language, Vision, Style, and IT (explained further in Sec. 4.2). While language and vision hallucinations have been explored earlier, and methods to alleviate them have been proposed, we show for the first time that Style and IT hallucinations exist and existing methods fail to alleviate them.
(2) Vision Hallucinations: Unlike language hallucinations, the tokens for vision hallucinations are generated by attending to the input image, and are shifted tokens. These are hallucinations caused when the model fails to accurately recognize specific components within an input image, or just inaccurate reasoning or IE, among other factors.
(3) Style Hallucinations. These are hallucinations caused by style imitation. Open-source VLLMs are often trained on instruction-tuning data synthesized from proprietary closed-source models. Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset ghosh2024closer ; gudibande2024the (e.g., lengthy answers, a summary line at the end of the response, etc.). However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding ghosh2024closer .
Figure 5: (Left) Frequency comparison of hallucination categories. (Right) Comparison for LLaVA-v1.5 with hallucination mitigation techniques. The top graph compares AMBER, and the bottom graph compares MMMU.
(4) Instruction Tuning (IT) Hallucinations. These are hallucinations caused by biases learned during the instruction tuning stage. LLMs are known to mimic the responses associated with the unfamiliar examples in the model’s instruction-tuning data (i.e., those unfamiliar to the pre-trained base model) ghosh2024closer . Formally, this can be defined as distribution shift. As visual instruction tuning is an extreme case of such a distribution shift, IT hallucinations are a major cause of hallucinations in LLMs.
Performance Comparison. Fig. 5 (top;left) compares the frequency of different categories of hallucinations, across various LVLMs for AMBER. Fig. 5 (top;right) displays the frequency of different hallucination categories for LLaVA-v1.5 alongside various hallucination mitigation techniques. Fig. 5 (bottom) offers a similar comparison for MMMU. Our key findings include: (1) Model-based improvements and mitigation strategies have significantly reduced language and vision hallucinations, but progress in style and IT hallucinations remains limited. (2) Specific mitigation techniques are more effective for certain types of hallucinations. For instance, decoding-based methods like VCD and OPERA excel at reducing language hallucinations, whereas object-grounding approaches like Woodpecker are more effective for vision hallucinations. Decoding-based techniques benefit from low-confidence hallucinated tokens and uncertainty encoded in logits as shown in Fig. 4. Further details can be found in Appendix H.1. (3) Existing methods fail to reduce hallucinations on non-VI benchmarks like MMMU. We attribute this to the algorithmic biases of these methods to alleviate VR hallucinations (for example, these algorithms are crafted for real-world scenes).
{mdframed}
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: While community efforts have enhanced the visual recognition capabilities of LVLMs, substantial challenges persist in mitigating hallucinations for images that depict scenarios beyond real-world scenes. Additionally, hallucinations can be distinctly categorized based on their occurrence patterns, and some categories have received lesser attention.
5 The Visual Perception Gap
As illustrated in the previous section, current progress in LVLMs primarily alleviates VR hallucinations. In this section, we delve into LVLM hallucinations in various benchmarks, suggesting that the issue may stem from alignment problems. We first show that for cognitive prompts, LVLMs do not tend to pay attention to the input image but respond only using language priors. Next, we show that this problem may arise from a visual perception gap, i.e., although LVLMs may accurately recognize visual elements in the input image and possess sufficient knowledge and reasoning skills to provide a factual response, they still fail to respond accurately to the prompt and hallucinate.
5.1 Assessing the Extent to Which LVLMs Attend to the Input Image
Figure 6: Base Rank Comparison between AMBER and MATH-Vision datasets as a function of token position in responses generated by CogVLM.
To assess the influence of the image on the LVLM response, we compare the Base Rank (defined in Section 4.2) of tokens across datasets. A lower value would signify that the model solely responds using language priors while a higher value would signify that the model pays attention to the image, i.e., it responds with tokens that are different from what the model would have responded with only language priors. Fig. 6 plots the Base Rank of tokens against the token position in response (average across the dataset). As we see, the Base Rank for AMBER, which prompts for an image description, is higher than MMU, which requires it to reason or extract knowledge based on the image. This hints towards an alignment issue: different from AMBER, for MMMU LVLMs are unable to perceive the image before responding, under-relying on visual cues and over-relying on language priors for generating responses. In Appendix H.1 we explain why decoding-based mitigation techniques discussed in Section 4.2 do not work for cognitive prompts.
5.2 Disentangling Visual Recognition from Other Cognitive Skills
Figure 7: (Left) Performance comparison of different LVLMs when prompted w/ original prompt vs rephrased prompts w/o image (-t). (Right) Performance comparison of different LVLMs for their ability to generate a faithful image description.
A fundamental goal of alignment fine-tuning is to enable an LVLM to master implicit visual perception. As discussed in Section 3, when faced with an information-seeking or reasoning prompt, an LVLM is expected to respond directly to the prompt and to have implicitly comprehended the accompanying image with respect to the prompt instruction. To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we aim to isolate visual recognition from other cognitive skills and evaluate them independently. To achieve this, we ask GPT-4 to rewrite the instructions in prompts from various visual information-seeking and reasoning benchmarks, such that they can be answered without the accompanying image. We show examples in Appendix H.6. Subsequently, we prompt LVLMs with these modified prompts without the image. Next, we prompt the LVLM to separately generate a description of the input image and evaluate this description for its faithfulness and detailedness (i.e., how well the image description can aid in a factual response to the original prompt). Evaluation prompts are provided in Appendix D. Fig. 7 (left) compares the performance of text-only prompts (denoted with -t in the key) with their original image-text counterpart across various LVLMs and Fig. 7 (right) shows their VR scores. As we see, in all cases, LVLMs fare well in VI, and are able to provide better factual responses than with the original prompt together with the input image. Thus, we conclude that while LVLMs demonstrate proficiency in visual recognition and have sufficient knowledge and reasoning abilities to generate factual responses, they encounter difficulties in employing these skills when responding to prompts that require their simultaneous application. This underscores a significant gap in alignment via lack of visual perception: LVLMs can recognize visual elements in images but fail to fully perceive them to accurately link what they recognize to their internal knowledge for reasoning or information extraction. We hypothesize a possible source of this issue to be the fact that most fine-tuning (or alignment) datasets for LVLMs consist of image-description pairs, which makes them struggle with accurate visual perception for non-description prompts.
5.3 Uncertainty in the Logit Space
Models Dataset k Variance Range Avg.
LLaVA-v1 AMBER 4.19 0.05 0.31 0.23
LLaVA-v1.5 AMBER 3.91 0.07 0.39 0.29
CogVLM AMBER 3.39 0.09 0.46 0.21
[…]
LLaVA-v1.5 MMMU 3.53 0.09 0.41 0.23
CogVLM MMMU 3.16 0.19 0.47 0.25
Table 1: Post-truncation probability statistics for various models using the elbow method. All values are computed across all hallucinated tokens in each dataset.
To investigate the origins of hallucinations, we analyze the logit space of hallucinated tokens. Specifically, after applying the softmax operation to the logits, we rank them by their probability scores. We then use the elbow method (detailed in Appendix H.5) to truncate the distribution, obtaining the top-k tokens along with their corresponding probabilities. This truncation approach, based on the elbow method, effectively identifies the tokens most likely to be selected during multinomial sampling. Table 1 shows several statistics of the top-k probabilities for hallucinated tokens (we only consider the first token of each word). As evident from the value of k, standard deviation, and the average, the top-k probability space for hallucinated tokens is dominated by a set of low and equally confident tokens. Thus, each of these tokens has almost a similar chance of being sampled with multi-nomial sampling.
{mdframed}
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Our analysis of disentangling VR from other cognitive skills reveals that LVLMs generally demonstrate accurate VR and possess the necessary skills to respond accurately. However, they struggle to effectively link they recognize to their internal knowledge when responding directly, which inhibits the accurate execution of these skills. This issue underlines what we identify as the visual perception gap.
6 Visual Description Grounding Decoding (VDGD)
Figure 8: Illustration of VDGD Hallucination Mitigation.
To bridge the visual perception gap and reduce hallucinations, we propose Visual Description Grounding Decoding (VDGD), a simple, effective, and training-free hallucination mitigation technique. Specifically, during the decoding process, VDGD first adds the image description as a prefix to the text instruction of the prompt and then grounds token generation to the description of the input image by selecting a token that deviates the least from the description. We measure token deviation from the description with a novel formulation that employs Kullback–Leibler divergence at its core. The description is generated by the LVLM itself. VDGD operates on 2 core hypotheses: (1) Since LVLMs rely on language priors for cognitive prompts (Section 5.1), adding the image description as a prefix to the instruction can lead to improved visual perception by explicitly conditioning the response on the description. (2) Re-scoring the equally and low-confidence tokens (Section 1) based on deviation from the description can increase the confidence of the accurate token from several low-confidence tokens xu2023look .
6.1 Methodology
Fig. 8 illustrates the VDGD methodology. For an input prompt, we first employ the LVLM to generate a description of the image accompanying the prompt (see Appendix D for the prompt). Next, we concatenate the generated description as a prefix to the original prompt.
Formally put, let the VLM, denoted as $p_{VLM}$, receive an input prompt of length $n$ symbolized as $X_{\text{pre}}=x_{1}\dots x_{n}$, with each $x_{i}$ representing a token from the vocabulary $\mathcal{V}$. The decoder’s objective is to produce responses of length $m$, represented as $X_{\text{res}}=x_{n+1},\dots,x_{n+m}$. During the decoding phase, the model iteratively decodes one token at a time, conditioning on the tokens that have been generated so far:
$p_{VLM}(X_{\text{resp}}\mid X_{\text{pre}})=\prod_{i=n+1}^{n+m}p_{VLM}(x_{i}\mid x_{<i})$ (1)
Here, $p_{VLM}(x_{i}\mid x_{<i})$ denotes the probability distribution for the next token, $x_{i}$, given the previous tokens. Inspired by our findings in Section 5.3, we first apply a plausibility constraint to truncate the vocabulary space of the next token $x_{i}$ li-etal-2023-contrastive . Essentially, this step cuts off tokens from the vocabulary $\mathcal{V}$ below the elbow and keeps only the top-k plausible tokens. This is done as follows:
$p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\begin{cases}p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right),&\text{ if }p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\geq\alpha\max_{w}p_{\operatorname{VLM}}\left(w\mid x_{<i}\right)\leq\alpha\\ -\infty,&\text{ Otherwise }\end{cases}$ (2)
where $\alpha$ is a hyper-parameter between [0,1]. Now let $\mathcal{V}_{K}^{x_{i}}$ be the set of top-K most plausible tokens from the vocabulary $\mathcal{V}$ after the truncation of $x_{i}$. Next, for each token $w_{k}\in\mathcal{V}_{K}^{x_{i}}$ in the top-K plausible tokens, we calculate the deviation of the token with the $n$ tokens of the prompt as follows:
$\mathrm{KL}_{w_{k}}^{x_{i}}=\min_{1\leq j\leq n}\mathrm{KL}\left(\operatorname{one-hot}(w_{k})\|p_{\operatorname{VLM}}\left(\cdot\mid x_{<j}\right)\right).$ (3)
where $\mathrm{KL}$ is the Kullback–Leibler divergence and $\operatorname{one-hot}(w_{i})\in\mathrm{R}^{\mathcal{V}}$ is the one-hot representation of token $w_{k}$ where each value is 0 except the position corresponding to the token-id of $w$, which is 1. Next, we replace the value of $w_{k}$ in the logit with $-\mathrm{KL}_{w_{k}}^{x_{i}}$ and apply softmax on the logits to obtain a probability distribution. Thus, tokens with larger KLD (or a large deviation) to the input prompt are less likely to be sampled given the softmax operation upon KL divergence. With this distribution, we can now apply any sampling technique to sample the next token.
6.2 VaLLu Benchmark
Figure 9: Distribution of task types in VaLLu.
Automatic evaluation of LVLMs, with powerful closed-source models, on multiple large benchmarks can be prohibitively expensive. Thus, we propose VaLLu, a benchmark that we build by amalgamating existing benchmarks. VaLLu comprises a total of 1500 instances, with 554 sourced from Oven, 535 from MMMU, 218 from MMC, 74 from MathVista, 60 from HallusionBench, 53 from MATH-Vision, and 14 from MME fu2023mme . We only include instances with instructions that require open-ended generations and do not include any binary Yes/No or Multi-choice questions. We control the size to ensure the evaluation is affordable while keeping the diversity of tasks and topics for comprehensive analysis. VaLLu is manually filtered (we found various noisy examples in existing benchmarks that do not have a definite ground-truth response and thus make evaluation difficult, we discuss this further in Appendix H.3), annotated (we annotate each instance for the type of skill required by the LVLM to generate a response, we illustrate this in Fig. 9) and answered (we accompany each instance with its ground-truth response).
6.3 Results and Analysis
Experimental Setup and Baselines. We compare VDGD against VCD, OPERA, Woodpecker, LRV, and LURE mitigation techniques. Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7). We employ greedy decoding for all methods as we find no difference in performance on sampling. We generate image descriptions for VDGD with the VCD mitigation technique and ablate the same.
Quantitative results. Table 2 compares the performance of VDGD with other baselines on the VaLLu Benchmark. We report factuality scores averaged across 3 runs for 3 random seeds. As we clearly see, VDGD not only outperforms all baselines but is also the only method that alleviates and improves performance on VaLLu. Results on other full benchmarks, including MMU, MathVista, and MATH-Vision, are provided in Appendix I.5, together with scores for other aspects beyond factuality where we show that VDGD does not compromise the depth and expressivity of responses. Removing KL divergence or just adding the description prefix results in a drop in performance.
Qualitative examples.Fig 10 compares responses from vanilla greedy decoding with VDGD. Grounding response generation results in more factual reasoning and information extraction.
Figure 10: Illustration of responses generated with VDGD compared with greedy decoding. More in Sec. H.2.
Methodology LLaVA-v1 LLaVA-v1.5 LLaVA-v1.6 mPLUG-Owl2 InternLM-X CogVLM
Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82
Vanilla-sampling 1.86 2.01 2.64 2.18 2.70 2.83
VCD 1.47 1.55 1.80 1.80 2.32 2.63
OPERA 2.04 2.05 2.62 2.26 2.65 2.85
Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91
LRV 2.03 2.10 2.64 2.24 2.64 2.78
LURE 1.98 2.03 2.65 2.19 2.59 2.88
[…]
VDGD (-) VCD $2.08_{\pm 0.09}$ $2.43_{\pm 0.15}$ $3.01_{\pm 0.08}$ $2.54_{\pm 0.07}$ $3.26_{\pm 0.12}$ $2.95_{\pm 0.06}$
VDGD (-) KLD $2.05_{\pm 0.10}$ $2.30_{\pm 0.10}$ $2.78_{\pm 0.12}$ $2.37_{\pm 0.06}$ $3.01_{\pm 0.02}$ $2.87_{\pm 0.08}$
Table 2: Comparison of VDGD with other hallucination mitigation techniques. All results are on the proposed VaLLu benchmark for the "Factuality" metric. Results on other benchmarks can be seen in Appendix I.5.
7 Conclusion, Limitations and Future Work
In this paper, we conduct extensive experiments to analyze hallucinations in LVLMs. Our findings reveal that although there has been considerable progress in reducing hallucinations related to visual recognition, reducing hallucinations for cognitive prompts has been limited. Further investigation into LVLM responses to cognitive prompts reveals a visual perception gap: despite having the required capabilities, LVLMs often produce hallucinations when responding to prompts. To address this issue, we introduce VDGD, a novel and training-free method to reduce hallucinations. Our results demonstrate that VDGD is the first method to effectively reduce hallucinations in responses that necessitate additional cognitive skills beyond just visual recognition.
As part of future work, we would like to work on the current limitations of our work, which include: (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
Appendix A Appendix
In the supplemental material,
  1. 1.
    B: We provide a detailed explanation of computational complexity of VDGD algorithm described in section 6.
  2. 2.
    C: We describe the impact of the paper on society and research communities.
  3. 3.
    D: We provide the prompts used for object detection, description, rephrasing, noise removal and evaluations required for our experiments.
  4. 4.
    E: We provide a detailed description of all the datasets utilized in this paper.
  5. 5.
    F: We provide a detailed description of all the models we have experimented with.
  6. 6.
    G: We provide a detailed description of all the prior-art hallucination mitigation techniques used as baselines.
  7. 7.
    H: We provide details about:
    • H.1: We provide detailed explanation about Logit Space for VR and Cognitive Prompts.
    • H.2: We provide examples comparing generations from Greedy, Decoding, VCD and VDGD.
    • H.3: We provide examples of noisy samples in datasets.
    • H.4: We provide and in-depth explanation of all the details of Algorithm 1.
    • H.5: We explain the elbow method used for logit analysis.
    • H.6: We provide examples of rephrased prompts which do not require image input.
    • H.7: We provide details of computational resources used for this paper.
  8. 8.
    I: We provide additional results for all the experiments described in the paper.
Appendix B VDGD Computational Complexity
Given a LVLM with a vocabulary size $V$, an input instruction $I$ with a description of length $M$ tokens, at each decoding step the complexity for equation (2) is $O(V)$ and let the number of plausible tokens after this step be $K$, the complexity for equation (3) is $O(VMK)$. The total computational complexity of VDGD at each decoding step is $O(V)+O(VMK)\approx O(VMK)$.
Appendix C Broader Impact
This paper addresses the broader impact of reducing hallucinations in Large Vision-Language Models (LVLMs), which are increasingly utilized in diverse applications such as automated content generation, assistive technologies, and decision-making systems. By focusing on the phenomenon of hallucination, particularly in responses to cognitive prompts, our work has the potential to enhance the reliability and safety of LVLMs in practical scenarios. The proposed Visual Description Grounded Decoding (VDGD) method, being training-free, offers an accessible and scalable solution that can be readily implemented across various models to mitigate risks associated with inaccurate or misleading outputs. Furthermore, the introduction of the VaLLu benchmark contributes to the field by providing a tool for comprehensive evaluation of cognitive capabilities in LVLMs, promoting transparency and fostering further research in the area. Collectively, these contributions aim to improve user trust in AI systems and encourage responsible AI development that prioritizes factual accuracy and understanding.
Appendix D Prompts
For image description benchmarks we employ the prompt illustrated in Fig. 11. For non-image description benchmarks, we modify this prompt slightly (we modify the part of the prompt for extracting hallucinated phrases) and we illustrate the same in Fig. 15. The prompt used to generate image descriptions for VDGD is illustrated in Fig. 14. Correctness in the prompts refers to Factuality in all tables in Section I. We name it so after an ablation study on the terminology that works better for prompting the judge (GPT-4). All prompt evaluations on GPT-4 are done with a temperature of 0.7.
Figure 11: Prompt used to evaluate image description benchmarks using GPT-4-Turbo as described in Section 4.1.
Figure 12: Prompt used to evaluate the performance of LLMs on rephrased instructions using GPT-4-Turbo as described in Section 5.2.
Figure 13: Prompt used to identify the visual elements in images using Llama3 for the algorithm described in Section 4.2.
Figure 14: Prompt used to describe images using LVLMs for VDGD as described in Section 6.
Figure 15: Prompt used to evaluate generations from benchmarks like VaLLu, MMMU, MathVista, and SynthDoG using GPT-4-Turbo for the results in Table 2.
Figure 16: Prompt used to evaluate generations for rephrased instructions on benchmarks like VaLLu, MMMU, MathVista, and SynthDoG using GPT-4-Turbo as described in Section 5.2.
Appendix E Dataset Details
VaLLu We propose VaLLu benchmark which is sourced from Oven, MMMU, MMC, MathVista, HallusionBench, MATH-Vision and MME. This dataset is licensed under all the licenses of the original benchmarks that it was sourced from.
AMBER AMBER [29] is a benchmark for evaluating hallucinations in both the generative task and discriminative task of MLLMs. It provides comprehensive coverage of evaluations for various types of hallucination, including existence, attribute, and relation. Licensed under Apache-2.0 License.
MMMU MMMU [39] benchmark evaluates multimodal models on multi-discipline tasks and demands college-level subject knowledge and deliberate reasoning. It consists of 11.5k questions, spanning 30 subjects and 183 subfields, focusing on advanced perception and reasoning on domain-specific knowledge. Licensed under Apache-2.0 License.
MMC MMC [18] is a human-annotated benchmark comprising 9 distinct tasks to evaluate LVLM’s reasoning capabilities to comprehend visual charts. It comprises tasks such as chart information extraction, chart reasoning, contextual chart understanding, etc. It also offers two evaluation methods: GPT4 free-format general ability evaluation and multiple-choice QA format chart understanding. Licensed for research purposes.
MathVista MathVista [23] is a robust mathematical reasoning evaluation benchmark that consists of challenging tasks that require fine-grained, deep visual recognition and compositional reasoning. It consists of 6141 examples derived from 31 multimodal datasets involving mathematics. Licensed under CC-BY-SA-4.0.
MATH-Vision MATH-Vision [30] dataset is a collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions. The dataset spans 16 distinct mathematical disciplines and graded across 5 levels of difficulty to evaluate the mathematical reasoning abilities of Multimodal Large Language Models(MLLM). Licensed under MIT License.
HallusionBench HallusionBench [9] is a comprehensive benchmark designed for the evaluation of image-context reasoning. The benchmark presents significant challenges to advanced large visual-language models (LVLMs), such as GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA- 1.5, by emphasizing nuanced understanding and interpretation of visual data. The benchmark comprises 346 images paired with 1129 questions, all meticulously crafted by human experts. Licensed under BSD-3-Clause.
MME Multimodal large language model Evaluation Benchmark (MME) [6] measures perception and cognition abilities on a total of 2 tasks(Perception and Congition) and 14 subtasks(Existence, Count, Position, Color, Poster, Celebrity, Scene, Landmark, Artwork, OCR, Commonsense Reasoning, Numerical Calculation, Text Translation and Code Reasoning). To prevent data leakage from use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed.
Oven Open-domain Visual Entity Recognition (OVEN) [12] is a task where a model needs to link an image onto a Wikipedia entity with respect to a text query. OVEN challenges models to select among six million possible Wikipedia entities, making it a general visual recognition benchmark with the largest number of labels. Licensed under MIT License.
SynthDoG SynthDoG [14] is created by generating synthetic data which consists of various components: background, document, text, and layout. The background image is sampled from ImageNet, and the document is sampled from the collected paper photos. The text is sampled from Wikipedia. Layout is constructed using randomly stacks grids. The english variant of the dataset consists of 65.5k training and 500 validation entries. Licensed under MIT License.
Appendix F Model Details
LLaVa-v1. LLaVa-v1 [22] [^1] https://huggingface.co/liuhaotian/llava-llama-2-7b-chat-lightning-lora-preview is an open-source LVLM trained by fine-tuning LLaMA/Vicuna (13B) on GPT-generated multimodal instruction-following data. It is initially trained on 558K filtered image-text pairs from LAION/CC/SBU and then, finetuned on 80K GPT-generated multimodal instruction-following data. Licensed under CC BY-SA 4.0 DEED.
LLaVa-1.5. LLaVa-1.5 [20] [^2] https://huggingface.co/liuhaotian/llava-v1.5-7b builds on LLaVA-v1’s architecture by replacing linear projection design with a two-layer MLP. It also includes additional academic-task-oriented VQA datasets for VQA, OCR, and region-level perception, to enhance the model’s multimodal capabilities. Licensed under CC BY-SA 4.0 DEED.
LLaVa-1.6. LLaVa-1.6 [21] [^3] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b re-uses the pretrained connector of LLaVA-1.5 and is instruction tuned on 158K GPT-generated multimodal data, 500K academic-task-oriented VQA data, 50K GPT-4V data and 40K ShareGPT data. LLaVa-1.6 has improved reasoning, OCR, and world knowledge. Licensed under CC BY-SA 4.0 DEED.
mPLUG-Owl2. mPLUG-Owl2 [36] [^4] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b is the first Multi-modal Large Language Model that performs well in both pure-text and multi-modal scenarios. mPLUG-Owl2 is pretrained 400 million image-text pairs from: Conceptual Captions (CC3M/CC12M), COCO, Laio-en, COYO, DataComp and for instruction tuning 5 types of data are used: image captioning, mage question answering, region-aware QA, multi-modal instruct data and text-only instruct. Licensed under MIT License.
InternLM-X. InternLM-X [40] [^5] https://huggingface.co/internlm/internlm-xcomposer2-vl-7b is a vision-language large model based on InternLM2 [3] for advanced text-image comprehension and composition. InternLM-X is pre-trained on 1.1 billion images alongside 77.7 billion text tokens, including both public datasets and in-house concept data followed by supervised fine-tuning using multi-task training and instruction tuning. Licensed under Apache-2.0 License.
CogVLM. CogVLM [31] [^6] https://huggingface.co/THUDM/cogvlm-chat-hf is a powerful open-source visual language model. CogVLM-17B has 10 billion vision parameters and 7 billion language parameters. CogVLM is pre-trained using 1.5B text-image pairs and instruction tuned on various visual question-answering datasets. CogVLM achieves state-of-the-art performance on 10 classic cross-modal benchmarks. Licensed under Apache-2.0 License.
Llama-2-7b Llama-2-7b [28] [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf is an open source fine-tuned generative text model with 7 billion parameters. It is an auto-regressive language model that uses an optimized transformer architecture. The fine-tuned version use supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) for helpfulness and safety. Licensed under Llama 2 community license agreement.
Vicuna-7b-v1.5 Vicuna-7b-v1.5 [45] [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5 is a 7b parameter open-sourced model and fine-tuned from Llama 2 with supervised instruction fine-tuning and is an auto-regressive language model based on the transformer architecture. The training data is around 125K conversations from ShareGPT. Licensed under Apache-2.0 License.
InternLM2-7B InternLM-7B [3] [^9] https://huggingface.co/internlm/internlm2-chat-7b InternLM2 is a open-source 7 billion parameter chat model with a 200k context window and achieves good performance in the“Needle-in-a-Haystack” test(on long-context tasks like LongBench and L-Eval) and is trained using Conditional Online RLHF (COOL RLHF). Licensed under Apache-2.0 License.
Appendix G Baseline Details
VCD. Visual Contrastive Decoding (VCD) [26] is a training-free method that contrasts output distributions of original/unaltered inputs and distorted visual inputs. VCD hypothesizes statistical bias and unimodal priors as the two essential causes of object hallucinations and reduces the dependence on them. This results in the generated content being closely grounded to visual inputs resulting more accurate outputs. Licensed under Apache-2.0 License.
OPERA. Over-trust Penalty and a Retrospection-Allocation (OPERA) [13] is a training-free approach that hypothesizes that hallucinations are closely tied to the knowledge aggregation patterns in the self-attention matrix and that generation of new tokens by models is majorly dependent on a few summary tokens(over-trust issue) but not all of the previous tokens. This methodology introduces a penalty term on the model logits during beam-search decoding to mitigate the over-trust issue and implements a rollback strategy to retrospect the presence of summary tokens. Licensed under MIT License.
Woodpecker. Woodpecker [38] is a training-free method that identifies and corrects hallucinations in the generated text. Unlike other training-free methodologies, Woodpecker works after the model generation process. This method works by creating a visual knowledge base of Question-Answer(QA) pairs which contain object-level and attribute-level claims about the input image followed by an LLM correcting hallucinations.
LRV. LRV [17], or Large-scale Robust Visual Instruction tuning, proposes a dataset of visual instructions generated by GPT4, covering 16 vision-and-language tasks with open-ended instructions and answers. Each instance has positive instruction samples and negative instructions for more robust visual instruction tuning. As a result, a model trained on this dataset proves to be robust to hallucinations. Licensed under BSD-3-Clause license.
LURE. LVLM Hallucination Revisor (LURE) [47] attributes object hallucination to three factors: co-occurrence, uncertainty, and object position, and develops an object hallucination revisor which converts potential hallucinations into accurate outputs. This method first creates a hallucination dataset by making modifications to the original data and trains the object hallucination revisor on this dataset. This method post-hoc rectifies object hallucinations in LVLMs by reconstructing less hallucinatory descriptions.
Appendix H Additional Details
H.1 Logit Space for VR and Cognitive Prompts
From our in-depth analysis of the logit space, we see that language hallucinations in LVLMs have a unique characteristic – the probability of the most confident token in the vocabulary marginally decreased from the probability of the same token in the base LLM when prompted with the same prior context (the most confident token by the LVLM will also be the most confident token by the language-only base LLM as the token on a language hallucination). Thus, decoding-based mitigation methods like OPERA and VDD benefit from this unique property, and contrasting tokens help mitigate hallucinations.
H.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
In this subsection, we illustrate several instances (figures 17, 20, 21, 18, 19) from VaLLu and compare their responses for LLaVa-1.5 with Greedy, VCD, and VDGD Decoding.
Figure 17: Qualitative example 1 for VaLLu benchmark.
Figure 18: Qualitative example 2 for VaLLu benchmark.
Figure 19: Qualitative example 3 for VaLLu benchmark.
Figure 20: Qualitative example 4 for VaLLu benchmark.
Figure 21: Qualitative example 5 for VaLLu benchmark.
H.3 Noisy Examples in Existing Datasets
To maintain a high quality of the benchmark and eliminate noisy examples we performed a manual review of the existing datasets. This was crucial to ensure that the data used in our paper was reliable. It was also essential for the validity of our research findings. We removed questions that did not have a definite ground truth answer. We also made sure that the questions were answerable using a combination of general knowledge and reasoning. This filtering was done by the paper authors and has been approved by our institution’s Institutional Review Board (IRB). We share a few examples (figures: 22, 23, 24, 25, 26, 27, 28, 29) of noise in the existing datasets below:
Figure 22: Noisy example 1 for MMMU benchmark.
Figure 23: Noisy example 2 for MMMU benchmark.
Figure 24: Noisy example 1 for MMC benchmark.
Figure 25: Noisy example 2 for MMC benchmark.
Figure 26: Noisy example 1 for MathVista benchmark.
Figure 27: Noisy example 2 for MathVista benchmark.
Figure 28: Noisy example 1 for HallusionBench benchmark.
Figure 29: Noisy example 1 for HallusionBench benchmark.
H.4 Additional Details on Algorithm 1
In this Section, we describe our algorithm verbally.
[…]
Obtaining Hallucinated Phrases. Fig. 11 illustrates the prompt used for AMBER evaluation, where an LVLM is prompted for a description of an input image. In addition to scores, our judge, GPT-4, also returns the exact hallucinated phrases by the LVLM. We divide these hallucinated phrases into 3 types: Object Hallucinations – There are hallucinated visual elements which are objects. Relation HallucinationsAction/Verb Hallucination – Hallucinated visual elements that are not objects but just actions or verbs, e.g., walking, etc. Relation Hallucination – Hallucinated visual elements that define relationships (spatial or other kinds) between objects.
Filtering visual elements in Hallucinated Phrases. We prompt LLaMa-3 with the image description output by the LVLM, to extract the phrases from the image description that are visual elements. This ensures filtering out stop-words and all other words that are not visual elements and thus aids our algorithm in getting a precise count of each category of hallucination.
Causal Categorization of Hallucinations. Next, based on the Base Rank formulation and the exact hallucinated phrases detected by our judge (GPT-4), we propose an algorithm (Algorithm 1) to automatically categorize identified hallucinations into the 4 different categories. Precisely, for every hallucinated phrase returned by the judge, we first check if the word in the phrase is a visual element. We do this using the list of visual elements output by LLaMa-3. Next, for the first token of each word of the hallucinated phrase, we check the Base Rank of the token.
  1. 1.
    If the Base Rank is 0, we attribute the phrase to be a language hallucination. This is as simple as the phrase was forced to be generated by the base LLM itself without looking at the image.
  2. 2.
    Next, if the Base Rank is more than 0 and the hallucination can be found in the top-k retrieved elements from the IT dataset the model was aligned using, we attribute it to be an IT hallucination [8]. To retrieve top-k image-instruction pairs that are closest to the prompt image, we employ CLIP [25] similarity. Specifically, we calculate the similarity between the image embedding of the image associated with the input prompt and the text embedding of the description/response associated with the IT instances. Next, for every prompt, we retrieve the top-25 IT instances and search the hallucinated phrase returned by the judge (in descriptions/responses) using string matching.
  3. 3.
    Next, if the Base Rank is more than 0, the hallucination is not an IT hallucination and the hallucination is a Relation Hallucination, we attribute it to be a style hallucination. LVLMs fine-tuned on datasets synthesized from stronger closed-source LVLMs generally learn to mimic their style. One such style is adding more details to an identified object, by relating it to other objects or world knowledge. Our in-depth analysis shows that LVLMs, striving for lengthier, more detailed responses to imitate style, hallucinate and generate wrongful relations for identified objects. Our relation hallucinations returned by the judge precisely capture this, and thus we attribute relation hallucinations with a base rank of more than 0 to be Style hallucinations.
  4. 4.
    Finally, if the Base Rank is more than 0, the hallucination is identified as an Obejct hallucination by the judge, we attribute it to be a Vision Hallucination. These are caused by ineffective recognition of objects in the image (due to low-resolution images, weak encoder, etc.)
H.5 Elbow Method for Logit Analysis
The Knee/Elbow Detection Method operates by plotting the cumulative distribution of the sorted probabilities and seeking the point that maximizes the distance from the line connecting the first and the last points of this cumulative plot. Conceptually, this corresponds to finding the point of maximum curvature on the plot, indicative of the most pronounced change in the distribution’s progression. It offers a robust approach by identifying the point in a sorted probability vector where the rate of decrease in probabilities shows a significant bend or ’elbow’. This method is particularly useful when the probabilities are distributed such that the top few are similar and significantly higher than the rest, which then decrease sharply. The method proves to be effective in our experiments with multinomial sampling from a distribution, where it allows us to focus on the probabilities that correspond to the tokens that are most likely to be selected by sampling. This method ensures that the chosen cutoff $k$ dynamically adjusts to the specific characteristics of each probability distribution encountered, aligning with the overarching goal of maintaining sampling relevance and computational efficiency.
H.6 Examples of Rephrased Prompts
To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we ask GPT-4 to rewrite the instructions in prompts in such a way that the base LLM can answer the question without any image input as described in section 5.2. We share a few examples (figures: 30, 31, 32, 33, 34, 35, 36, 37) of rephrased prompts in the existing datasets below:
Figure 30: Rephrased prompt example for VaLLu benchmark.
Figure 31: Rephrased prompt example for VaLLu benchmark.
Figure 32: Rephrased prompt example for MMMU benchmark.
Figure 33: Rephrased prompt example for MMMU benchmark.
Figure 34: Rephrased prompt example for MathVista benchmark.
Figure 35: Rephrased prompt example for MathVista benchmark.
Figure 36: Rephrased prompt 1 for SynthDoG benchmark.
Figure 37: Rephrased prompt 2 for SynthDoG benchmark.
Methodology VP VP+Info-Seek VP+Info-Seek+Reasoning VP+Reasoning
LLaVA-v1 1.62 2.60 1.67 1.70
LLaVA-v1.5 2.59 3.18 1.50 1.76
LLaVA-v1.6 3.46 3.68 1.67 2.15
mPLUG-Owl2 2.56 3.23 1.67 1.94
InternLM-X 3.75 3.90 2.33 2.56
CogVLM 3.00 3.37 1.25 2.08
Table 3: Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark for the "Factuality" metric.
Dataset Model k Variance Range Avg.
AMBER LLaVA-v1 4.19 0.05 0.31 0.23
LLaVA-v1.5 3.91 0.07 0.39 0.29
LLaVA-v1.6 3.78 0.08 0.43 0.32
mPLUG-Owl2 3.65 0.06 0.36 0.28
InternLM-X 3.56 0.09 0.44 0.36
CogVLM 3.16 0.91 0.47 0.38
MMMU LLaVA-v1 3.68 0.08 0.39 0.27
LLaVA-v1.5 3.53 0.09 0.41 0.23
LLaVA-v1.6 3.40 0.09 0.42 0.24
mPLUG-Owl2 3.32 0.10 0.44 0.26
InternLM-X 3.26 0.14 0.46 0.28
CogVLM 3.16 0.19 0.47 0.25
Table 4: Post-truncation probability statistics for various models using the elbow method. All values are computed across all hallucinated tokens in each dataset.
Dataset Model Helpfulness Clarity Factuality Depth Engagement
VaLLu LLaVA-v1.5-t 3.05 4.02 2.83 2.70 3.85
mPLUG-Owl2-t 3.70 4.15 3.14 3.11 3.24
InternLM-X-t 4.28 4.58 3.66 3.62 3.57
MMMU LLaVA-v1.5-t 1.97 3.25 1.69 1.87 3.25
mPLUG-Owl2-t 2.48 3.05 1.63 2.22 2.64
InternLM-X-t 3.51 3.93 2.43 2.98 3.01
MathVista LLaVA-v1.5-t 2.53 3.58 2.21 2.10 3.44
mPLUG-Owl2-t 3.56 4.05 3.04 2.52 2.79
InternLM-X-t 4.39 4.65 3.99 2.91 2.95
Oven LLaVA-v1.5-t 3.38 4.48 2.95 2.32 3.35
mPLUG-Owl2-t 3.62 4.56 2.99 2.47 3.37
InternLM-X-t 3.80 4.58 3.76 2.89 3.52
Table 5: Performance comparison of various LVLMs on rephrased prompts without images as described in Section 5.2
Dataset Model Helpfulness Clarity Factuality Depth Engagement
AMBER LLaVA-v1 3.79 4.46 3.2 3.27 0.96
LLaVA-1.5 4.29 4.73 3.91 3.44 1.26
LLaVA-1.6 4.30 4.78 3.85 4.29 1.27
mPLUG-Owl2 4.02 4.58 3.48 3.27 0.71
InternLM-X 4.59 4.86 4.37 4.00 1.40
CogVLM 4.68 4.93 4.49 3.96 1.28
GPT-4-Turbo 4.89 4.96 4.79 4.64 1.50
OCR LLaVA-v1 1.11 2.09 1.11 1.14 0.47
LLaVA-1.5 1.07 2.24 1.17 1.15 0.48
LLaVA-1.6 2.51 2.92 2.63 1.99 1.11
mPLUG-Owl2 1.21 1.15 1.16 1.10 0.64
InternLM-X 1.11 1.12 1.12 1.03 0.67
CogVLM 3.12 3.04 2.68 2.53 1.47
GPT-4-Turbo 4.15 4.02 3.78 3.98 1.67
MMMU LLaVA-v1 1.77 2.73 1.50 1.65 0.46
LLaVA-1.5 2.45 3.34 2.00 2.23 0.59
LLaVA-1.6 3.54 3.98 2.95 3.39 0.99
mPLUG-Owl2 2.94 3.64 2.56 2.33 0.82
InternLM-X 3.60 4.10 3.24 3.15 1.05
CogVLM 4.03 4.38 3.57 3.43 1.03
GPT-4-Turbo 4.86 4.92 4.73 4.56 1.15
MathVista LLaVA-v1 1.95 2.84 1.95 1.85 0.76
LLaVA-1.5 2.57 3.47 2.23 2.36 0.98
LLaVA-1.6 3.68 4.07 3.17 3.48 1.12
mPLUG-Owl2 3.14 3.85 2.78 2.46 1.03
InternLM-X 3.91 4.32 3.45 3.35 1.23
CogVLM 4.20 4.52 3.78 3.66 1.24
GPT-4-Turbo 4.54 4.97 4.12 4.10 1.35
MMC LLaVA-v1 1.81 2.90 1.51 1.75 0.56
LLaVA-1.5 2.66 3.49 2.21 2.42 0.71
LLaVA-1.6 4.09 4.32 3.50 3.90 1.09
mPLUG-Owl2 3.63 4.15 3.32 2.60 0.97
InternLM-X 3.72 4.17 3.28 3.25 1.10
CogVLM 4.27 4.58 3.91 3.92 1.32
GPT-4-Turbo 4.81 4.89 4.57 4.49 1.12
Table 6: Performance comparison of various LVLMs on popular benchmarks for image description. While AMBER has real-world scenes, others do not. Experiment described in Section 4.1.
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25
Vanilla-sampling 2.64 3.73 1.86 1.94 1.27
VDD 1.70 2.55 1.47 1.81 1.59
OPERA 2.60 3.75 2.04 1.89 1.30
Woodpecker 2.68 3.81 2.01 1.95 1.28
LRV 2.64 3.78 1.98 1.94 1.36
LURE 2.60 3.54 2.03 1.89 1.40
VDGD (ours) 2.38 3.39 2.16 2.46 1.34
LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25
Vanilla-sampling 2.57 3.65 2.01 1.92 1.22
VDD 1.94 2.99 1.55 1.65 1.05
OPERA 2.73 3.74 2.05 2.04 1.22
Woodpecker 2.74 3.87 2.05 2.03 1.29
LRV 2.76 3.82 2.10 2.03 1.28
LURE 2.72 3.76 2.03 2.05 1.27
VDGD (ours) 2.97 3.91 2.64 2.27 1.54
LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43
Vanilla-sampling 3.10 3.85 2.64 2.63 1.40
VDD 2.01 2.96 1.80 2.10 1.20
OPERA 3.10 4.01 2.62 2.60 1.42
Woodpecker 3.16 4.00 2.67 2.58 1.44
LRV 3.15 3.98 2.65 2.62 1.39
LURE 3.11 4.02 2.64 2.64 1.48
VDGD (ours) 3.51 4.04 3.16 2.87 1.58
mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28
Vanilla-sampling 2.72 3.60 2.18 1.91 1.24
VDD 2.16 2.95 1.80 1.85 1.26
OPERA 2.92 3.96 2.26 2.08 1.34
Woodpecker 2.89 3.90 2.23 2.02 1.24
LRV 2.90 3.94 2.19 2.05 1.28
LURE 2.92 3.89 2.24 2.06 1.32
VDGD (ours) 3.14 3.98 2.72 2.22 1.50
InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47
Vanilla-sampling 3.13 4.04 2.70 2.41 1.56
VDD 2.56 3.26 2.32 2.19 1.41
OPERA 3.14 4.20 2.65 2.32 1.46
Woodpecker 3.13 4.12 2.60 2.28 1.41
LRV 3.06 4.19 2.59 2.31 1.45
LURE 3.12 4.10 2.64 2.40 1.49
VDGD (ours) 3.57 4.37 3.45 2.65 1.68
CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55
Vanilla-sampling 3.26 4.21 2.83 2.40 1.55
VDD 2.80 3.62 2.63 2.10 1.58
OPERA 3.40 4.27 2.85 2.46 1.57
Woodpecker 3.44 4.30 2.91 2.45 1.59
LRV 3.42 4.26 2.88 2.42 1.54
LURE 3.39 4.20 2.78 2.41 1.52
VDGD (ours) 3.33 4.15 3.01 2.51 1.60
GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
Table 7: Performance comparison of various LVLMs on VaLLu benchmark. This is an extension of results shown in Table 2
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00
Vanilla-sampling 1.82 2.82 1.38 1.56 1.06
VDD 1.70 2.67 1.22 1.40 0.95
OPERA 1.88 2.87 1.45 1.61 1.07
Woodpecker 1.89 2.88 1.48 1.60 1.06
LRV 1.83 2.85 1.40 1.57 1.08
LURE 1.91 2.90 1.52 1.63 1.09
VDGD (ours) 2.09 2.99 1.64 1.74 1.14
LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09
Vanilla-sampling 1.83 2.88 1.42 1.67 1.08
VDD 1.70 2.52 1.34 1.58 0.97
OPERA 2.18 3.08 1.62 1.75 1.22
Woodpecker 2.15 2.93 1.57 1.63 1.18
LRV 2.16 2.88 1.60 1.67 1.16
LURE 2.03 2.87 1.46 1.71 1.15
VDGD (ours) 2.30 3.17 1.85 1.86 1.20
LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97
Vanilla-sampling 1.80 2.63 1.50 1.78 0.96
VDD 1.68 2.57 1.36 1.68 0.94
OPERA 1.86 2.70 1.57 1.84 0.99
Woodpecker 1.90 2.74 1.63 1.89 1.04
LRV 1.88 2.72 1.61 1.85 1.02
LURE 1.92 2.73 1.65 1.86 1.03
VDGD (ours) 2.11 2.89 1.81 1.92 1.05
mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16
Vanilla-sampling 2.25 3.12 1.59 1.66 1.14
VDD 2.10 2.98 1.42 1.56 1.04
OPERA 2.35 3.23 1.67 1.74 1.18
Woodpecker 2.28 3.17 1.61 1.69 1.15
LRV 2.34 3.24 1.66 1.72 1.14
LURE 2.32 3.21 1.63 1.67 1.19
VDGD (ours) 2.48 3.36 1.82 1.78 1.23
InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63
Vanilla-sampling 2.97 3.76 2.17 2.45 1.64
VDD 2.76 3.54 2.03 2.26 1.57
OPERA 3.09 3.86 2.20 2.43 1.66
Woodpecker 3.07 3.83 2.19 2.46 1.68
LRV 3.11 3.88 2.24 2.49 1.70
LURE 3.10 3.89 2.22 2.48 1.71
VDGD (ours) 3.21 4.11 2.57 2.78 1.79
CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22
Vanilla-sampling 2.44 3.37 1.68 1.84 1.23
VDD 2.22 3.15 1.53 1.68 1.17
OPERA 2.56 3.47 1.76 1.88 1.27
Woodpecker 2.57 3.47 1.78 1.90 1.26
LRV 2.58 3.49 1.80 1.92 1.25
LURE 2.54 3.42 1.75 1.85 1.22
VDGD (ours) 2.65 3.56 1.92 1.97 1.29
GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
Table 8: Performance comparison of various LVLMs on MMMU benchmark. This is an extension of results shown in Figure 7.
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42
Vanilla-sampling 3.72 4.56 2.24 2.27 1.43
VDD 3.54 4.32 2.05 2.07 1.36
OPERA 3.80 4.63 2.29 2.35 1.50
Woodpecker 3.78 4.60 2.26 2.30 1.47
LRV 3.76 4.64 2.30 2.38 1.49
LURE 3.74 4.59 2.25 2.28 1.46
VDGD (ours) 3.92 4.78 2.52 2.42 1.54
LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32
Vanilla-sampling 3.68 4.39 2.38 2.04 1.36
VDD 3.38 4.09 2.35 1.87 1.10
OPERA 3.74 4.46 2.26 1.93 1.32
Woodpecker 3.46 4.53 2.34 1.97 1.38
LRV 3.53 4.42 2.39 1.96 1.27
LURE 3.33 4.41 2.37 1.91 1.29
VDGD (ours) 3.84 4.61 2.56 2.15 1.45
LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50
Vanilla-sampling 3.69 4.32 2.45 2.28 1.48
VDD 3.54 4.11 2.34 2.18 1.35
OPERA 3.80 4.47 2.59 2.46 1.57
Woodpecker 3.78 4.42 2.56 2.43 1.53
LRV 3.75 4.40 2.54 2.42 1.52
LURE 3.79 4.41 2.52 2.41 1.53
VDGD (ours) 3.90 4.53 2.79 2.51 1.58
mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33
Vanilla-sampling 3.80 4.45 2.51 2.12 1.34
VDD 3.62 4.27 2.35 1.98 1.24
OPERA 3.85 4.52 2.57 2.20 1.40
Woodpecker 3.84 4.56 2.60 2.24 1.42
LRV 3.82 4.54 2.59 2.22 1.41
LURE 3.86 4.58 2.61 2.25 1.45
VDGD (ours) 3.99 4.65 2.76 2.36 1.52
InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55
Vanilla-sampling 4.01 4.68 2.83 2.37 1.54
VDD 3.82 4.37 2.66 2.10 1.44
OPERA 4.13 4.76 2.92 2.42 1.59
Woodpecker 4.07 4.72 2.88 2.40 1.58
LRV 4.10 4.73 2.89 2.40 1.56
LURE 4.14 4.75 2.91 2.43 1.60
VDGD (ours) 4.28 4.89 3.12 2.54 1.66
CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44
Vanilla-sampling 3.73 4.38 2.74 2.61 1.42
VDD 3.58 4.27 2.64 2.50 1.37
OPERA 3.76 4.45 2.82 2.69 1.47
Woodpecker 3.75 4.44 2.79 2.64 1.45
LRV 3.80 4.51 2.86 2.70 1.49
LURE 3.82 4.53 2.88 2.73 1.52
VDGD (ours) 3.94 4.64 2.99 2.80 1.60
GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61
Table 9: Performance comparison of various LVLMs on MathVista benchmark. This is an extension of results shown in Figure 7.
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23
Vanilla-sampling 3.14 4.42 2.45 2.20 1.26
VDD 2.98 4.10 2.20 2.07 1.16
OPERA 3.20 4.48 2.50 2.22 1.28
Woodpecker 3.21 4.49 2.46 2.21 1.25
LRV 3.25 4.51 2.51 2.23 1.29
LURE 3.20 4.47 2.49 2.20 1.28
VDGD (ours) 3.42 4.59 2.78 2.30 1.30
LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19
Vanilla-sampling 3.24 4.40 2.65 2.58 1.10
VDD 2.82 4.10 2.44 2.39 1.20
OPERA 3.22 4.40 2.49 2.58 1.18
Woodpecker 3.26 4.42 2.46 2.64 1.22
LRV 3.25 4.38 2.52 2.59 1.24
LURE 3.24 4.35 2.51 2.60 1.21
VDGD (ours) 3.47 4.56 2.84 2.51 1.80
LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35
Vanilla-sampling 3.40 4.38 3.10 3.27 1.30
VDD 3.26 4.19 2.96 3.13 1.20
OPERA 3.49 4.48 3.20 3.38 1.40
Woodpecker 3.45 4.45 3.21 3.35 1.42
LRV 3.42 4.40 3.18 3.34 1.39
LURE 3.38 4.36 3.18 3.28 1.36
VDGD (ours) 3.64 4.71 3.49 3.42 1.45
mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25
Vanilla-sampling 3.24 4.47 2.84 2.50 1.24
VDD 3.07 4.21 2.67 2.40 1.17
OPERA 3.29 4.52 2.92 2.58 1.27
Woodpecker 3.27 4.50 2.90 2.52 1.25
LRV 3.23 4.49 2.88 2.50 1.24
LURE 3.26 4.51 2.91 2.51 1.28
VDGD (ours) 3.34 4.64 3.15 2.70 1.32
InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37
Vanilla-sampling 3.27 4.53 3.33 2.57 1.35
VDD 3.08 4.24 3.10 2.44 1.30
OPERA 3.37 4.65 3.44 2.65 1.43
Woodpecker 3.38 4.63 3.45 2.66 1.44
LRV 3.40 4.69 3.47 2.70 1.46
LURE 3.36 4.64 3.42 2.68 1.42
VDGD (ours) 3.54 4.78 3.68 2.79 1.49
CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38
Vanilla-sampling 3.59 4.50 3.37 2.78 1.41
VDD 3.36 4.32 3.21 2.58 1.30
OPERA 3.68 4.60 3.42 2.84 1.45
Woodpecker 3.70 4.58 3.45 2.85 1.42
LRV 3.65 4.57 3.40 2.82 1.40
LURE 3.72 4.62 3.46 2.83 1.41
VDGD (ours) 3.89 4.82 3.59 2.89 1.49
GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
Table 10: Performance comparison of various LVLMs on Oven benchmark. This is an extension of results shown in Figure 7.
H.7 Compute Infrastructure
All our analysis, inference and baseline experiments are conducted on a node of 4 NVIDIA RTX A6000 GPUs, with 128GB RAM and 10 CPU cores. The evaluations are conducted using gpt-4-turbo-2024-04-09 model.
Appendix I Additional Results
I.1 Comparison of various LVLMs on rephrased prompts
Table 5 provides quantitative values for rephrased prompts without image input for various LVLMs, detailed in section 5.2 of the main paper.
I.2 Comparison of various LVLMs on image description
Table 6 provides quantitative values for faithful image description generations for various LVLMs on multiple benchmarks, detailed in section 5.2 of the main paper.
I.3 Comparison of post-truncation probability statistics on AMBER and MMMU benchmarks
Table 4 provides post-truncation probabilities for all LVLMs on AMBER and MMMU, this is an extension of the statistic shown in section 5.3 of the main paper.
I.4 Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark
Table 3 provides comparison of LVLMs on different capabilities on the proposed VaLLu benchmark.
I.5 Comparison of VaLLu with Mitigation Baselines
Table 7 provides quantitative values for comparison of VDGD and other baselines on the VaLLu benchmark, detailed in section 6.3 of the main paper.
I.6 Comparison of MMMU with Mitigation Baselines
Table 8 provides quantitative values for comparison of VDGD and other baselines on the MMMU benchmark, detailed in section 4.1 of the main paper.
I.7 Comparison of MathVista with Mitigation Baselines
Table 9 provides quantitative values for comparison of VDGD and other baselines on the MathVista benchmark, detailed in section 4.1 of the main paper.
I.8 Comparison of Oven with Mitigation Baselines
Table 10 provides quantitative values for comparison of VDGD and other baselines on the Oven benchmark, detailed in section 4.1 of the main paper.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contribution and are true to the best of our knowledge.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The paper’s last section i.e 7 Conclusion, Limitations and Future Work discusses the limitation of our work.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: This paper does not make any theoretical assumptions and does not describe any theoretical proofs.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: To the best of our knowledge we described our experimental methods, algorithms and evaluation prompts in the paper in detail.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We have released the code for all the experiments, algorithm and our novel contributions in an anonymous git reopsitory as described in 7. We have also uploaded the benchmark proposed in the paper under Supplementary Materials.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We have included all the evaluation prompts and hyperparameters used for testing. The paper does not provide any training methodologies and as such does not specify any training details.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We have provided error bars for the results in the main paper. Due to the expensive nature of evaluation, we provided error bars only for the main results in the paper.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We have details about all the compute resources we have used in section H.7.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: To the best of our knowledge the paper adheres to all the ethical guidelines provided in the above link.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We have described the broader/societal impacts of the paper in section C.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [No]
  44. Justification: We do not release any new data or models. All data and models used are from existing resources. The VaLLu benchmark is an amalgamation of existing benchmarks. We use licenses and release the dataset under licenses issued by the original benchmark.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All the licenses for models, datasets and baselines used in the paper are provided in appendix sections F, E and G.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We introduced a new benchmark in this paper. The benchmark data and metadata file have been uploaded in the supplementary folder in an anonymized zip file.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [No]
  56. Justification: We did not use crowdsourcing and research with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [Yes]
  60. Justification: We obtained an IRB approval to filter out noisy examples in the datasets we used as described in section H.3.

### v2

@@
-VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap
+Visual Description Grounding Reduces Hallucinations and Boosts Reasoning in LVLMs
 Abstract
-Recent interest in Large Vision-Language Models (LVLMs) for practical applications is moderated by the significant challenge of hallucination or the inconsistency between the factual information and the generated text.
-In this paper, we first perform an in-depth analysis of hallucinations and discover several novel insights about how and when LVLMs hallucinate.
-From our analysis, we show that: (1) The community’s efforts have been primarily targeted towards reducing hallucinations related to visual recognition (VR) prompts (e.g., prompts that only require describing the image), thereby ignoring hallucinations for cognitive prompts (e.g., prompts that require additional skills like reasoning on contents of the image).
-(2) LVLMs lack visual perception, i.e., they can see but not necessarily understand or perceive the input image.
-We analyze responses to cognitive prompts and show that LVLMs hallucinate due to a perception gap: although LVLMs accurately recognize visual elements in the input image and possess sufficient cognitive skills, they struggle to respond accurately and hallucinate.
-To overcome this shortcoming, we propose Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method for alleviating hallucinations.
-Specifically, we first describe the image and add it as a prefix to the instruction.
-Next, during auto-regressive decoding, we sample from the plausible candidates according to their KL-Divergence (KLD) to the description, where lower KLD is given higher preference.
-Experimental results on several benchmarks and LVLMs show that VDGD improves significantly over other baselines in reducing hallucinations.
-We also propose VaLLu, a benchmark for the comprehensive evaluation of the cognitive capabilities of LVLMs.
+Large Vision-Language Models (LVLMs) often produce responses that misalign with factual information, a phenomenon known as hallucinations.
+While hallucinations are well-studied, the exact causes behind them remain underexplored.
+In this paper, we first investigate the root causes of hallucinations in LVLMs.
+Our findings reveal that existing mitigation techniques primarily reduce hallucinations for visual recognition prompts—those that require simple descriptions of visual elements—but fail for cognitive prompts that demand deliberate reasoning.
+We identify the core issue as a lack of true visual perception in LVLMs: although they can accurately recognize visual elements, they struggle to fully interpret these elements in the context of the input prompt and effectively link this recognition to their internal knowledge, which is critical for reasoning.
+To address this gap, we introduce Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method designed to enhance visual perception and improve reasoning capabilities in LVLMs.
+VDGD works by first generating a detailed description of the image and appending it as a prefix to the instruction.
+During response generation, tokens are sampled based on their KL divergence to the description, favoring candidates with lower divergence.
+Experimental results on multiple visual reasoning benchmarks and LVLMs demonstrate that VDGD consistently outperforms existing baselines 2% - 33%.
+Finally, we introduce VaLLu, a benchmark designed for comprehensive evaluation of the cognitive capabilities of LVLMs.
+Code is available here: https://github.com/Sreyan88/VDGD  [^1] Corresponding Author: sreyang@umd.edu.
+* denotes equal technical contribution.
 1 Introduction
-Large Language Models (LLMs) pre-trained at on web-scale text data with the next token prediction objective implicitly compress world knowledge in their parameters zhao2023survey .
-These models learn general-purpose representations, which can then be aligned with the desired response characteristics zhang2023instruction .
-This step generally involves fine-tuning on instruction-response pairs, also known as instruction tuning (IT) wei2022finetuned , and is followed by an optional step of reinforcement learning from human feedback (RLHF) bai2022training .
+Large Language Models (LLMs) pre-trained on web-scale text data with the next token prediction objective implicitly compress world knowledge in their parameters (Zhao et al., 2023).
+These models learn general-purpose representations, which can then be aligned with the desired response characteristics (Zhang et al., 2024b).
+This step generally involves fine-tuning on instruction-response pairs, also known as instruction tuning (IT) (Wei et al., 2022), and is followed by an optional step of reinforcement learning from human feedback (RLHF) (Bai et al., 2022).
 Such aligned LLMs can be engaged through text-based inputs, or prompts, directing them to perform various tasks, including those based on information retrieval and reasoning.
-Recent advancements in the research community have demonstrated success in multi-modal alignment, whereby fine-tuning LLMs with multi-modal instruction-response pairs enables them to execute tasks that conform to the text prompt as well as additional multi-modal inputs yin2023survey .
-However, similar to text-only LLMs, Multi-modal Large Language Models (MLLMs) are also prone to hallucination liu2024survey , or the misalignment between factual visual content and corresponding textual responses generated.
-This is a prominent concern that impacts the reliability and applicability of MLLMs across domains.
-In this paper, we investigate hallucinations in Large Vision Language Models (LVLMs) to explore the underlying causes and mechanisms.
-The research community has made significant attempts to reduce LVLM hallucinations liu2024survey .
-These attempts include better architectures liu2023improved ; wang2023cogvlm , better alignment techniques chen2024far , and explicit hallucination mitigation techniques yin2023woodpecker ; zhang2024debiasing ; wang2024mitigating .
-Each approach has demonstrated superior performance over others on benchmark datasets.
-In this paper, we conduct a comprehensive analysis of LVLM hallucinations to explore the mechanisms and conditions under which LVLMs hallucinate and to measure the community’s progress in addressing these issues.
-Our results reveal the following:
-  1. 1.
-    Prior work has majorly focused on alleviating hallucinations for visual recognition (VR) prompts (e.g., object hallucinations in prompts that demand image descriptions).
-While this has led to significant improvements in VR benchmarks (like AMBER), improvements in cognitive prompts that require additional skills, like reasoning and information seeking, have remained stagnant.
-Additionally, we perform an in-depth analysis to show that there has been relatively less progress in alleviating hallucinations for certain aspects (like descriptions of non-real world scenes) and categories (described in Section 4.2) of VR hallucinations.
-  2. 2.
-    For cognitive prompts we perform controlled experiments to show that LVLMs suffer from a visual perception gap: although they can accurately recognize visual elements in the input image and possess sufficient knowledge and reasoning skills to provide a factual response, they fail to fully perceive the image to generate a factual response.
-  3. 3.
-    Inspired by our findings, we propose a Visual Description Grounding Decoding (VDGD), a simple and novel training-free technique to alleviate LVLM hallucinations.
+Recent advancements in the research community have demonstrated success in multi-modal alignment, whereby fine-tuning LLMs with multi-modal instruction-response pairs enables them to execute tasks that conform to the text prompt as well as additional multi-modal inputs (Yin et al., 2023a).
+Since their introduction, the research community has extensively evaluated the capabilities of Large Vision-Language Models (LVLMs) across areas such as visual recognition (Wang et al., 2023a), perception (Fu et al., 2024), and reasoning (Yue et al., 2024).
+Researchers have made significant efforts to enhance these capabilities by refining model architectures and training techniques (Zhang et al., 2024a) or by implementing strategies to mitigate hallucinations  [^2] Hallucination refers to the mismatch between factual content and the model’s generated responses.
+Mitigation aims to reduce these discrepancies for more accurate outputs and improve task performance.
+(Huang et al., 2023a).
+Among these approaches, training-free hallucination mitigation techniques stand out as a cost-effective and accessible method to enhance LVLM performance without the need for extensive computational resources.
+However, current hallucination mitigation techniques primarily focus on reducing hallucinations and improving the performance of visual recognition tasks (e.g., describing a scene, OCR, etc).
+Our extensive experiments show that these techniques fall short when applied to cognitive prompts requiring reasoning or knowledge extraction, indicating a significant gap in their ability to address hallucinations in more complex, reasoning-intensive scenarios.
+To this end, we propose Visual Description Grounded Decoding (VDGD), a simple and novel training-free technique designed to improve LVLM reasoning by reducing hallucinations for cognitive prompts.
+Much like humans, who often write down their observations of an image in their own words and refer back to them while tackling complex reasoning tasks, we hypothesize that grounding the LVLM’s response generation in an explicit visual description can significantly aid its reasoning capabilities.
 Specifically, we first generate a description of the input image.
@@
 Finally, to predict the next word, we sample from the plausible candidates according to their KLD to the description, where lower KLD is given higher preference.
-VDGD improves performance on standard benchmarks over other mitigation techniques from literature.
-2 Experimental Setup
-LVLMs.
-For the scope of this work, we experiment with 6 different types of LVLMs, namely LLaVA-v1 liu2023visual , LLaVA-v1.5 liu2023improved , LLaVA-v1.6 liu2024llavanext , mPLUG-Owl2 ye2023mplugowl2 , InternLM-X zhang2023internlm and CogVLM wang2023cogvlm , each utilizing a 7B parameter language model, though the number of vision parameters varies.
-To accurately evaluate their capabilities in visual recognition and perception, we ensure that all models use the same foundational language model.
-More details can be found in Appendix F.
-Evaluation Setup.
-For evaluating LVLMs, we perform a mix of expert human evaluation and automatic evaluation using GPT-4 Turbo (gpt-4-turbo-2024-04-09).
-Recent research indicates that using ChatGPT and GPT-4 to score and assess outputs from LLMs aligns well with human evaluations, offering the added benefit of cost reduction zhou2023lima ; ghosh2024closer ; lin2023unlocking ; chan2024chateval ; xu-etal-2023-instructscore ; zheng2023judging .
-We devise an explainable and multi-aspect evaluation prompt lin2023unlocking ; ghosh2024closer , which prompts the judge (GPT-4) to assign a score between 1 and 5 to an LVLM’s response to an instruction on five aspects: helpfulness, clarity, factuality, depth, and engagement.
-It also provides an explanation for each score and extracts hallucinated phrases.
-Prompt is in Appendix D.
-With hallucinations as our primary focus, all results shown in the main paper are for the factuality metric.
-The metric is designed to penalize hallucinations and evaluate correctness in open-ended generations.
-Scores for all metrics and all graphs are in Appendix I.
-Evaluation Benchmarks.
-We employ a range of benchmarks for LVLM evaluation that assess the diverse capabilities of a model.
-These include AMBER wang2023llm (visual recognition), SynthDoG kim2022donut (OCR), Oven hu2023open (information seeking), MMMU yue2023mmmu (general and holistic reasoning), MMC liu2023mmc , MathVista lu2023mathvista and MATH-Vision wang2024measuring (mathematical reasoning), HallusionBench guan2023hallusionbench , and MME fu2023mme (chart understanding).
-We only use the sub-sets of these datasets that demand open-ended generation for response, as LLMs have been shown not to be robust to MCQ zheng2024large .
-Additionally, since automatic evaluation on all these datasets can be expensive, we propose VaLLu, detailed in Section 6.2.
-3 Current Progress on Solving LVLM Hallucinations
+By continuously referencing its descriptive summary of the visual content, the LVLM maintains alignment between the visual input and its generated responses, thereby reducing hallucinations.
+To summarize, our contributions are as follows:
+  1. 1.
+    We conduct extensive controlled experiments to investigate the causes of hallucinations in LVLMs, categorizing them into distinct types.
+We demonstrate that existing hallucination mitigation techniques are effective only for visual recognition (VR) prompts—specifically, those that involve object recognition and image descriptions—while proving ineffective for cognitive prompts that require reasoning.
+We further show that these techniques are biased towards reducing language hallucinations in real-world scenes but fail to address hallucinations in non-real-world contexts and other categories.
+  2. 2.
+    Identifying the core issue, we reveal that LVLMs suffer from a critical visual perception gap.
+Although these models can recognize visual elements and possess the reasoning skills needed for factual response generation, they struggle to fully perceive the visual information in the context of the prompt, resulting in hallucinated outputs for cognitive reasoning tasks.
+  3. 3.
+    We introduce VDGD, a novel and training-free method that enhances LVLM reasoning by grounding response generation in visual descriptions.
+We evaluate VDGD on X benchmarks with prompts that require deliberate reasoning and how VDGD significantly outperforms existing techniques, improving performance by 2% - 33% by reducing hallucinations.
+  4. 4.
+    Finally, we introduce VaLLu, a meticulously curated benchmark designed to evaluate the cognitive capabilities of LVLMs.
+Comprising 1,500 carefully selected instances, VaLLu is constructed by refining existing benchmarks and rigorously removing noisy and misleading examples identified through our thorough analysis.
+VaLLu emphasizes a balanced mix of diverse cognitive tasks, focusing exclusively on questions that require open-ended responses.
+2 Related Work
+Large Vision-Language Models In recent years, Large Vision-Language Models (LVLMs) have seen rapid advancements with the release of numerous new models.
+These developments have been driven by innovations in model architectures (Liu et al., 2023b; Ye et al., 2023; Wang et al., 2023b), training methods (Liu et al., 2023b), alignment techniques (Chen et al., 2024b), and data augmentation techniques (Liu et al., 2023b; Wang et al., 2024b).
+To assess the capabilities of these models, the research community has introduced a variety of benchmarks specifically designed to evaluate different aspects, including visual recognition and understanding in diverse scenarios (Liu et al., 2023a; Kim et al., 2022), reasoning (Lu et al., 2023; Yue et al., 2024), knowledge-based information retrieval (Hu et al., 2023), and other specialized tasks.
+Reducing Hallucinations to Improve Response Quality: Hallucination mitigation broadly aims to enhance the accuracy of model responses.
+While numerous studies have proposed methods to reduce hallucinations and improve response accuracy (Sicong Leng et al., 2023; Huang et al., 2023b; Yin et al., 2023b; Zhou et al., 2023b) and evaluated their effectiveness (Wang et al., 2023a; Guan et al., 2023), most focus on mitigating object hallucinations in visual recognition tasks.
+While these efforts are valuable, they fall short of assessing the true reasoning capabilities of LVLMs.
+Moreover, despite the growing focus on evaluation and mitigation, the underlying causes of hallucinations remain largely unexplored.
+3 How Close Are We to Reliable LVLM Responses?
+Experimental Setup for Analysis.
+We evaluate six LVLMs—LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, and CogVLM—all built on a 7B parameter language model.
+The models are tested across seven benchmarks: AMBER (visual recognition), SynthDoG (OCR), MMMU (expert-level reasoning), MathVista and MATH-Vision (mathematical reasoning), MMC (chart understanding), and MME and HallusionBench.
+To address potential limitations in existing evaluation metrics, which may not fully capture reasoning skills and often rely on string-matching techniques, we employ a combination of expert human evaluation and LLM-as-a-Judge (using GPT-4-turbo-2024-04-09).
+This paradigm has been extensively followed by prior studies in LLM evaluation (Zhou et al., 2023a; Ghosh et al., 2024; Zheng et al., 2023) and hallucination evaluation (Sicong Leng et al., 2023; Yu et al., 2023; Liu et al., 2023c).
+We devise an evaluation prompt that penalizes hallucinations and assigns scores from 1 to 5 based on factual correctness (see Appendix G).
+This approach also helps us standardize scoring across benchmarks and has a higher correlation ($\approx$0.97%) to expert human evaluations (see Section D).
+Quantitative results for all graphs in the paper are in Appendix L.
+Preliminaries.
 As LVLMs advance and become more accessible, their application extends beyond merely generating image descriptions.
-Thus, to better understand hallucinations, we break down information processing by an LVLM, when prompted with an input image and a text instruction, into its constituent steps:
+Thus, to better understand visual perception and reasoning, we break down information processing by an LVLM into its constituent steps:
 {mdframed}
@@
 […]
-As depicted in Fig.
-1, diverse prompts applied to a single image may evaluate the LVLM for various skills.
-Visual Recognition (VR) involves recognizing visual elements within the image (like objects, backgrounds, etc.), their interrelations, and their unique characteristics.
+[…]
+[trim= 0cm 2.25cm 0cm 1cm, width=]perception.pdf
 Figure 1: Depending on the text instruction, an LVLM might be assessed on one or more different capabilities.
-A prompt that just seeks a description of the image or details about a specific element explicitly evaluates only visual recognition capabilities.
-Visual Perception (VP), which follows and moves beyond VR, entails interpreting and contextualizing the observed elements in the context of the entire image fu2024blink .
-For tasks that extend beyond basic image description, accurate visual perception is crucial, surpassing the mere recognition of individual elements.
-If the prompt seeks further knowledge-specific information related to the elements in the image, then the next step involves extracting relevant information (IE) from the knowledge learned during pre-training.
-Finally, some prompts may necessitate the LVLM to reason over the visual data, textual prompt, and any implicitly extracted information to deduce a response.
-While VR relies on the performance of the vision encoder, the abilities to extract accurate information and reason are based on the strength of the foundational language model used in alignment tuning (as alignment does not teach a model new knowledge or skills ghosh2024closer ; zhou2023lima ).
-VP, which is done implicitly by the model, is a skill taught during alignment tuning and serves as a bridge between VR and other cognitive abilities.
-barrow1978recovering ; torralba2002depth ; hartley2003multiple
-A closer look at the progress in mitigating hallucinations. Fig. 2 (left) compares
+Figure 1 demonstrates that different prompts applied to a single image can assess various LVLM skills.
+Visual Recognition (VR) focuses on identifying visual elements and their relationships within the image, such as describing objects or specific details.
+Visual Perception (VP) extends beyond VR by interpreting and contextualizing these elements within the broader scene (Fu et al., 2024), essential for tasks requiring more than basic recognition.
+Prompts that demand knowledge-specific insights (also known as information-seeking prompts) engage in knowledge extraction (KE) learned from pre-training.
+Finally, prompts requiring reasoning involve combining visual data, textual prompts, and extracted knowledge to generate a response.
+While recognition depends on the vision encoder, knowledge extraction, and reasoning rely on the foundational language model.
+VP, implicitly learned during alignment (Ghosh et al., 2024; Zhou et al., 2023a), bridges VR and higher cognitive functions, facilitating comprehensive understanding of the image in context to the prompt (Barrow et al., 1978; Torralba & Oliva, 2002; Hartley & Zisserman, 2003).
+3.1 Do Hallucination Mitigation Techniques improve Cognitive Abilities?
+[…]
+[trim= 0cm 0.75cm 0cm 0.5cm, width=]radar1-2.pdf
 Figure 2: (Left) Performance comparison of different LVLMs on various benchmarks. (Right) Performance comparison of different hallucination mitigation techniques applied to LLaVA-1.5.
-the performance of LVLMs across 5 benchmarks.
+Fig.
+2 (left) compares the performance of LVLMs across 5 benchmarks.
 As we clearly see, while newer models that were trained using improved architectures and alignment training data boosted performance on the AMBER benchmark (for recognition), performance on other reasoning and information-seeking benchmarks remained stagnant.
 Fig.
-2 (right) compares the performance of LLaVA-1.5 across the same 5 benchmarks when employed with explicit hallucination mitigation techniques, namely, Visual Contrastive Decoding (VCD) damonlpsg2023vcd , OPERA huang2023opera , Woodpecker yin2023woodpecker and LURE zhou2023analyzing .
-As we clearly see, while all these methods boost performance on AMBER, similar to the previous experiments, performance on other reasoning and information-seeking benchmarks remains stagnant.
-{mdframed}
-[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Current advancements in reducing hallucinations via improved model architectures, training data, and explicit mitigation strategies are limited to improving the visual recognition capabilities of LVLMs.
-Progress in addressing hallucinations beyond those related to visual recognition has been relatively quite less.
-4 A Closer Look at Visual Recognition Hallucinations
-As we have seen in the earlier section, advancements in LVLMs and hallucination mitigation techniques have led to considerable improvements in visual recognition capabilities by these models.
-In the next few subsections, we take a closer look at visual recognition hallucinations to show that the models and methods discussed above only alleviate hallucinations in some specific cases.
-4.1 Moving Beyond Real-World Scenes
-As LVLMs are now being employed on a wide range of tasks, considerable research efforts are being made to evaluate the efficacy of LVLMs to reason on images beyond real world
+2 (right) compares the performance of LLaVA-1.5 across the same 5 benchmarks when employed with explicit hallucination mitigation techniques, with strategies ranging from improving model architectures to explicit response correction strategies, namely, Visual Contrastive Decoding (VCD) (Sicong Leng et al., 2023), OPERA (Huang et al., 2023b), Woodpecker (Yin et al., 2023b) and LURE (Zhou et al., 2023b).
+As we clearly see, while all these methods boost performance on AMBER, performance on other reasoning and information-seeking benchmarks remains stagnant.
+3.2 Do Mitigation Techniques Generalize Beyond Real-World Scenes?
+[…]
+[trim= 0cm 0.75cm 0cm 0.5cm, width=]radar3-4.pdf
 Figure 3: (Left) Performance comparison of different LVLMs on various benchmarks when prompted to only describe the image.
 (Right) Performance comparison of different hallucination mitigation techniques applied to LLaVA-1.5.
-scenes, i.e., charts liu2023mmc or math problems lu2024mathvista .
+As LVLMs are now being employed on a wide range of tasks, considerable research efforts are being made to evaluate the efficacy of LVLMs to reason on images beyond real-world scenes, i.e., charts (Liu et al., 2023a) or math problems (Lu et al., 2024).
 However, how often do LVLMs hallucinate while describing images beyond real-world scenes?
 Fig 3 (left) compares the performance of various LVLMs across 5 benchmarks.
-While MMMU and MathVista include images with mathematical figures, MMC includes images with charts from diverse sources, and SynthDoG prompts the model for OCR.
-For AMBER, MMMU, MathVista, and MMC, we prompt the LVLM with the prompt “Describe the image.” and for SynthDoG we prompt the LVLM with the prompt Extract text from scanned PDFs, photos, and even handwriting..
+MMMU and MathVista include images with mathematical figures and MMC includes charts from diverse sources.
+On the other hand, SynthDoG prompts the model for OCR and AMBER for image descriptions for real-world scenes.
 As we clearly see, compared to AMBER, models hallucinate more often when describing visuals beyond real-world scenes.
 Fig 3 (right) further shows that hallucination mitigation techniques improve performance on AMBER but rarely other benchmarks.
-4.2 Causal Categorization of Visual Recognition Hallucinations
+[…]
+[width=]vllm.drawio5.pdf
+Figure 4: Types of Visual Recognition Hallucinations.
+We define Algo.
+1 to divide VR hallucinations into 4 different categories automatically: Language, Vision, Style, and IT (explained further in Sec.
+3.3).
+While language and vision hallucinations have been explored earlier, and methods to alleviate them have been proposed, we show for the first time that Style and IT hallucinations exist and existing methods fail to alleviate them.
+3.3 Are all hallucinations equal?
 Overview.
 In our detailed analysis of model responses on the AMBER dataset, we identified intriguing patterns in which hallucinations occur.
-Although the community has primarily focused on the broader topic of object hallucinations liu2024survey , our findings reveal that visual recognition hallucinations can be classified into several subcategories based on their specific occurrence patterns.
+Although the community has primarily focused on the broader topic of object hallucinations (Liu et al., 2024b), our findings reveal that visual recognition hallucinations can be classified into several subcategories based on their specific occurrence patterns.
+Specifically, we perform the token-distribution analysis proposed by  (Lin et al., 2023).
 Algorithm 1 Categorizing Visual Hallucinations
@@
 […]
-(1) Language Hallucinations: These are hallucinations that are caused by language priors and are unshifted tokens (tokens that would have been output by the non-aligned LLM knowledge through prior context without an image).
-During auto-regressive decoding, the model might over-rely on the prior tokens in the response rather than the input image.
-Thus, this may leads the model to mention details in the response from its pre-trained knowledge that is not actually faithful to the image.
-Figure 4: Types of Visual Recognition Hallucinations.
-We define Algo.
-1 to divide VR hallucinations into 4 different categories automatically: Language, Vision, Style, and IT (explained further in Sec.
-4.2).
-While language and vision hallucinations have been explored earlier, and methods to alleviate them have been proposed, we show for the first time that Style and IT hallucinations exist and existing methods fail to alleviate them.
-(2) Vision Hallucinations: Unlike language hallucinations, the tokens for vision hallucinations are generated by attending to the input image, and are shifted tokens.
-These are hallucinations caused when the model fails to accurately recognize specific components within an input image, or just inaccurate reasoning or IE, among other factors.
+(1) Language Hallucinations: These are hallucinations that are caused when LVLMs over-rely on the language priors (or prior tokens in the response) rather than the input image.
+We attribute all hallucinated tokens that are unshifted as language hallucinations.
+[…]
+[…]
+Figure 5: (Left) Frequency comparison of hallucination categories.
+(Right) Comparison for LLaVA-v1.5 with hallucination mitigation techniques.
+The top graph compares AMBER, and the bottom graph compares MMMU.
+(2) Vision Hallucinations: Unlike language hallucinations, the tokens for vision hallucinations are generated by attending to the input image and are shifted or marginal tokens.
+These are hallucinations caused when the model fails to accurately recognize visual elements in an input image or inaccurate reasoning or knowledge extraction, among other factors.
 (3) Style Hallucinations.
@@
 Open-source VLLMs are often trained on instruction-tuning data synthesized from proprietary closed-source models.
-Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset ghosh2024closer ; gudibande2024the (e.g., lengthy answers, a summary line at the end of the response, etc.).
-However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding ghosh2024closer .
-Figure 5: (Left) Frequency comparison of hallucination categories.
-(Right) Comparison for LLaVA-v1.5 with hallucination mitigation techniques.
-The top graph compares AMBER, and the bottom graph compares MMMU.
+Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset (Ghosh et al., 2024; Gudibande et al., 2024) (e.g., lengthy answers, a summary line at the end of the response, etc.).
+However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding (Ghosh et al., 2024).
 (4) Instruction Tuning (IT) Hallucinations.
 These are hallucinations caused by biases learned during the instruction tuning stage.
-LLMs are known to mimic the responses associated with the unfamiliar examples in the model’s instruction-tuning data (i.e., those unfamiliar to the pre-trained base model) ghosh2024closer .
-Formally, this can be defined as distribution shift.
+LLMs are known to mimic the responses associated with the unfamiliar examples in the model’s instruction-tuning data (i.e., those unfamiliar to the pre-trained base model) (Ghosh et al., 2024).
+Formally, this can be defined as a distribution shift.
 As visual instruction tuning is an extreme case of such a distribution shift, IT hallucinations are a major cause of hallucinations in LLMs.
@@
 Fig.
-5 (top;left) compares the frequency of different categories of hallucinations, across various LVLMs for AMBER.
+5 (top;left) compares the frequency of different categories of hallucinations across various LVLMs for AMBER.
 Fig.
@@
 5 (bottom) offers a similar comparison for MMMU.
-Our key findings include: (1) Model-based improvements and mitigation strategies have significantly reduced language and vision hallucinations, but progress in style and IT hallucinations remains limited.
+Our key findings include: (1) Model-based improvements and mitigation strategies reduce language and vision hallucinations, but progress in style and IT hallucinations remains limited.
 (2) Specific mitigation techniques are more effective for certain types of hallucinations.
-For instance, decoding-based methods like VCD and OPERA excel at reducing language hallucinations, whereas object-grounding approaches like Woodpecker are more effective for vision hallucinations.
-Decoding-based techniques benefit from low-confidence hallucinated tokens and uncertainty encoded in logits as shown in Fig.
-4.
-Further details can be found in Appendix H.1.
-(3) Existing methods fail to reduce hallucinations on non-VI benchmarks like MMMU.
-We attribute this to the algorithmic biases of these methods to alleviate VR hallucinations (for example, these algorithms are crafted for real-world scenes).
+For e.g., decoding-based methods like VCD and OPERA excel at reducing language hallucinations, whereas grounding approaches like Woodpecker are more effective for vision hallucinations.
+Decoding-based techniques benefit from low-confidence hallucinated tokens and uncertainty encoded in logits discussed next.
+More discussion is in Appendix K.1.
+(3) Existing methods fail to reduce hallucinations on reasoning benchmarks like MMMU.
+We attribute this to the algorithmic biases of these methods that are crafted to just reduce object hallucinations.
 {mdframed}
-[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: While community efforts have enhanced the visual recognition capabilities of LVLMs, substantial challenges persist in mitigating hallucinations for images that depict scenarios beyond real-world scenes.
-Additionally, hallucinations can be distinctly categorized based on their occurrence patterns, and some categories have received lesser attention.
-5 The Visual Perception Gap
-As illustrated in the previous section, current progress in LVLMs primarily alleviates VR hallucinations.
-In this section, we delve into LVLM hallucinations in various benchmarks, suggesting that the issue may stem from alignment problems.
-We first show that for cognitive prompts, LVLMs do not tend to pay attention to the input image but respond only using language priors.
-Next, we show that this problem may arise from a visual perception gap, i.e., although LVLMs may accurately recognize visual elements in the input image and possess sufficient knowledge and reasoning skills to provide a factual response, they still fail to respond accurately to the prompt and hallucinate.
-5.1 Assessing the Extent to Which LVLMs Attend to the Input Image
-Figure 6: Base Rank Comparison between AMBER and MATH-Vision datasets as a function of token position in responses generated by CogVLM.
-To assess the influence of the image on the LVLM response, we compare the Base Rank (defined in Section 4.2) of tokens across datasets.
-A lower value would signify that the model solely responds using language priors while a higher value would signify that the model pays attention to the image, i.e., it responds with tokens that are different from what the model would have responded with only language priors.
+[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Current hallucination mitigation techniques mainly improve LVLMs’ visual recognition skills but are less effective in improving other cognitive abilities, such as reasoning.
+Additionally, these improvements are largely limited to real-world scenes and specific types of hallucinations, leaving areas like non-real-world scenes largely unexplored.
+4 The Visual Perception Gap: LVLMs can see but not perceive
+It is evident that current hallucination mitigation techniques primarily enhance visual recognition but fail to improve other cognitive abilities, such as reasoning or knowledge extraction.
+In this section, we investigate the underlying cause of this limitation.
+Our analysis reveals that LVLMs often rely on language priors rather than attending to the input image when generating responses to reasoning prompts.
+This leads us to identify a critical issue: the visual perception gap.
+While LVLMs can accurately recognize visual elements and possess the necessary knowledge and reasoning skills to respond factually, they struggle to perceive and interpret these elements in relation to the input prompt.
+This gap in perception causes hallucinations, resulting in incorrect responses.
+4.1 Visual Blind Spots: LVLMs Ignore the Input Image for Reasoning
+[…]
+[trim= 0.5cm 1.5cm 0cm 1cm, width=]TR_line_cogvlm_first_token.pdf
+Figure 6: Base Rank Comparison between AMBER and MATH-Vision datasets as a function of token position in responses (for CogVLM).
+With findings from previous sections, we now assess the influence of the image on the LVLM response.
+Precisely, we compare the Base Rank (defined in Section 3.3) of tokens across datasets.
+A lower value would signify that the model solely responds using language priors, while a higher value would signify that the model pays attention to the image, i.e., it responds with tokens different from what the model would have responded with only language priors.
 Fig.
 6 plots the Base Rank of tokens against the token position in response (average across the dataset).
-As we see, the Base Rank for AMBER, which prompts for an image description, is higher than MMU, which requires it to reason or extract knowledge based on the image.
-This hints towards an alignment issue: different from AMBER, for MMMU LVLMs are unable to perceive the image before responding, under-relying on visual cues and over-relying on language priors for generating responses.
-In Appendix H.1 we explain why decoding-based mitigation techniques discussed in Section 4.2 do not work for cognitive prompts.
+As we see, the Base Rank for AMBER, which prompts for an image description, is higher than Math-Vision, which requires it to reason or extract knowledge based on the image.
+This hints towards an alignment issue: different from AMBER, for Math-Vision LVLMs are unable to perceive the image before responding, under-relying on visual cues and over-relying on language priors for generating responses.
+In Appendix K.1 we explain why decoding-based mitigation techniques discussed in Section 3.3 do not work for reaosning in cognitive prompts.
 4.2 Disentangling Visual Recognition from other Cognitive Skills
+A key objective of alignment fine-tuning is to equip LVLMs with implicit visual perception skills.
+As outlined in Section 3, LVLMs should interpret images in context and respond accurately to information-seeking or reasoning prompts.
+To determine if hallucinations arise from the base LLM’s limitations or alignment gaps, we aim to isolate visual recognition from other cognitive skills and evaluate them separately.
+We use GPT-4 to rephrase prompts from visual information-seeking and reasoning benchmarks, modifying them so they can be answered without the image (examples in Appendix K.7 and prompt used in Appendix G).
+[…]
+[trim= 0cm 0.5cm 0cm 0.5cm, width=]lm_radar_combined.pdf
 Figure 7: (Left) Performance comparison of different LVLMs when prompted w/ original prompt vs rephrased prompts w/o image (-t).
 (Right) Performance comparison of different LVLMs for their ability to generate a faithful image description.
-A fundamental goal of alignment fine-tuning is to enable an LVLM to master implicit visual perception.
-As discussed in Section 3, when faced with an information-seeking or reasoning prompt, an LVLM is expected to respond directly to the prompt and to have implicitly comprehended the accompanying image with respect to the prompt instruction.
-To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we aim to isolate visual recognition from other cognitive skills and evaluate them independently.
-To achieve this, we ask GPT-4 to rewrite the instructions in prompts from various visual information-seeking and reasoning benchmarks, such that they can be answered without the accompanying image.
-We show examples in Appendix H.6.
-Subsequently, we prompt LVLMs with these modified prompts without the image.
-Next, we prompt the LVLM to separately generate a description of the input image and evaluate this description for its faithfulness and detailedness (i.e., how well the image description can aid in a factual response to the original prompt).
-Evaluation prompts are provided in Appendix D.
-Fig.
-7 (left) compares the performance of text-only prompts (denoted with -t in the key) with their original image-text counterpart across various LVLMs and Fig.
-7 (right) shows their VR scores.
-As we see, in all cases, LVLMs fare well in VI, and are able to provide better factual responses than with the original prompt together with the input image.
-Thus, we conclude that while LVLMs demonstrate proficiency in visual recognition and have sufficient knowledge and reasoning abilities to generate factual responses, they encounter difficulties in employing these skills when responding to prompts that require their simultaneous application.
-This underscores a significant gap in alignment via lack of visual perception: LVLMs can recognize visual elements in images but fail to fully perceive them to accurately link what they recognize to their internal knowledge for reasoning or information extraction.
-We hypothesize a possible source of this issue to be the fact that most fine-tuning (or alignment) datasets for LVLMs consist of image-description pairs, which makes them struggle with accurate visual perception for non-description prompts.
+LVLMs are first evaluated on these text-only prompts and then on generating independent image descriptions (evaluation prompt in Appendix G).
+Figure 7 (left) compares the performance of LVLMs on the text-only prompts (marked with -t) against their original image-text versions, while Figure 7 (right) shows VR scores.
+Our results indicate that LVLMs excel in visual recognition and perform better on factual responses when using text-only prompts than with combined image-text inputs.
+This suggests that LVLMs have the necessary knowledge and reasoning skills but struggle to integrate these with visual perception in prompts requiring both.
+We identify this as a perception gap: LVLMs can recognize visual elements but fail to fully perceive them in a way that connects recognition to internal knowledge, which is essential for reasoning or information extraction.
+This challenge likely arises due to several factors: (i) training datasets that are largely composed of image-description pairs, (ii) alignment datasets that contain strong visual cues within the natural-language prompts.
 4.3 Uncertainty in the Logit Space
-Models Dataset k Variance Range Avg.
-LLaVA-v1 AMBER 4.19 0.05 0.31 0.23
-LLaVA-v1.5 AMBER 3.91 0.07 0.39 0.29
-CogVLM AMBER 3.39 0.09 0.46 0.21
 […]
-LLaVA-v1.5 MMMU 3.53 0.09 0.41 0.23
-CogVLM MMMU 3.16 0.19 0.47 0.25
-Table 1: Post-truncation probability statistics for various models using the elbow method. All values are computed across all hallucinated tokens in each dataset.
+[…]
+Table 1: Post-truncation probability statistics using the elbow method. All values are in the format: non-/hallucinated averaged across all tokens in the dataset.
 To investigate the origins of hallucinations, we analyze the logit space of hallucinated tokens.
 Specifically, after applying the softmax operation to the logits, we rank them by their probability scores.
-We then use the elbow method (detailed in Appendix H.5) to truncate the distribution, obtaining the top-k tokens along with their corresponding probabilities.
+We then use the elbow method (detailed in Appendix K.6) to truncate the distribution, obtaining the top-k tokens along with their corresponding probabilities.
 This truncation approach, based on the elbow method, effectively identifies the tokens most likely to be selected during multinomial sampling.
@@
 {mdframed}
-[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Our analysis of disentangling VR from other cognitive skills reveals that LVLMs generally demonstrate accurate VR and possess the necessary skills to respond accurately.
-However, they struggle to effectively link they recognize to their internal knowledge when responding directly, which inhibits the accurate execution of these skills.
+[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: LVLMs generally demonstrate accurate recognition and cognition skills to respond and reason accurately.
+However, they struggle to effectively link they recognize to their internal knowledge during the reasoning process, which results in inaccurate responses.
 This issue underlines what we identify as the visual perception gap.
 5 Visual Description Grounding Decoding (VDGD)
-Figure 8: Illustration of VDGD Hallucination Mitigation.
-To bridge the visual perception gap and reduce hallucinations, we propose Visual Description Grounding Decoding (VDGD), a simple, effective, and training-free hallucination mitigation technique.
-Specifically, during the decoding process, VDGD first adds the image description as a prefix to the text instruction of the prompt and then grounds token generation to the description of the input image by selecting a token that deviates the least from the description.
-We measure token deviation from the description with a novel formulation that employs Kullback–Leibler divergence at its core.
-The description is generated by the LVLM itself.
-VDGD operates on 2 core hypotheses: (1) Since LVLMs rely on language priors for cognitive prompts (Section 5.1), adding the image description as a prefix to the instruction can lead to improved visual perception by explicitly conditioning the response on the description.
-(2) Re-scoring the equally and low-confidence tokens (Section 1) based on deviation from the description can increase the confidence of the accurate token from several low-confidence tokens xu2023look .
+[…]
+[trim= 0cm 0.25cm 0cm 0.9cm, width=0.95]vllm-kl4.pdf
+Figure 8: Illustration of our proposed VDGD method.
+Main Motivation.
+To bridge the visual perception gap and reduce hallucinations for cognitive prompts, we propose Visual Description Grounding Decoding (VDGD), a simple, effective, and training-free hallucination mitigation technique.
+During decoding, VDGD first appends the image description (generated by the LVLM itself) as a prefix to the prompt’s text instruction, then grounds token generation by selecting the token that deviates the least from the description.
+Token deviation is measured using a novel formulation based on Kullback–Leibler divergence of the token with the description generated by the LVLM itself.
+Much like humans, who often write down their key observations in an image and refer to them while solving complex reasoning tasks, we hypothesize that grounding the LVLM’s responses in a detailed visual description can significantly enhance its reasoning process.
+Additionally, re-scoring the equally and low-confidence tokens (discussed in Section 1) based on deviation from the description can increase the confidence of the accurate token from several low-confidence tokens (Xu et al., 2023).
 5.1 Methodology
@@
 8 illustrates the VDGD methodology.
-For an input prompt, we first employ the LVLM to generate a description of the image accompanying the prompt (see Appendix D for the prompt).
+For an input prompt, we first employ the LVLM to generate a description of the image accompanying the prompt (see Appendix G for the prompt).
 Next, we concatenate the generated description as a prefix to the original prompt.
@@
 Here, $p_{VLM}(x_{i}\mid x_{<i})$ denotes the probability distribution for the next token, $x_{i}$, given the previous tokens.
-Inspired by our findings in Section 5.3, we first apply a plausibility constraint to truncate the vocabulary space of the next token $x_{i}$ li-etal-2023-contrastive .
+Inspired by our findings in Section 4.3, we first apply a plausibility constraint to truncate the vocabulary space of the next token $x_{i}$ (Li et al., 2023).
 Essentially, this step cuts off tokens from the vocabulary $\mathcal{V}$ below the elbow and keeps only the top-k plausible tokens.
 This is done as follows:
-$p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\begin{cases}p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right),&\text{ if }p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\geq\alpha\max_{w}p_{\operatorname{VLM}}\left(w\mid x_{<i}\right)\leq\alpha\\ -\infty,&\text{ Otherwise }\end{cases}$ (2)
+$p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\begin{cases}p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right),&\text{ if }p_{\text{VLM}}\left(x_{i}\mid x_{<i}\right)\geq\alpha\max_{w}p_{\operatorname{VLM}}\left(w\mid x_{<i}\right)\\ -\infty,&\text{ Otherwise }\end{cases}$ (2)
 where $\alpha$ is a hyper-parameter between [0,1].
@@
 With this distribution, we can now apply any sampling technique to sample the next token.
-6.2 VaLLu Benchmark
-Figure 9: Distribution of task types in VaLLu.
+[…]
+!
+Benchmark Baseline LLaVA-v1 LLaVA-1.5 LLaVA-1.6 mPLUG-Owl2 InternLM-X CogVLM CogVLM2 Qwen2-VL MMMU Vanilla-greedy 1.26 1.35 1.42 1.40 2.01 1.66 2.08 2.39 Vanilla-sampling 1.27 1.44 1.40 1.41 2.05 1.64 2.10 2.35 VDD 1.34 1.52 1.44 1.53 2.22 1.68 2.14 2.64 OPERA 1.30 1.43 1.57 1.64 2.25 1.62 2.21 2.44 Woodpecker 1.32 1.44 1.63 1.61 2.12 1.65 2.14 2.56 LRV 1.29 1.49 1.61 1.58 2.08 1.59 2.18 2.48 LURE 1.31 1.47 1.60 1.64 2.27 1.66 2.27 2.71 PAI 1.42 1.39 1.44 1.56 2.23 1.65 2.29 2.64 HALC 1.40 1.54 1.63 1.65 2.15 1.67 2.24 2.69 VDGD (ours) 1.49 (+18%) 1.62 (+20%) 1.75 (+23%) 1.72 (+23%) 2.39 (+19%) 1.71 (+3%) 2.47 (+19%) 2.91 (+22%) MathVista Vanilla-greedy 1.56 1.65 2.00 1.92 2.56 2.15 2.45 2.97 Vanilla-sampling 1.54 1.68 1.91 1.94 2.52 2.19 2.46 2.95 VDD 1.67 1.68 2.07 2.11 2.59 2.53 2.81 3.19 OPERA 1.64 1.83 2.04 2.04 2.62 2.30 2.53 3.22 Woodpecker 1.72 2.10 2.19 2.13 2.43 2.43 2.59 3.13 LRV 1.68 1.68 2.17 1.99 2.61 2.20 2.62 3.30 LURE 1.59 2.03 2.21 2.07 2.37 2.45 2.49 3.13 PAI 1.68 1.85 1.97 2.08 2.63 2.20 2.60 3.39 HALC 1.71 2.05 2.17 1.97 2.54 2.24 2.77 3.38 VDGD (ours) 1.84 (+18%) 2.19 (+33%) 2.44 (+22%) 2.24 (+17%) 2.88 (+13%) 2.62 (+22%) 2.93 (+20%) 3.59 (+21%) MMBench Vanilla-greedy 2.67 2.89 3.11 2.98 3.58 3.19 3.71 3.93 Vanilla-sampling 2.69 2.92 3.12 3.02 3.59 3.22 3.68 3.95 VDD 2.80 3.03 3.09 3.22 3.93 3.28 3.81 4.20 OPERA 2.81 2.94 3.17 3.10 3.95 3.22 3.93 4.05 Woodpecker 2.83 2.96 3.14 3.07 3.84 3.33 3.83 4.19 LRV 2.77 2.99 3.13 3.23 3.84 3.25 3.77 4.05 LURE 2.72 2.97 3.13 3.26 3.72 3.21 3.82 4.16 PAI 2.75 3.00 3.04 3.16 3.94 3.35 3.97 4.05 HALC 2.73 2.93 3.02 3.18 3.81 3.27 4.04 4.08 VDGD (ours) 2.93 (+10%) 3.12 (+8%) 3.26 (+5%) 3.32 (+11%) 4.08 (+14%) 3.42 (+7%) 4.15 (+12%) 4.31 (+10%) MME Vanilla-greedy 3.32 3.54 3.65 3.49 3.75 3.57 3.64 3.86 Vanilla-sampling 3.34 3.53 3.62 3.48 3.77 3.58 3.66 3.85 VDD 3.46 3.61 3.77 3.59 3.96 3.67 3.95 4.21 OPERA 3.42 3.59 3.82 3.54 3.93 3.60 3.88 4.09 Woodpecker 3.37 3.55 3.76 3.63 3.82 3.59 3.92 4.15 LRV 3.43 3.63 3.78 3.52 3.95 3.61 3.96 3.99 LURE 3.42 3.62 3.87 3.67 3.93 3.72 3.77 4.10 PAI 3.38 3.56 3.81 3.55 3.89 3.69 3.82 4.04 HALC 3.47 3.58 3.83 3.71 4.01 3.59 3.80 4.14 VDGD (ours) 3.59 (+8%) 3.70 (+5%) 3.99 (+9%) 3.82 (+9%) 4.12 (+10%) 3.79 (+6%) 4.09 (+12%) 4.34 (+12%) Oven Vanilla-greedy 2.44 2.66 3.13 2.86 3.35 3.35 3.48 3.64 Vanilla-sampling 2.45 2.65 3.10 2.84 3.33 3.37 3.45 3.61 VDD 2.20 2.44 2.96 2.67 3.10 3.21 3.39 3.59 OPERA 2.50 2.49 3.20 2.92 3.44 3.42 3.53 3.68 Woodpecker 2.46 2.46 3.21 2.90 3.45 3.45 3.52 3.70 LRV 2.51 2.52 3.18 2.88 3.47 3.40 3.57 3.74 LURE 2.49 2.51 3.18 2.91 3.42 3.46 3.55 3.72 PAI 2.54 2.60 3.26 2.95 3.48 3.48 3.60 3.78 HALC 2.52 2.57 3.24 2.94 3.50 3.45 3.61 3.80 VDGD (ours) 2.78 (+14%) 2.84 (+7%) 3.49 (+12%) 3.15 (+10%) 3.68 (+10%) 3.59 (+7%) 3.75 (+8%) 3.92 (+7%) LLaVA-Bench Vanilla-greedy 3.01 3.12 3.23 3.13 4.12 3.68 4.15 4.37 Vanilla-sampling 2.98 3.10 3.27 3.17 4.08 3.71 4.17 4.33 VDD 2.93 3.16 3.18 3.15 3.89 3.57 4.13 4.28 OPERA 3.08 3.18 3.30 3.22 4.16 3.75 4.22 4.40 Woodpecker 3.10 3.22 3.36 3.23 4.21 3.73 4.21 4.39 LRV 3.16 3.19 3.30 3.22 4.27 3.72 4.20 4.49 LURE 3.18 3.26 3.35 3.24 4.25 3.64 4.27 4.51 PAI 3.20 3.28 3.38 3.25 4.30 3.74 4.32 4.54 HALC 3.25 3.27 3.37 3.22 4.24 3.76 4.35 4.53 VDGD (ours) 3.36 (+12%) 3.47 (+11%) 3.52 (+9%) 3.38 (+8%) 4.45 (+8%) 3.96 (+8%) 4.50 (+8%) 4.78 (+9%) MMVET Vanilla-greedy 2.39 2.52 2.72 2.87 3.39 3.18 3.47 3.64 Vanilla-sampling 2.34 2.50 2.70 2.84 3.32 3.35 3.42 3.62 VDD 2.29 2.61 2.67 2.92 3.29 3.42 3.39 3.58 OPERA 2.48 2.67 2.77 3.01 3.37 3.39 3.52 3.69 Woodpecker 2.47 2.64 2.82 3.06 3.40 3.44 3.50 3.72 LRV 2.58 2.56 2.79 2.92 3.38 3.42 3.51 3.75 LURE 2.54 2.63 2.81 3.12 3.41 3.46 3.56 3.71 PAI 2.62 2.60 2.83 3.18 3.37 3.49 3.59 3.78 HALC 2.66 2.70 2.88 3.15 3.44 3.47 3.54 3.76 VDGD (ours) 2.85 (+19%) 3.01 (+19%) 3.23 (+19%) 3.45 (+20%) 3.78 (+12%) 3.57 (+12%) 3.74 (+8%) 3.99 (+10%) VaLLu Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82 3.23 3.45 Vanilla-sampling 1.86 2.01 1.64 2.18 2.70 2.83 3.21 3.40 VDD 1.47 1.55 1.80 1.80 2.32 2.63 3.19 3.42 OPERA 2.04 2.05 2.62 2.26 2.65 2.85 3.22 3.47 Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91 3.28 3.52 LRV 1.98 2.10 2.65 2.19 2.59 2.88 3.32 3.51 LURE 2.03 2.03 2.64 2.24 2.64 2.78 3.28 3.48 PAI 2.04 2.23 2.78 2.34 2.76 2.86 3.31 3.54 HALC 2.02 2.22 2.74 2.32 2.72 2.89 3.29 3.52 VDGD (ours) 2.16 (+11%) 2.64 (+30%) 3.16 (+20%) 2.72 (+24%) 3.45 (+30%) 3.01 (+7%) 3.48 (+8%) 3.67 (+6%)
+Table 2: Performance comparison of VDGD with various baselines. VDGD outperforms by 2%-33%.
+5.2 Experimental Setup
+Datasets and Evaluation Metrics.
+For evaluation, we employ a variety of standard benchmarks focused on reasoning and information-seeking tasks.
+These include LLaVA-Bench, MM-Vet (Yu et al., 2023), MMBench (Liu et al., 2023d), MME (Fu et al., 2023), MathVista (test-mini subset), MathVision, and MMMU (validation set).
+For each benchmark, we employ our proposed GPT evaluation technique described in Section 3.
+Beyond being more robust to the diverse formats of model generation, we also find GPT scores to have a higher correlation to expert-human evaluation ($\approx$0.97; Section D has more details with original benchmark evaluation metrics that has a correlation of $\approx$0.92).
+We are inspired by a wealth of prior studies in hallucination evaluation (Sicong Leng et al., 2023; Yu et al., 2023; Liu et al., 2023c).
+All results are averaged across 3 runs.
+Baselines.
+We compare VDGD against VCD, OPERA, Woodpecker, LRV, LURE, HALC (Chen et al., 2024a) and PAI (Liu et al., 2024d) hallucination mitigation techniques.
+Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla-greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7).
+We employ greedy decoding for all methods as we find no difference in performance on sampling.
+We generate image descriptions for VDGD using the VCD mitigation technique.
+For LVLMs, we employ LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, CogVLM, Qwen2-VL (Wang et al., 2024b) and CogVLM2 (Hong et al., 2024).
+Proposed VaLLu Benchmark.
 Automatic evaluation of LVLMs, with powerful closed-source models, on multiple large benchmarks can be prohibitively expensive.
 Thus, we propose VaLLu, a benchmark that we build by amalgamating existing benchmarks.
-VaLLu comprises a total of 1500 instances, with 554 sourced from Oven, 535 from MMMU, 218 from MMC, 74 from MathVista, 60 from HallusionBench, 53 from MATH-Vision, and 14 from MME fu2023mme .
+[…]
+[trim= 0cm 1cm 0cm 0cm, width=]chart-11.pdf
+Figure 9: Distribution of task types in VaLLu.
+VaLLu comprises a total of 1500 instances, with 554 sourced from Oven, 535 from MMMU, 218 from MMC, 74 from MathVista, 60 from HallusionBench, 53 from MATH-Vision, and 14 from MME.
 We only include instances with instructions that require open-ended generations and do not include any binary Yes/No or Multi-choice questions.
 We control the size to ensure the evaluation is affordable while keeping the diversity of tasks and topics for comprehensive analysis.
-VaLLu is manually filtered (we found various noisy examples in existing benchmarks that do not have a definite ground-truth response and thus make evaluation difficult, we discuss this further in Appendix H.3), annotated (we annotate each instance for the type of skill required by the LVLM to generate a response, we illustrate this in Fig.
-9) and answered (we accompany each instance with its ground-truth response).
+VaLLu is manually filtered for noisy examples commonly found in existing benchmarks (we discuss this further in Appendix K.3), annotated with additional meta-data (illustrated in Fig.
+9) and paired with an expert-provided response.
 5.3 Results and Analysis
-Experimental Setup and Baselines.
-We compare VDGD against VCD, OPERA, Woodpecker, LRV, and LURE mitigation techniques.
-Additionally, we compare two vanilla decoding methods (without any additional mitigation techniques): Vanilla-greedy with vanilla greedy decoding and Vanilla-sampling with multinomial sampling-based decoding (top-p=0.5 and temperature=0.7).
-We employ greedy decoding for all methods as we find no difference in performance on sampling.
-We generate image descriptions for VDGD with the VCD mitigation technique and ablate the same.
 Quantitative results.
-Table 2 compares the performance of VDGD with other baselines on the VaLLu Benchmark.
-We report factuality scores averaged across 3 runs for 3 random seeds.
+Table 9 compares the performance of VDGD with other baselines.
 As we clearly see, VDGD not only outperforms all baselines but is also the only method that alleviates and improves performance on VaLLu.
-Results on other full benchmarks, including MMU, MathVista, and MATH-Vision, are provided in Appendix I.5, together with scores for other aspects beyond factuality where we show that VDGD does not compromise the depth and expressivity of responses.
-Removing KL divergence or just adding the description prefix results in a drop in performance.
-Qualitative examples.Fig 10 compares responses from vanilla greedy decoding with VDGD. Grounding response generation results in more factual reasoning and information extraction.
-Figure 10: Illustration of responses generated with VDGD compared with greedy decoding. More in Sec. H.2.
-Methodology LLaVA-v1 LLaVA-v1.5 LLaVA-v1.6 mPLUG-Owl2 InternLM-X CogVLM
-Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82
-Vanilla-sampling 1.86 2.01 2.64 2.18 2.70 2.83
-VCD 1.47 1.55 1.80 1.80 2.32 2.63
-OPERA 2.04 2.05 2.62 2.26 2.65 2.85
-Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91
-LRV 2.03 2.10 2.64 2.24 2.64 2.78
-LURE 1.98 2.03 2.65 2.19 2.59 2.88
+Results on other full benchmarks are provided in Appendix L.6, together with scores for other aspects beyond factuality, where we show that VDGD does not compromise the depth and expressivity of responses.
+Ablation results in Appendix C demonstrate that removing KL divergence leads to drop in performance.
+Additionally, using only the description prefix without the image leads to performance drops, confirming that VDGD enhances reasoning through grounding rather than making the model rely completely on the description alone.
+Qualitative examples.
+Fig 10 compares responses from vanilla greedy decoding with VDGD.
+Grounding response generation results in more factual responses.
+Appendix B illustrates 2 failure cases of VDGD where incorrect descriptions can lead VDGD to hallucinate.
 […]
-VDGD (-) VCD $2.08_{\pm 0.09}$ $2.43_{\pm 0.15}$ $3.01_{\pm 0.08}$ $2.54_{\pm 0.07}$ $3.26_{\pm 0.12}$ $2.95_{\pm 0.06}$
-VDGD (-) KLD $2.05_{\pm 0.10}$ $2.30_{\pm 0.10}$ $2.78_{\pm 0.12}$ $2.37_{\pm 0.06}$ $3.01_{\pm 0.02}$ $2.87_{\pm 0.08}$
-Table 2: Comparison of VDGD with other hallucination mitigation techniques.
-All results are on the proposed VaLLu benchmark for the "Factuality" metric.
-Results on other benchmarks can be seen in Appendix I.5.
+[width=]vllm_sk.pdf
+Figure 10: Illustration of responses generated with VDGD compared with greedy decoding. More in Sec. K.2.
 6 Conclusion, Limitations and Future Work
@@
 Our results demonstrate that VDGD is the first method to effectively reduce hallucinations in responses that necessitate additional cognitive skills beyond just visual recognition.
-As part of future work, we would like to work on the current limitations of our work, which include: (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
+As part of future work, we would like to work on the current limitations of our work, which include (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
+7 Reproducibility Statement
+We provide our code here: https://anonymous.4open.science/r/VDGD-1E04/.
+All codes will be open-sourced upon paper acceptance, including all checkpoints.
+All experimental details, including training parameters and hyper-parameters, are provided in Section 5.3.
+The Appendix has more comprehensive details about our experimental setup.
 Appendix A Appendix
@@
   1. 1.
-    B: We provide a detailed explanation of computational complexity of VDGD algorithm described in section 6.
+    B: We illustrate failure cases for VDGD.
   2. 2.
-    C: We describe the impact of the paper on society and research communities.
+    C: We provide results for ablation of VDGDs key components.
   3. 3.
-    D: We provide the prompts used for object detection, description, rephrasing, noise removal and evaluations required for our experiments.
+    D: We provide correlation between GPT scores and actual metrics.
   4. 4.
-    E: We provide a detailed description of all the datasets utilized in this paper.
+    E: We provide a detailed explanation of computational complexity of VDGD algorithm described in section 5.
   5. 5.
-    F: We provide a detailed description of all the models we have experimented with.
+    F: We describe the impact of the paper on society and research communities.
   6. 6.
-    G: We provide a detailed description of all the prior-art hallucination mitigation techniques used as baselines.
+    G: We provide the prompts used for object detection, description, rephrasing, noise removal and evaluations required for our experiments.
   7. 7.
-    H: We provide details about:
-    • H.1: We provide detailed explanation about Logit Space for VR and Cognitive Prompts.
-    • H.2: We provide examples comparing generations from Greedy, Decoding, VCD and VDGD.
-    • H.3: We provide examples of noisy samples in datasets.
-    • H.4: We provide and in-depth explanation of all the details of Algorithm 1.
-    • H.5: We explain the elbow method used for logit analysis.
-    • H.6: We provide examples of rephrased prompts which do not require image input.
-    • H.7: We provide details of computational resources used for this paper.
+    H: We provide a detailed description of all the datasets utilized in this paper.
   8. 8.
-    I: We provide additional results for all the experiments described in the paper.
-Appendix B VDGD Computational Complexity
-Given a LVLM with a vocabulary size $V$, an input instruction $I$ with a description of length $M$ tokens, at each decoding step the complexity for equation (2) is $O(V)$ and let the number of plausible tokens after this step be $K$, the complexity for equation (3) is $O(VMK)$.
+    I: We provide a detailed description of all the models we have experimented with.
+  9. 9.
+    J: We provide a detailed description of all the prior-art hallucination mitigation techniques used as baselines.
+  10. 10.
+    K: We provide details about:
+    • K.1: We provide detailed explanation about Logit Space for VR and Cognitive Prompts.
+    • K.2: We provide examples comparing generations from Greedy, Decoding, VCD and VDGD.
+    • K.3: We provide examples of noisy samples in datasets.
+    • K.4: We provide and in-depth explanation of all the details of Algorithm 1.
+    • K.6: We explain the elbow method used for logit analysis.
+    • K.7: We provide examples of rephrased prompts which do not require image input.
+    • K.8: We provide details of computational resources used for this paper.
+  11. 11.
+    L: We provide additional results for all the experiments described in the paper.
+Appendix B VDGD Failure Cases
+Fig.
+11 illustrates 2 failure cases of VDGD.
+We show that VDGD can lead to incorrect responses due to an incorrectly generated image description.
+However, we also argue that as LVLMs get better at describing images, the performance of VDGD decoding will improve.
+[…]
+[width=]failure_cases.pdf
+Figure 11: Illustration of failure cases of VDGD.
+Appendix C VDGD Ablation Results
+To demonstrate the effectiveness of VDGD, we conduct ablation studies on its key components: (i) VDGD (- VCD): Image descriptions are generated without using Visual Contrastive Decoding (VCD), (ii) VDGD (- KLD): KLD-based sampling is removed, and only the description is appended, (iii) VDGD (- Image): VDGD is applied without the original image input; only the image description is appended to the prompt.
+Results are presented on the VDGD benchmark in Table 3.
+Removing KL Divergence leads to a substantial performance drop.
+We notice a similar phenomenon by removing the input image with the steepest drop.
+This shows that the input image is still essential for the model.
+[…]
+[…]
+Table 3: Ablation results for VDGD. VDGD sees a drop in performance when the KLD decoding objective is removed and a major drop when no input image is provided.
+Appendix D Correlation between GPT scores and actual metrics
+Background of Experts in Human Evaluation.
+For the expert human evaluation of model generations, we engaged a group of 5 PhD students with at least four years of research experience in computer vision.
+These experts were tasked with evaluating the outputs of the models for accuracy, relevance, and factual correctness.
+To ensure thorough evaluation, they were provided with internet access to verify factual information when necessary.
+The evaluation process was conducted in an anonymized format to maintain objectivity, with the experts working independently and without prior knowledge of the model specifics.
+We selected 500 samples or the size of the benchmark, whichever was lower.
+Appendix E VDGD Computational Complexity
+Given an LVLM with a vocabulary size $V$, an input instruction $I$ with a description of length $M$ tokens, at each decoding step, the complexity for equation (2) is $O(V)$ and let the number of plausible tokens after this step be $K$, the complexity for equation (3) is $O(VMK)$.
 The total computational complexity of VDGD at each decoding step is $O(V)+O(VMK)\approx O(VMK)$.
-Appendix C Broader Impact
+Appendix F Broader Impact
 This paper addresses the broader impact of reducing hallucinations in Large Vision-Language Models (LVLMs), which are increasingly utilized in diverse applications such as automated content generation, assistive technologies, and decision-making systems.
@@
 The proposed Visual Description Grounded Decoding (VDGD) method, being training-free, offers an accessible and scalable solution that can be readily implemented across various models to mitigate risks associated with inaccurate or misleading outputs.
-Furthermore, the introduction of the VaLLu benchmark contributes to the field by providing a tool for comprehensive evaluation of cognitive capabilities in LVLMs, promoting transparency and fostering further research in the area.
+Furthermore, the introduction of the VaLLu benchmark contributes to the field by providing a tool for the comprehensive evaluation of cognitive capabilities in LVLMs, promoting transparency and fostering further research in the area.
 Collectively, these contributions aim to improve user trust in AI systems and encourage responsible AI development that prioritizes factual accuracy and understanding.
-Appendix D Prompts
-For image description benchmarks we employ the prompt illustrated in Fig.
-11.
-For non-image description benchmarks, we modify this prompt slightly (we modify the part of the prompt for extracting hallucinated phrases) and we illustrate the same in Fig.
+Appendix G Prompts
+For image description benchmarks, we employ the prompt illustrated in Fig.
+12.
+For non-image description benchmarks, we modify this prompt slightly (we modify the part of the prompt for extracting hallucinated phrases), and we illustrate the same in Fig.
+16.
+The prompt used to generate image descriptions for VDGD is illustrated in Fig.
 15.
-The prompt used to generate image descriptions for VDGD is illustrated in Fig.
-14.
-Correctness in the prompts refers to Factuality in all tables in Section I.
+Correctness in the prompts refers to Factuality in all tables in Section L.
 We name it so after an ablation study on the terminology that works better for prompting the judge (GPT-4).
 All prompt evaluations on GPT-4 are done with a temperature of 0.7.
+[…]
+[width=]evalutaion_prompt.pdf
 Figure 12: Prompt used to evaluate image description benchmarks using GPT-4-Turbo as described in Section 3.2.
+[…]
+[width=]rephrase_prompt.pdf
 Figure 13: Prompt used to evaluate the performance of LLMs on rephrased instructions using GPT-4-Turbo as described in Section 4.2.
+[…]
+[width=]llama3_prompt.pdf
 Figure 14: Prompt used to identify the visual elements in images using Llama3 for the algorithm described in Section 3.3.
+[…]
+[width=]image_desc_prompt.pdf
 Figure 15: Prompt used to describe images using LVLMs for VDGD as described in Section 5.
-Figure 15: Prompt used to evaluate generations from benchmarks like VaLLu, MMMU, MathVista, and SynthDoG using GPT-4-Turbo for the results in Table 2.
+[…]
+[width=]maths_eval_prompt.pdf
+Figure 16: Prompt used to evaluate generations from cognitive benchmarks in Table 9.
+[…]
+[width=]prompt_rebut.pdf
+Figure 17: GPT evaluation prompt for evaluating informativeness of image descriptions generated by LVLMs.
+[…]
+[width=]text-only-eval.pdf
 Figure 18: Prompt used to evaluate generations for rephrased instructions on benchmarks like VaLLu, MMMU, MathVista, and SynthDoG using GPT-4-Turbo as described in Section 4.2.
-Appendix E Dataset Details
+Appendix H Dataset Details
 VaLLu We propose VaLLu benchmark which is sourced from Oven, MMMU, MMC, MathVista, HallusionBench, MATH-Vision and MME.
 This dataset is licensed under all the licenses of the original benchmarks that it was sourced from.
-AMBER AMBER [29] is a benchmark for evaluating hallucinations in both the generative task and discriminative task of MLLMs.
+AMBER AMBER (Wang et al., 2023a) is a benchmark for evaluating hallucinations in both the generative task and discriminative task of MLLMs.
 It provides comprehensive coverage of evaluations for various types of hallucination, including existence, attribute, and relation.
 Licensed under Apache-2.0 License.
-MMMU MMMU [39] benchmark evaluates multimodal models on multi-discipline tasks and demands college-level subject knowledge and deliberate reasoning.
+MMMU MMMU (Yue et al., 2024) benchmark evaluates multimodal models on multi-discipline tasks and demands college-level subject knowledge and deliberate reasoning.
 It consists of 11.5k questions, spanning 30 subjects and 183 subfields, focusing on advanced perception and reasoning on domain-specific knowledge.
 Licensed under Apache-2.0 License.
-MMC MMC [18] is a human-annotated benchmark comprising 9 distinct tasks to evaluate LVLM’s reasoning capabilities to comprehend visual charts.
+MMC MMC (Liu et al., 2023a) is a human-annotated benchmark comprising 9 distinct tasks to evaluate LVLM’s reasoning capabilities to comprehend visual charts.
 It comprises tasks such as chart information extraction, chart reasoning, contextual chart understanding, etc.
@@
 Licensed for research purposes.
-MathVista MathVista [23] is a robust mathematical reasoning evaluation benchmark that consists of challenging tasks that require fine-grained, deep visual recognition and compositional reasoning.
+MathVista MathVista (Lu et al., 2023) is a robust mathematical reasoning evaluation benchmark that consists of challenging tasks that require fine-grained, deep visual recognition and compositional reasoning.
 It consists of 6141 examples derived from 31 multimodal datasets involving mathematics.
 Licensed under CC-BY-SA-4.0.
-MATH-Vision MATH-Vision [30] dataset is a collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions.
-The dataset spans 16 distinct mathematical disciplines and graded across 5 levels of difficulty to evaluate the mathematical reasoning abilities of Multimodal Large Language Models(MLLM).
+MATH-Vision MATH-Vision (Wang et al., 2024a) dataset is a collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions.
+The dataset spans 16 distinct mathematical disciplines and is graded across 5 levels of difficulty to evaluate the mathematical reasoning abilities of Multimodal Large Language Models(MLLM).
 Licensed under MIT License.
-HallusionBench HallusionBench [9] is a comprehensive benchmark designed for the evaluation of image-context reasoning.
+HallusionBench HallusionBench (Guan et al., 2023) is a comprehensive benchmark designed for the evaluation of image-context reasoning.
 The benchmark presents significant challenges to advanced large visual-language models (LVLMs), such as GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA- 1.5, by emphasizing nuanced understanding and interpretation of visual data.
@@
 Licensed under BSD-3-Clause.
-MME Multimodal large language model Evaluation Benchmark (MME) [6] measures perception and cognition abilities on a total of 2 tasks(Perception and Congition) and 14 subtasks(Existence, Count, Position, Color, Poster, Celebrity, Scene, Landmark, Artwork, OCR, Commonsense Reasoning, Numerical Calculation, Text Translation and Code Reasoning).
+LlavaBench LlavaBench (Liu et al., 2023c) is a challenging evaluation suite designed to assess visual-language alignment and instruction-following capabilities of large visual-language models.
+It consists of two benchmarks: LLaVA-Bench (COCO), which includes 30 randomly selected images from COCO-Val-2014 paired with 90 questions covering conversational, descriptive, and complex reasoning tasks, and LLaVA-Bench (In-the-Wild), featuring 24 diverse images from various domains, including memes, paintings, and sketches, with 60 questions.
+MM-Vet MM-Vet (Yu et al., 2023) is an evaluation benchmark designed to assess large multimodal models on complex multimodal tasks.
+It defines six core vision-language capabilities—Recognition, Knowledge, OCR, Spatial Awareness, Language Generation, and Math—and examines their integration across 16 emergent tasks.
+MM-Vet employs an LLM-based evaluator for open-ended outputs, providing unified scoring across diverse question types and answer styles.
+The benchmark offers insights into the capabilities of different large multimodal model paradigms and models beyond simple performance ranking.
+MME Multimodal large language model Evaluation Benchmark (MME) (Fu et al., 2023) measures perception and cognition abilities on a total of 2 tasks(Perception and Congition) and 14 subtasks(Existence, Count, Position, Color, Poster, Celebrity, Scene, Landmark, Artwork, OCR, Commonsense Reasoning, Numerical Calculation, Text Translation and Code Reasoning).
 To prevent data leakage from use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed.
-Oven Open-domain Visual Entity Recognition (OVEN) [12] is a task where a model needs to link an image onto a Wikipedia entity with respect to a text query.
+Oven Open-domain Visual Entity Recognition (OVEN) (Hu et al., 2023) is a task where a model needs to link an image onto a Wikipedia entity with respect to a text query.
 OVEN challenges models to select among six million possible Wikipedia entities, making it a general visual recognition benchmark with the largest number of labels.
 Licensed under MIT License.
-SynthDoG SynthDoG [14] is created by generating synthetic data which consists of various components: background, document, text, and layout.
+SynthDoG SynthDoG (Kim et al., 2022) is created by generating synthetic data which consists of various components: background, document, text, and layout.
 The background image is sampled from ImageNet, and the document is sampled from the collected paper photos.
@@
 Licensed under MIT License.
-Appendix F Model Details
+Appendix I Model Details
 LLaVa-v1.
-LLaVa-v1 [22] [^1] https://huggingface.co/liuhaotian/llava-llama-2-7b-chat-lightning-lora-preview is an open-source LVLM trained by fine-tuning LLaMA/Vicuna (13B) on GPT-generated multimodal instruction-following data.
+LLaVa-v1 (Liu et al., 2023c) [^3] https://huggingface.co/liuhaotian/llava-llama-2-7b-chat-lightning-lora-preview is an open-source LVLM trained by fine-tuning LLaMA/Vicuna (13B) on GPT-generated multimodal instruction-following data.
 It is initially trained on 558K filtered image-text pairs from LAION/CC/SBU and then, finetuned on 80K GPT-generated multimodal instruction-following data.
@@
 LLaVa-1.5.
-LLaVa-1.5 [20] [^2] https://huggingface.co/liuhaotian/llava-v1.5-7b builds on LLaVA-v1’s architecture by replacing linear projection design with a two-layer MLP.
+LLaVa-1.5 (Liu et al., 2023b) [^4] https://huggingface.co/liuhaotian/llava-v1.5-7b builds on LLaVA-v1’s architecture by replacing linear projection design with a two-layer MLP.
 It also includes additional academic-task-oriented VQA datasets for VQA, OCR, and region-level perception, to enhance the model’s multimodal capabilities.
@@
 LLaVa-1.6.
-LLaVa-1.6 [21] [^3] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b re-uses the pretrained connector of LLaVA-1.5 and is instruction tuned on 158K GPT-generated multimodal data, 500K academic-task-oriented VQA data, 50K GPT-4V data and 40K ShareGPT data.
+LLaVa-1.6 (Liu et al., 2024c) [^5] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b re-uses the pretrained connector of LLaVA-1.5 and is instruction tuned on 158K GPT-generated multimodal data, 500K academic-task-oriented VQA data, 50K GPT-4V data and 40K ShareGPT data.
 LLaVa-1.6 has improved reasoning, OCR, and world knowledge.
@@
 mPLUG-Owl2.
-mPLUG-Owl2 [36] [^4] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b is the first Multi-modal Large Language Model that performs well in both pure-text and multi-modal scenarios.
+mPLUG-Owl2 (Ye et al., 2023) [^6] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b is the first Multi-modal Large Language Model that performs well in both pure-text and multi-modal scenarios.
 mPLUG-Owl2 is pretrained 400 million image-text pairs from: Conceptual Captions (CC3M/CC12M), COCO, Laio-en, COYO, DataComp and for instruction tuning 5 types of data are used: image captioning, mage question answering, region-aware QA, multi-modal instruct data and text-only instruct.
@@
 InternLM-X.
-InternLM-X [40] [^5] https://huggingface.co/internlm/internlm-xcomposer2-vl-7b is a vision-language large model based on InternLM2 [3] for advanced text-image comprehension and composition.
+InternLM-X (Zhang et al., 2023) [^7] https://huggingface.co/internlm/internlm-xcomposer2-vl-7b is a vision-language large model based on InternLM2 (Cai et al., 2024) for advanced text-image comprehension and composition.
 InternLM-X is pre-trained on 1.1 billion images alongside 77.7 billion text tokens, including both public datasets and in-house concept data followed by supervised fine-tuning using multi-task training and instruction tuning.
@@
 CogVLM.
-CogVLM [31] [^6] https://huggingface.co/THUDM/cogvlm-chat-hf is a powerful open-source visual language model.
+CogVLM (Wang et al., 2023b) [^8] https://huggingface.co/THUDM/cogvlm-chat-hf is a powerful open-source visual language model.
 CogVLM-17B has 10 billion vision parameters and 7 billion language parameters.
@@
 Licensed under Apache-2.0 License.
-Llama-2-7b Llama-2-7b [28] [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf is an open source fine-tuned generative text model with 7 billion parameters.
+Llama-2-7b.
+Llama-2-7b (Touvron et al., 2023) [^9] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf is an open source fine-tuned generative text model with 7 billion parameters.
 It is an auto-regressive language model that uses an optimized transformer architecture.
@@
 Licensed under Llama 2 community license agreement.
-Vicuna-7b-v1.5 Vicuna-7b-v1.5 [45] [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5 is a 7b parameter open-sourced model and fine-tuned from Llama 2 with supervised instruction fine-tuning and is an auto-regressive language model based on the transformer architecture.
+Vicuna-7b-v1.5.
+Vicuna-7b-v1.5 (Zheng et al., 2023) [^10] https://huggingface.co/lmsys/vicuna-7b-v1.5 is a 7b parameter open-sourced model and fine-tuned from Llama 2 with supervised instruction fine-tuning and is an auto-regressive language model based on the transformer architecture.
 The training data is around 125K conversations from ShareGPT.
 Licensed under Apache-2.0 License.
-InternLM2-7B InternLM-7B [3] [^9] https://huggingface.co/internlm/internlm2-chat-7b InternLM2 is a open-source 7 billion parameter chat model with a 200k context window and achieves good performance in the“Needle-in-a-Haystack” test(on long-context tasks like LongBench and L-Eval) and is trained using Conditional Online RLHF (COOL RLHF).
+InternLM2-7B.
+InternLM-7B (Cai et al., 2024) [^11] https://huggingface.co/internlm/internlm2-chat-7b InternLM2 is a open-source 7 billion parameter chat model with a 200k context window and achieves good performance in the“Needle-in-a-Haystack” test(on long-context tasks like LongBench and L-Eval) and is trained using Conditional Online RLHF (COOL RLHF).
 Licensed under Apache-2.0 License.
-Appendix G Baseline Details
+CogVLM2.
+CogVLM2 (Hong et al., 2024) [^12] https://huggingface.co/THUDM/cogvlm2-llama3-chat-19B CogVLM2 is a new generation of visual language models based on Meta-LLaMA-3-8B-Instruct, offering significant improvements over previous versions.
+It supports 8K content length, image resolutions up to 1344x1344, and excels in both English and Chinese.
+CogVLM2 has achieved state-of-the-art performance on benchmarks like TextVQA, DocVQA, and MM-Vet.
+The open-source models, including CogVLM2 and CogVLM2-Video, are available on GitHub, contributing to advancements in image and video understanding.
+Licensed under Apache-2.0 License.
+Qwen2-VL.
+Qwen2-VL (Wang et al., 2024b) [^13] https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct Qwen2-VL is the latest iteration of the Qwen-VL model, featuring state-of-the-art performance in image and video understanding tasks.
+It excels in processing images of various resolutions and aspect ratios and supports long-form video understanding for tasks like question answering and content creation.
+Key architectural updates include Naive Dynamic Resolution and Multimodal Rotary Position Embedding (M-ROPE), improving its multimodal processing capabilities.
+Available in 2, 7, and 72 billion parameter variants.
+Licensed under Apache-2.0 License.
+Appendix J Baseline Details
 VCD.
-Visual Contrastive Decoding (VCD) [26] is a training-free method that contrasts output distributions of original/unaltered inputs and distorted visual inputs.
+Visual Contrastive Decoding (VCD) (Sicong Leng et al., 2023) is a training-free method that contrasts output distributions of original/unaltered inputs and distorted visual inputs.
 VCD hypothesizes statistical bias and unimodal priors as the two essential causes of object hallucinations and reduces the dependence on them.
@@
 OPERA.
-Over-trust Penalty and a Retrospection-Allocation (OPERA) [13] is a training-free approach that hypothesizes that hallucinations are closely tied to the knowledge aggregation patterns in the self-attention matrix and that generation of new tokens by models is majorly dependent on a few summary tokens(over-trust issue) but not all of the previous tokens.
+Over-trust Penalty and a Retrospection-Allocation (OPERA) (Huang et al., 2023b) is a training-free approach that hypothesizes that hallucinations are closely tied to the knowledge aggregation patterns in the self-attention matrix and that generation of new tokens by models is majorly dependent on a few summary tokens(over-trust issue) but not all of the previous tokens.
 This methodology introduces a penalty term on the model logits during beam-search decoding to mitigate the over-trust issue and implements a rollback strategy to retrospect the presence of summary tokens.
@@
 Woodpecker.
-Woodpecker [38] is a training-free method that identifies and corrects hallucinations in the generated text.
+Woodpecker (Yin et al., 2023b) is a training-free method that identifies and corrects hallucinations in the generated text.
 Unlike other training-free methodologies, Woodpecker works after the model generation process.
@@
 LRV.
-LRV [17], or Large-scale Robust Visual Instruction tuning, proposes a dataset of visual instructions generated by GPT4, covering 16 vision-and-language tasks with open-ended instructions and answers.
+LRV (Liu et al., 2024a), or Large-scale Robust Visual Instruction tuning, proposes a dataset of visual instructions generated by GPT4, covering 16 vision-and-language tasks with open-ended instructions and answers.
 Each instance has positive instruction samples and negative instructions for more robust visual instruction tuning.
@@
 LURE.
-LVLM Hallucination Revisor (LURE) [47] attributes object hallucination to three factors: co-occurrence, uncertainty, and object position, and develops an object hallucination revisor which converts potential hallucinations into accurate outputs.
+LVLM Hallucination Revisor (LURE) (Zhou et al., 2023b) attributes object hallucination to three factors: co-occurrence, uncertainty, and object position, and develops an object hallucination revisor which converts potential hallucinations into accurate outputs.
 This method first creates a hallucination dataset by making modifications to the original data and trains the object hallucination revisor on this dataset.
 This method post-hoc rectifies object hallucinations in LVLMs by reconstructing less hallucinatory descriptions.
-Appendix H Additional Details
-H.1 Logit Space for VR and Cognitive Prompts
+PAI.
+Pay Attention to Image (PAI) (Liu et al., 2024d) is a training-free algorithm designed to reduce hallucinations in LVLMs.
+PAI intervenes during the inference process to make it more image-centric by amplifying attention to image tokens in the self-attention layers of LVLMs.
+By adjusting attention weights and subtracting logits of text-only inputs from multi-modal inputs, PAI mitigates the issue of “text inertia,” where outputs are overly influenced by context text.
+This method enhances image representation during generation, leading to more accurate, visually grounded outputs without requiring additional training or external tools.
+HALC.
+HALC (Chen et al., 2024a) is a novel decoding algorithm designed to mitigate object hallucinations (OH) in LVLMs.
+It integrates a robust auto-focal grounding mechanism to correct hallucinated tokens locally and a specialized beam search algorithm to reduce OH globally while preserving text generation quality.
+HALC can be easily integrated into any LVLM without additional training, addressing all three types of OH—existence, attribute, and relationship.
+Extensive experiments show that HALC outperforms state-of-the-art methods across multiple benchmarks.
+Appendix K Additional Details
+K.1 Logit Space for VR and Cognitive Prompts
 From our in-depth analysis of the logit space, we see that language hallucinations in LVLMs have a unique characteristic – the probability of the most confident token in the vocabulary marginally decreased from the probability of the same token in the base LLM when prompted with the same prior context (the most confident token by the LVLM will also be the most confident token by the language-only base LLM as the token on a language hallucination).
 Thus, decoding-based mitigation methods like OPERA and VDD benefit from this unique property, and contrasting tokens help mitigate hallucinations.
-H.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
-In this subsection, we illustrate several instances (figures 17, 20, 21, 18, 19) from VaLLu and compare their responses for LLaVa-1.5 with Greedy, VCD, and VDGD Decoding.
+K.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
+In this subsection, we illustrate several instances (figures 20, 22, 23, 21, 24) from VaLLu and compare their responses for LLaVa-1.5 with Greedy, VCD, and VDGD Decoding.
+[…]
+[width=]qual_with_desc.pdf
+Figure 19: Qualitative example for VaLLu benchmark also showing the descriptions.
+[…]
+[width=]qual_app_img_1.pdf
 Figure 20: Qualitative example 1 for VaLLu benchmark.
+[…]
+[width=]qual_app_img_4.pdf
 Figure 21: Qualitative example 2 for VaLLu benchmark.
+[…]
+[width=]qual_app_img_2.pdf
 Figure 22: Qualitative example 3 for VaLLu benchmark.
+[…]
+[width=]qual_app_img_3.pdf
 Figure 23: Qualitative example 4 for VaLLu benchmark.
+[…]
+[width=]qual_app_img_5.pdf
 Figure 24: Qualitative example 5 for VaLLu benchmark.
-H.3 Noisy Examples in Existing Datasets
-To maintain a high quality of the benchmark and eliminate noisy examples we performed a manual review of the existing datasets.
+K.3 Noisy Examples in Existing Datasets
+To maintain a high quality of the benchmark and eliminate noisy examples, we performed a manual review of the existing datasets.
 This was crucial to ensure that the data used in our paper was reliable.
@@
 This filtering was done by the paper authors and has been approved by our institution’s Institutional Review Board (IRB).
-We share a few examples (figures: 22, 23, 24, 25, 26, 27, 28, 29) of noise in the existing datasets below:
+We share a few examples (figures: 25, 26, 27, 28, 29, 30, 31, 32) of noise in the existing datasets below:
+[…]
+[width=]mmmu_noisy_img_1.pdf
 Figure 25: Noisy example 1 for MMMU benchmark.
+[…]
+[width=]mmmu_noisy_img_2.pdf
 Figure 26: Noisy example 2 for MMMU benchmark.
+[…]
+[width=]mmc_noisy_img_1.pdf
 Figure 27: Noisy example 1 for MMC benchmark.
+[…]
+[width=]mmc_noisy_img_2.pdf
 Figure 28: Noisy example 2 for MMC benchmark.
+[…]
+[width=]mathvista_noisy_img_1.pdf
 Figure 29: Noisy example 1 for MathVista benchmark.
+[…]
+[width=]mathvista_noisy_img_2.pdf
 Figure 30: Noisy example 2 for MathVista benchmark.
+[…]
+[width=]hallu_noisy_img_1.pdf
 Figure 31: Noisy example 1 for HallusionBench benchmark.
+[…]
+[width=]hallu_noisy_img_2.pdf
 Figure 32: Noisy example 1 for HallusionBench benchmark.
-H.4 Additional Details on Algorithm 1
+K.4 Additional Details on Algorithm 1
 In this Section, we describe our algorithm verbally.
@@
   2. 2.
-    Next, if the Base Rank is more than 0 and the hallucination can be found in the top-k retrieved elements from the IT dataset the model was aligned using, we attribute it to be an IT hallucination [8].
-To retrieve top-k image-instruction pairs that are closest to the prompt image, we employ CLIP [25] similarity.
+    Next, if the Base Rank is more than 0 and the hallucination can be found in the top-k retrieved elements from the IT dataset the model was aligned using, we attribute it to be an IT hallucination (Ghosh et al., 2024).
+To retrieve top-k image-instruction pairs that are closest to the prompt image, we employ CLIP (Radford et al., 2021) similarity.
 Specifically, we calculate the similarity between the image embedding of the image associated with the input prompt and the text embedding of the description/response associated with the IT instances.
 Next, for every prompt, we retrieve the top-25 IT instances and search the hallucinated phrase returned by the judge (in descriptions/responses) using string matching.
+We show how varying values of k change the final counts for our hallucination categories in Appendix K.5.
   3. 3.
@@
 These are caused by ineffective recognition of objects in the image (due to low-resolution images, weak encoder, etc.)
-H.5 Elbow Method for Logit Analysis
+K.5 Results for various values of k for IT Hallucinations
+Fig. 33 shows that even for higher and lower values of k, our hypothesis and claims stay unchanged as changing the value of k does not affect the overall frequency distribution of categories.
+[…]
+[width=0.5]chart-2.pdf
+Figure 33: Frequency comparison of hallucination categories for top-k=10,25,40
+K.6 Elbow Method for Logit Analysis
 The Knee/Elbow Detection Method operates by plotting the cumulative distribution of the sorted probabilities and seeking the point that maximizes the distance from the line connecting the first and the last points of this cumulative plot.
@@
 This method ensures that the chosen cutoff $k$ dynamically adjusts to the specific characteristics of each probability distribution encountered, aligning with the overarching goal of maintaining sampling relevance and computational efficiency.
-H.6 Examples of Rephrased Prompts
+K.7 Examples of Rephrased Prompts
 To determine whether LVLM hallucinations stem from the limitations of the base LLM or from alignment issues, we ask GPT-4 to rewrite the instructions in prompts in such a way that the base LLM can answer the question without any image input as described in section 4.2.
-We share a few examples (figures: 30, 31, 32, 33, 34, 35, 36, 37) of rephrased prompts in the existing datasets below:
+We share a few examples (figures: 34, 35, 36, 37, 38, 39, 40, 41) of rephrased prompts in the existing datasets below:
+[…]
+[width=]vallu_rephrased_img_1.pdf
 Figure 34: Rephrased prompt example for VaLLu benchmark.
+[…]
+[width=]vallu_rephrased_img_2.pdf
 Figure 35: Rephrased prompt example for VaLLu benchmark.
+[…]
+[width=]mmmu_rephrase_img_1.pdf
 Figure 36: Rephrased prompt example for MMMU benchmark.
+[…]
+[width=]mmmu_rephrase_img_2.pdf
 Figure 37: Rephrased prompt example for MMMU benchmark.
+[…]
+[width=]math_vista_rephrase_img_1.pdf
 Figure 38: Rephrased prompt example for MathVista benchmark.
+[…]
+[width=]math_vista_rephrase_img_2.pdf
 Figure 39: Rephrased prompt example for MathVista benchmark.
+[…]
+[width=]donut_rephrase_img_1.pdf
 Figure 40: Rephrased prompt 1 for SynthDoG benchmark.
+[…]
+[width=]donut_rephrase_img_2.pdf
 Figure 41: Rephrased prompt 2 for SynthDoG benchmark.
-Methodology VP VP+Info-Seek VP+Info-Seek+Reasoning VP+Reasoning
-LLaVA-v1 1.62 2.60 1.67 1.70
-LLaVA-v1.5 2.59 3.18 1.50 1.76
-LLaVA-v1.6 3.46 3.68 1.67 2.15
-mPLUG-Owl2 2.56 3.23 1.67 1.94
-InternLM-X 3.75 3.90 2.33 2.56
-CogVLM 3.00 3.37 1.25 2.08
-Table 3: Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark for the "Factuality" metric.
-Dataset Model k Variance Range Avg.
-AMBER LLaVA-v1 4.19 0.05 0.31 0.23
-LLaVA-v1.5 3.91 0.07 0.39 0.29
-LLaVA-v1.6 3.78 0.08 0.43 0.32
-mPLUG-Owl2 3.65 0.06 0.36 0.28
-InternLM-X 3.56 0.09 0.44 0.36
-CogVLM 3.16 0.91 0.47 0.38
-MMMU LLaVA-v1 3.68 0.08 0.39 0.27
-LLaVA-v1.5 3.53 0.09 0.41 0.23
-LLaVA-v1.6 3.40 0.09 0.42 0.24
-mPLUG-Owl2 3.32 0.10 0.44 0.26
-InternLM-X 3.26 0.14 0.46 0.28
-CogVLM 3.16 0.19 0.47 0.25
-Table 4: Post-truncation probability statistics for various models using the elbow method. All values are computed across all hallucinated tokens in each dataset.
-Dataset Model Helpfulness Clarity Factuality Depth Engagement
-VaLLu LLaVA-v1.5-t 3.05 4.02 2.83 2.70 3.85
-mPLUG-Owl2-t 3.70 4.15 3.14 3.11 3.24
-InternLM-X-t 4.28 4.58 3.66 3.62 3.57
-MMMU LLaVA-v1.5-t 1.97 3.25 1.69 1.87 3.25
-mPLUG-Owl2-t 2.48 3.05 1.63 2.22 2.64
-InternLM-X-t 3.51 3.93 2.43 2.98 3.01
-MathVista LLaVA-v1.5-t 2.53 3.58 2.21 2.10 3.44
-mPLUG-Owl2-t 3.56 4.05 3.04 2.52 2.79
-InternLM-X-t 4.39 4.65 3.99 2.91 2.95
-Oven LLaVA-v1.5-t 3.38 4.48 2.95 2.32 3.35
-mPLUG-Owl2-t 3.62 4.56 2.99 2.47 3.37
-InternLM-X-t 3.80 4.58 3.76 2.89 3.52
-Table 5: Performance comparison of various LVLMs on rephrased prompts without images as described in Section 5.2
-Dataset Model Helpfulness Clarity Factuality Depth Engagement
-AMBER LLaVA-v1 3.79 4.46 3.2 3.27 0.96
-LLaVA-1.5 4.29 4.73 3.91 3.44 1.26
-LLaVA-1.6 4.30 4.78 3.85 4.29 1.27
-mPLUG-Owl2 4.02 4.58 3.48 3.27 0.71
-InternLM-X 4.59 4.86 4.37 4.00 1.40
-CogVLM 4.68 4.93 4.49 3.96 1.28
-GPT-4-Turbo 4.89 4.96 4.79 4.64 1.50
-OCR LLaVA-v1 1.11 2.09 1.11 1.14 0.47
-LLaVA-1.5 1.07 2.24 1.17 1.15 0.48
-LLaVA-1.6 2.51 2.92 2.63 1.99 1.11
-mPLUG-Owl2 1.21 1.15 1.16 1.10 0.64
-InternLM-X 1.11 1.12 1.12 1.03 0.67
-CogVLM 3.12 3.04 2.68 2.53 1.47
-GPT-4-Turbo 4.15 4.02 3.78 3.98 1.67
-MMMU LLaVA-v1 1.77 2.73 1.50 1.65 0.46
-LLaVA-1.5 2.45 3.34 2.00 2.23 0.59
-LLaVA-1.6 3.54 3.98 2.95 3.39 0.99
-mPLUG-Owl2 2.94 3.64 2.56 2.33 0.82
-InternLM-X 3.60 4.10 3.24 3.15 1.05
-CogVLM 4.03 4.38 3.57 3.43 1.03
-GPT-4-Turbo 4.86 4.92 4.73 4.56 1.15
-MathVista LLaVA-v1 1.95 2.84 1.95 1.85 0.76
-LLaVA-1.5 2.57 3.47 2.23 2.36 0.98
-LLaVA-1.6 3.68 4.07 3.17 3.48 1.12
-mPLUG-Owl2 3.14 3.85 2.78 2.46 1.03
-InternLM-X 3.91 4.32 3.45 3.35 1.23
-CogVLM 4.20 4.52 3.78 3.66 1.24
-GPT-4-Turbo 4.54 4.97 4.12 4.10 1.35
-MMC LLaVA-v1 1.81 2.90 1.51 1.75 0.56
-LLaVA-1.5 2.66 3.49 2.21 2.42 0.71
-LLaVA-1.6 4.09 4.32 3.50 3.90 1.09
-mPLUG-Owl2 3.63 4.15 3.32 2.60 0.97
-InternLM-X 3.72 4.17 3.28 3.25 1.10
-CogVLM 4.27 4.58 3.91 3.92 1.32
-GPT-4-Turbo 4.81 4.89 4.57 4.49 1.12
-Table 6: Performance comparison of various LVLMs on popular benchmarks for image description. While AMBER has real-world scenes, others do not. Experiment described in Section 4.1.
-Model Methodology Helpfulness Clarity Factuality Depth Engagement
-LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25
-Vanilla-sampling 2.64 3.73 1.86 1.94 1.27
-VDD 1.70 2.55 1.47 1.81 1.59
-OPERA 2.60 3.75 2.04 1.89 1.30
-Woodpecker 2.68 3.81 2.01 1.95 1.28
-LRV 2.64 3.78 1.98 1.94 1.36
-LURE 2.60 3.54 2.03 1.89 1.40
-VDGD (ours) 2.38 3.39 2.16 2.46 1.34
-LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25
-Vanilla-sampling 2.57 3.65 2.01 1.92 1.22
-VDD 1.94 2.99 1.55 1.65 1.05
-OPERA 2.73 3.74 2.05 2.04 1.22
-Woodpecker 2.74 3.87 2.05 2.03 1.29
-LRV 2.76 3.82 2.10 2.03 1.28
-LURE 2.72 3.76 2.03 2.05 1.27
-VDGD (ours) 2.97 3.91 2.64 2.27 1.54
-LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43
-Vanilla-sampling 3.10 3.85 2.64 2.63 1.40
-VDD 2.01 2.96 1.80 2.10 1.20
-OPERA 3.10 4.01 2.62 2.60 1.42
-Woodpecker 3.16 4.00 2.67 2.58 1.44
-LRV 3.15 3.98 2.65 2.62 1.39
-LURE 3.11 4.02 2.64 2.64 1.48
-VDGD (ours) 3.51 4.04 3.16 2.87 1.58
-mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28
-Vanilla-sampling 2.72 3.60 2.18 1.91 1.24
-VDD 2.16 2.95 1.80 1.85 1.26
-OPERA 2.92 3.96 2.26 2.08 1.34
-Woodpecker 2.89 3.90 2.23 2.02 1.24
-LRV 2.90 3.94 2.19 2.05 1.28
-LURE 2.92 3.89 2.24 2.06 1.32
-VDGD (ours) 3.14 3.98 2.72 2.22 1.50
-InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47
-Vanilla-sampling 3.13 4.04 2.70 2.41 1.56
-VDD 2.56 3.26 2.32 2.19 1.41
-OPERA 3.14 4.20 2.65 2.32 1.46
-Woodpecker 3.13 4.12 2.60 2.28 1.41
-LRV 3.06 4.19 2.59 2.31 1.45
-LURE 3.12 4.10 2.64 2.40 1.49
-VDGD (ours) 3.57 4.37 3.45 2.65 1.68
-CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55
-Vanilla-sampling 3.26 4.21 2.83 2.40 1.55
-VDD 2.80 3.62 2.63 2.10 1.58
-OPERA 3.40 4.27 2.85 2.46 1.57
-Woodpecker 3.44 4.30 2.91 2.45 1.59
-LRV 3.42 4.26 2.88 2.42 1.54
-LURE 3.39 4.20 2.78 2.41 1.52
-VDGD (ours) 3.33 4.15 3.01 2.51 1.60
-GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
-Table 7: Performance comparison of various LVLMs on VaLLu benchmark. This is an extension of results shown in Table 2
-Model Methodology Helpfulness Clarity Factuality Depth Engagement
-LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00
-Vanilla-sampling 1.82 2.82 1.38 1.56 1.06
-VDD 1.70 2.67 1.22 1.40 0.95
-OPERA 1.88 2.87 1.45 1.61 1.07
-Woodpecker 1.89 2.88 1.48 1.60 1.06
-LRV 1.83 2.85 1.40 1.57 1.08
-LURE 1.91 2.90 1.52 1.63 1.09
-VDGD (ours) 2.09 2.99 1.64 1.74 1.14
-LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09
-Vanilla-sampling 1.83 2.88 1.42 1.67 1.08
-VDD 1.70 2.52 1.34 1.58 0.97
-OPERA 2.18 3.08 1.62 1.75 1.22
-Woodpecker 2.15 2.93 1.57 1.63 1.18
-LRV 2.16 2.88 1.60 1.67 1.16
-LURE 2.03 2.87 1.46 1.71 1.15
-VDGD (ours) 2.30 3.17 1.85 1.86 1.20
-LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97
-Vanilla-sampling 1.80 2.63 1.50 1.78 0.96
-VDD 1.68 2.57 1.36 1.68 0.94
-OPERA 1.86 2.70 1.57 1.84 0.99
-Woodpecker 1.90 2.74 1.63 1.89 1.04
-LRV 1.88 2.72 1.61 1.85 1.02
-LURE 1.92 2.73 1.65 1.86 1.03
-VDGD (ours) 2.11 2.89 1.81 1.92 1.05
-mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16
-Vanilla-sampling 2.25 3.12 1.59 1.66 1.14
-VDD 2.10 2.98 1.42 1.56 1.04
-OPERA 2.35 3.23 1.67 1.74 1.18
-Woodpecker 2.28 3.17 1.61 1.69 1.15
-LRV 2.34 3.24 1.66 1.72 1.14
-LURE 2.32 3.21 1.63 1.67 1.19
-VDGD (ours) 2.48 3.36 1.82 1.78 1.23
-InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63
-Vanilla-sampling 2.97 3.76 2.17 2.45 1.64
-VDD 2.76 3.54 2.03 2.26 1.57
-OPERA 3.09 3.86 2.20 2.43 1.66
-Woodpecker 3.07 3.83 2.19 2.46 1.68
-LRV 3.11 3.88 2.24 2.49 1.70
-LURE 3.10 3.89 2.22 2.48 1.71
-VDGD (ours) 3.21 4.11 2.57 2.78 1.79
-CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22
-Vanilla-sampling 2.44 3.37 1.68 1.84 1.23
-VDD 2.22 3.15 1.53 1.68 1.17
-OPERA 2.56 3.47 1.76 1.88 1.27
-Woodpecker 2.57 3.47 1.78 1.90 1.26
-LRV 2.58 3.49 1.80 1.92 1.25
-LURE 2.54 3.42 1.75 1.85 1.22
-VDGD (ours) 2.65 3.56 1.92 1.97 1.29
-GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
-Table 8: Performance comparison of various LVLMs on MMMU benchmark. This is an extension of results shown in Figure 7.
-Model Methodology Helpfulness Clarity Factuality Depth Engagement
-LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42
-Vanilla-sampling 3.72 4.56 2.24 2.27 1.43
-VDD 3.54 4.32 2.05 2.07 1.36
-OPERA 3.80 4.63 2.29 2.35 1.50
-Woodpecker 3.78 4.60 2.26 2.30 1.47
-LRV 3.76 4.64 2.30 2.38 1.49
-LURE 3.74 4.59 2.25 2.28 1.46
-VDGD (ours) 3.92 4.78 2.52 2.42 1.54
-LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32
-Vanilla-sampling 3.68 4.39 2.38 2.04 1.36
-VDD 3.38 4.09 2.35 1.87 1.10
-OPERA 3.74 4.46 2.26 1.93 1.32
-Woodpecker 3.46 4.53 2.34 1.97 1.38
-LRV 3.53 4.42 2.39 1.96 1.27
-LURE 3.33 4.41 2.37 1.91 1.29
-VDGD (ours) 3.84 4.61 2.56 2.15 1.45
-LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50
-Vanilla-sampling 3.69 4.32 2.45 2.28 1.48
-VDD 3.54 4.11 2.34 2.18 1.35
-OPERA 3.80 4.47 2.59 2.46 1.57
-Woodpecker 3.78 4.42 2.56 2.43 1.53
-LRV 3.75 4.40 2.54 2.42 1.52
-LURE 3.79 4.41 2.52 2.41 1.53
-VDGD (ours) 3.90 4.53 2.79 2.51 1.58
-mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33
-Vanilla-sampling 3.80 4.45 2.51 2.12 1.34
-VDD 3.62 4.27 2.35 1.98 1.24
-OPERA 3.85 4.52 2.57 2.20 1.40
-Woodpecker 3.84 4.56 2.60 2.24 1.42
-LRV 3.82 4.54 2.59 2.22 1.41
-LURE 3.86 4.58 2.61 2.25 1.45
-VDGD (ours) 3.99 4.65 2.76 2.36 1.52
-InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55
-Vanilla-sampling 4.01 4.68 2.83 2.37 1.54
-VDD 3.82 4.37 2.66 2.10 1.44
-OPERA 4.13 4.76 2.92 2.42 1.59
-Woodpecker 4.07 4.72 2.88 2.40 1.58
-LRV 4.10 4.73 2.89 2.40 1.56
-LURE 4.14 4.75 2.91 2.43 1.60
-VDGD (ours) 4.28 4.89 3.12 2.54 1.66
-CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44
-Vanilla-sampling 3.73 4.38 2.74 2.61 1.42
-VDD 3.58 4.27 2.64 2.50 1.37
-OPERA 3.76 4.45 2.82 2.69 1.47
-Woodpecker 3.75 4.44 2.79 2.64 1.45
-LRV 3.80 4.51 2.86 2.70 1.49
-LURE 3.82 4.53 2.88 2.73 1.52
-VDGD (ours) 3.94 4.64 2.99 2.80 1.60
-GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61
-Table 9: Performance comparison of various LVLMs on MathVista benchmark. This is an extension of results shown in Figure 7.
-Model Methodology Helpfulness Clarity Factuality Depth Engagement
-LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23
-Vanilla-sampling 3.14 4.42 2.45 2.20 1.26
-VDD 2.98 4.10 2.20 2.07 1.16
-OPERA 3.20 4.48 2.50 2.22 1.28
-Woodpecker 3.21 4.49 2.46 2.21 1.25
-LRV 3.25 4.51 2.51 2.23 1.29
-LURE 3.20 4.47 2.49 2.20 1.28
-VDGD (ours) 3.42 4.59 2.78 2.30 1.30
-LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19
-Vanilla-sampling 3.24 4.40 2.65 2.58 1.10
-VDD 2.82 4.10 2.44 2.39 1.20
-OPERA 3.22 4.40 2.49 2.58 1.18
-Woodpecker 3.26 4.42 2.46 2.64 1.22
-LRV 3.25 4.38 2.52 2.59 1.24
-LURE 3.24 4.35 2.51 2.60 1.21
-VDGD (ours) 3.47 4.56 2.84 2.51 1.80
-LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35
-Vanilla-sampling 3.40 4.38 3.10 3.27 1.30
-VDD 3.26 4.19 2.96 3.13 1.20
-OPERA 3.49 4.48 3.20 3.38 1.40
-Woodpecker 3.45 4.45 3.21 3.35 1.42
-LRV 3.42 4.40 3.18 3.34 1.39
-LURE 3.38 4.36 3.18 3.28 1.36
-VDGD (ours) 3.64 4.71 3.49 3.42 1.45
-mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25
-Vanilla-sampling 3.24 4.47 2.84 2.50 1.24
-VDD 3.07 4.21 2.67 2.40 1.17
-OPERA 3.29 4.52 2.92 2.58 1.27
-Woodpecker 3.27 4.50 2.90 2.52 1.25
-LRV 3.23 4.49 2.88 2.50 1.24
-LURE 3.26 4.51 2.91 2.51 1.28
-VDGD (ours) 3.34 4.64 3.15 2.70 1.32
-InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37
-Vanilla-sampling 3.27 4.53 3.33 2.57 1.35
-VDD 3.08 4.24 3.10 2.44 1.30
-OPERA 3.37 4.65 3.44 2.65 1.43
-Woodpecker 3.38 4.63 3.45 2.66 1.44
-LRV 3.40 4.69 3.47 2.70 1.46
-LURE 3.36 4.64 3.42 2.68 1.42
-VDGD (ours) 3.54 4.78 3.68 2.79 1.49
-CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38
-Vanilla-sampling 3.59 4.50 3.37 2.78 1.41
-VDD 3.36 4.32 3.21 2.58 1.30
-OPERA 3.68 4.60 3.42 2.84 1.45
-Woodpecker 3.70 4.58 3.45 2.85 1.42
-LRV 3.65 4.57 3.40 2.82 1.40
-LURE 3.72 4.62 3.46 2.83 1.41
-VDGD (ours) 3.89 4.82 3.59 2.89 1.49
-GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
-Table 10: Performance comparison of various LVLMs on Oven benchmark. This is an extension of results shown in Figure 7.
-H.7 Compute Infrastructure
+K.8 Compute Infrastructure
 All our analysis, inference and baseline experiments are conducted on a node of 4 NVIDIA RTX A6000 GPUs, with 128GB RAM and 10 CPU cores.
 The evaluations are conducted using gpt-4-turbo-2024-04-09 model.
-Appendix I Additional Results
-I.1 Comparison of various LVLMs on rephrased prompts
+Appendix L Additional Results
+L.1 Comparison of various LVLMs on rephrased prompts
 Table 4 provides quantitative values for rephrased prompts without image input for various LVLMs, detailed in section 4.2 of the main paper.
-I.2 Comparison of various LVLMs on image description
+[…]
+!
+Dataset Model Helpfulness Clarity Factuality Depth Engagement VaLLu LLaVA-v1.5-t 3.05 4.02 2.83 2.70 3.85 mPLUG-Owl2-t 3.70 4.15 3.14 3.11 3.24 InternLM-X-t 4.28 4.58 3.66 3.62 3.57 MMMU LLaVA-v1.5-t 1.97 3.25 1.69 1.87 3.25 mPLUG-Owl2-t 2.48 3.05 1.63 2.22 2.64 InternLM-X-t 3.51 3.93 2.43 2.98 3.01 MathVista LLaVA-v1.5-t 2.53 3.58 2.21 2.10 3.44 mPLUG-Owl2-t 3.56 4.05 3.04 2.52 2.79 InternLM-X-t 4.39 4.65 3.99 2.91 2.95 Oven LLaVA-v1.5-t 3.38 4.48 2.95 2.32 3.35 mPLUG-Owl2-t 3.62 4.56 2.99 2.47 3.37 InternLM-X-t 3.80 4.58 3.76 2.89 3.52
+Table 4: Performance comparison of various LVLMs on rephrased prompts without images as described in Section 4.2
+L.2 Comparison of various LVLMs on VaLLu benchmark with no image input.
+Table 5 provides comparison of various LVLMs on VaLLU benchmark when prompted with no image but image descriptions are provided.
+[…]
+!
+Methodology LLaVA-v1 LLaVA-v1.5 LLaVA-v1.6 mPLUG-Owl2 InternLM-X CogVLM Vanilla-greedy 1.67 1.78 2.07 1.76 2.02 2.23 VCD 1.31 1.29 1.40 1.42 2.04 2.21 OPERA 1.65 1.43 2.29 1.91 2.30 2.44 Woodpecker 1.77 1.85 2.29 2.02 2.41 2.60
+Table 5: Comparison of LVLMs performance when prompted without an image but image descriptions are provided on VaLLu benchmark. The values correspond to the "Factuality" metric.
+L.3 Comparison of various LVLMs on image description
 Table 6 provides quantitative values for faithful image description generations for various LVLMs on multiple benchmarks, detailed in section 4.2 of the main paper.
-I.3 Comparison of post-truncation probability statistics on AMBER and MMMU benchmarks
+L.4 Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark
+Table 7 provides a comparison of LVLMs on different capabilities on the proposed VaLLu benchmark.
+[…]
+!
+Dataset Model Helpfulness Clarity Factuality Depth Engagement AMBER LLaVA-v1 3.79 4.46 3.2 3.27 0.96 LLaVA-1.5 4.29 4.73 3.91 3.44 1.26 LLaVA-1.6 4.30 4.78 3.85 4.29 1.27 mPLUG-Owl2 4.02 4.58 3.48 3.27 0.71 InternLM-X 4.59 4.86 4.37 4.00 1.40 CogVLM 4.68 4.93 4.49 3.96 1.28 GPT-4-Turbo 4.89 4.96 4.79 4.64 1.50 OCR LLaVA-v1 1.11 2.09 1.11 1.14 0.47 LLaVA-1.5 1.07 2.24 1.17 1.15 0.48 LLaVA-1.6 2.51 2.92 2.63 1.99 1.11 mPLUG-Owl2 1.21 1.15 1.16 1.10 0.64 InternLM-X 1.11 1.12 1.12 1.03 0.67 CogVLM 3.12 3.04 2.68 2.53 1.47 GPT-4-Turbo 4.15 4.02 3.78 3.98 1.67 MMMU LLaVA-v1 1.77 2.73 1.50 1.65 0.46 LLaVA-1.5 2.45 3.34 2.00 2.23 0.59 LLaVA-1.6 3.54 3.98 2.95 3.39 0.99 mPLUG-Owl2 2.94 3.64 2.56 2.33 0.82 InternLM-X 3.60 4.10 3.24 3.15 1.05 CogVLM 4.03 4.38 3.57 3.43 1.03 GPT-4-Turbo 4.86 4.92 4.73 4.56 1.15 MathVista LLaVA-v1 1.95 2.84 1.95 1.85 0.76 LLaVA-1.5 2.57 3.47 2.23 2.36 0.98 LLaVA-1.6 3.68 4.07 3.17 3.48 1.12 mPLUG-Owl2 3.14 3.85 2.78 2.46 1.03 InternLM-X 3.91 4.32 3.45 3.35 1.23 CogVLM 4.20 4.52 3.78 3.66 1.24 GPT-4-Turbo 4.54 4.97 4.12 4.10 1.35 MMC LLaVA-v1 1.81 2.90 1.51 1.75 0.56 LLaVA-1.5 2.66 3.49 2.21 2.42 0.71 LLaVA-1.6 4.09 4.32 3.50 3.90 1.09 mPLUG-Owl2 3.63 4.15 3.32 2.60 0.97 InternLM-X 3.72 4.17 3.28 3.25 1.10 CogVLM 4.27 4.58 3.91 3.92 1.32 GPT-4-Turbo 4.81 4.89 4.57 4.49 1.12
+Table 6: Performance comparison of various LVLMs on popular benchmarks for image description. While AMBER has real-world scenes, others do not. Experiment described in Section 3.2.
+[…]
+0.95!
+Methodology VP VP+Info-Seek VP+Info-Seek+Reasoning VP+Reasoning LLaVA-v1 1.62 2.60 1.67 1.70 LLaVA-v1.5 2.59 3.18 1.50 1.76 LLaVA-v1.6 3.46 3.68 1.67 2.15 mPLUG-Owl2 2.56 3.23 1.67 1.94 InternLM-X 3.75 3.90 2.33 2.56 CogVLM 3.00 3.37 1.25 2.08
+Table 7: Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark for the "Factuality" metric.
+L.5 Comparison of post-truncation probability statistics on AMBER and MMMU benchmarks
 Table 8 provides post-truncation probabilities for all LVLMs on AMBER and MMMU, this is an extension of the statistic shown in section 4.3 of the main paper.
-I.4 Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark
-Table 3 provides comparison of LVLMs on different capabilities on the proposed VaLLu benchmark.
-I.5 Comparison of VaLLu with Mitigation Baselines
+[…]
+0.7!
+Dataset Model k Variance Range Avg.
+AMBER LLaVA-v1 4.19 0.05 0.31 0.23 LLaVA-v1.5 3.91 0.07 0.39 0.29 LLaVA-v1.6 3.78 0.08 0.43 0.32 mPLUG-Owl2 3.65 0.06 0.36 0.28 InternLM-X 3.56 0.09 0.44 0.36 CogVLM 3.16 0.91 0.47 0.38 MMMU LLaVA-v1 3.68 0.08 0.39 0.27 LLaVA-v1.5 3.53 0.09 0.41 0.23 LLaVA-v1.6 3.40 0.09 0.42 0.24 mPLUG-Owl2 3.32 0.10 0.44 0.26 InternLM-X 3.26 0.14 0.46 0.28 CogVLM 3.16 0.19 0.47 0.25
+Table 8: Post-truncation probability statistics for various models using the elbow method. All values are computed across all hallucinated tokens in each dataset.
+L.6 Comparison of VaLLu with Mitigation Baselines
 Table 9 provides quantitative values for comparison of VDGD and other baselines on the VaLLu benchmark, detailed in section 5.3 of the main paper.
-I.6 Comparison of MMMU with Mitigation Baselines
+L.7 Comparison of MMMU with Mitigation Baselines
 Table 10 provides quantitative values for comparison of VDGD and other baselines on the MMMU benchmark, detailed in section 3.2 of the main paper.
-I.7 Comparison of MathVista with Mitigation Baselines
+L.8 Comparison of MathVista with Mitigation Baselines
 Table 11 provides quantitative values for comparison of VDGD and other baselines on the MathVista benchmark, detailed in section 3.2 of the main paper.
-I.8 Comparison of Oven with Mitigation Baselines
+L.9 Comparison of Oven with Mitigation Baselines
 Table 12 provides quantitative values for comparison of VDGD and other baselines on the Oven benchmark, detailed in section 3.2 of the main paper.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contribution and are true to the best of our knowledge.
-  5. 2.
-    Limitations
-  6. Question: Does the paper discuss the limitations of the work performed by the authors?
-  7. Answer: [Yes]
-  8. Justification: The paper’s last section i.e 7 Conclusion, Limitations and Future Work discusses the limitation of our work.
-  9. 3.
-    Theory Assumptions and Proofs
-  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  11. Answer: [N/A]
-  12. Justification: This paper does not make any theoretical assumptions and does not describe any theoretical proofs.
-  13. 4.
-    Experimental Result Reproducibility
-  14.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  15. Answer: [Yes]
-  16. Justification: To the best of our knowledge we described our experimental methods, algorithms and evaluation prompts in the paper in detail.
-  17. 5.
-    Open access to data and code
-  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  19. Answer: [Yes]
-  20.
-Justification: We have released the code for all the experiments, algorithm and our novel contributions in an anonymous git reopsitory as described in 7.
-We have also uploaded the benchmark proposed in the paper under Supplementary Materials.
-  21. 6.
-    Experimental Setting/Details
-  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  23. Answer: [Yes]
-  24.
-Justification: We have included all the evaluation prompts and hyperparameters used for testing.
-The paper does not provide any training methodologies and as such does not specify any training details.
-  25. 7.
-    Experiment Statistical Significance
-  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  27. Answer: [Yes]
-  28. Justification: We have provided error bars for the results in the main paper. Due to the expensive nature of evaluation, we provided error bars only for the main results in the paper.
-  29. 8.
-    Experiments Compute Resources
-  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  31. Answer: [Yes]
-  32. Justification: We have details about all the compute resources we have used in section H.7.
-  33. 9.
-    Code Of Ethics
-  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  35. Answer: [Yes]
-  36. Justification: To the best of our knowledge the paper adheres to all the ethical guidelines provided in the above link.
-  37. 10.
-    Broader Impacts
-  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  39. Answer: [Yes]
-  40. Justification: We have described the broader/societal impacts of the paper in section C.
-  41. 11.
-    Safeguards
-  42.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  43. Answer: [No]
-  44.
-Justification: We do not release any new data or models.
-All data and models used are from existing resources.
-The VaLLu benchmark is an amalgamation of existing benchmarks.
-We use licenses and release the dataset under licenses issued by the original benchmark.
-  45. 12.
-    Licenses for existing assets
-  46.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  47. Answer: [Yes]
-  48. Justification: All the licenses for models, datasets and baselines used in the paper are provided in appendix sections F, E and G.
-  49. 13.
-    New Assets
-  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  51. Answer: [Yes]
-  52. Justification: We introduced a new benchmark in this paper. The benchmark data and metadata file have been uploaded in the supplementary folder in an anonymized zip file.
-  53. 14.
-    Crowdsourcing and Research with Human Subjects
-  54.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  55. Answer: [No]
-  56. Justification: We did not use crowdsourcing and research with human subjects.
-  57. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  58.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  59. Answer: [Yes]
-  60. Justification: We obtained an IRB approval to filter out noisy examples in the datasets we used as described in section H.3.
+[…]
+!
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25 Vanilla-sampling 2.64 3.73 1.86 1.94 1.27 VDD 1.70 2.55 1.47 1.81 1.59 OPERA 2.60 3.75 2.04 1.89 1.30 Woodpecker 2.68 3.81 2.01 1.95 1.28 LRV 2.64 3.78 1.98 1.94 1.36 LURE 2.60 3.54 2.03 1.89 1.40 VDGD (ours) 2.38 3.39 2.16 2.46 1.34 LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25 Vanilla-sampling 2.57 3.65 2.01 1.92 1.22 VDD 1.94 2.99 1.55 1.65 1.05 OPERA 2.73 3.74 2.05 2.04 1.22 Woodpecker 2.74 3.87 2.05 2.03 1.29 LRV 2.76 3.82 2.10 2.03 1.28 LURE 2.72 3.76 2.03 2.05 1.27 VDGD (ours) 2.97 3.91 2.64 2.27 1.54 LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43 Vanilla-sampling 3.10 3.85 2.64 2.63 1.40 VDD 2.01 2.96 1.80 2.10 1.20 OPERA 3.10 4.01 2.62 2.60 1.42 Woodpecker 3.16 4.00 2.67 2.58 1.44 LRV 3.15 3.98 2.65 2.62 1.39 LURE 3.11 4.02 2.64 2.64 1.48 VDGD (ours) 3.51 4.04 3.16 2.87 1.58 mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28 Vanilla-sampling 2.72 3.60 2.18 1.91 1.24 VDD 2.16 2.95 1.80 1.85 1.26 OPERA 2.92 3.96 2.26 2.08 1.34 Woodpecker 2.89 3.90 2.23 2.02 1.24 LRV 2.90 3.94 2.19 2.05 1.28 LURE 2.92 3.89 2.24 2.06 1.32 VDGD (ours) 3.14 3.98 2.72 2.22 1.50 InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47 Vanilla-sampling 3.13 4.04 2.70 2.41 1.56 VDD 2.56 3.26 2.32 2.19 1.41 OPERA 3.14 4.20 2.65 2.32 1.46 Woodpecker 3.13 4.12 2.60 2.28 1.41 LRV 3.06 4.19 2.59 2.31 1.45 LURE 3.12 4.10 2.64 2.40 1.49 VDGD (ours) 3.57 4.37 3.45 2.65 1.68 CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55 Vanilla-sampling 3.26 4.21 2.83 2.40 1.55 VDD 2.80 3.62 2.63 2.10 1.58 OPERA 3.40 4.27 2.85 2.46 1.57 Woodpecker 3.44 4.30 2.91 2.45 1.59 LRV 3.42 4.26 2.88 2.42 1.54 LURE 3.39 4.20 2.78 2.41 1.52 VDGD (ours) 3.33 4.15 3.01 2.51 1.60 GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
+Table 9: Performance comparison of various LVLMs on VaLLu benchmark. This is an extension of results shown in Table 9
+[…]
+!
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00 Vanilla-sampling 1.82 2.82 1.38 1.56 1.06 VDD 1.70 2.67 1.22 1.40 0.95 OPERA 1.88 2.87 1.45 1.61 1.07 Woodpecker 1.89 2.88 1.48 1.60 1.06 LRV 1.83 2.85 1.40 1.57 1.08 LURE 1.91 2.90 1.52 1.63 1.09 VDGD (ours) 2.09 2.99 1.64 1.74 1.14 LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09 Vanilla-sampling 1.83 2.88 1.42 1.67 1.08 VDD 1.70 2.52 1.34 1.58 0.97 OPERA 2.18 3.08 1.62 1.75 1.22 Woodpecker 2.15 2.93 1.57 1.63 1.18 LRV 2.16 2.88 1.60 1.67 1.16 LURE 2.03 2.87 1.46 1.71 1.15 VDGD (ours) 2.30 3.17 1.85 1.86 1.20 LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97 Vanilla-sampling 1.80 2.63 1.50 1.78 0.96 VDD 1.68 2.57 1.36 1.68 0.94 OPERA 1.86 2.70 1.57 1.84 0.99 Woodpecker 1.90 2.74 1.63 1.89 1.04 LRV 1.88 2.72 1.61 1.85 1.02 LURE 1.92 2.73 1.65 1.86 1.03 VDGD (ours) 2.11 2.89 1.81 1.92 1.05 mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16 Vanilla-sampling 2.25 3.12 1.59 1.66 1.14 VDD 2.10 2.98 1.42 1.56 1.04 OPERA 2.35 3.23 1.67 1.74 1.18 Woodpecker 2.28 3.17 1.61 1.69 1.15 LRV 2.34 3.24 1.66 1.72 1.14 LURE 2.32 3.21 1.63 1.67 1.19 VDGD (ours) 2.48 3.36 1.82 1.78 1.23 InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63 Vanilla-sampling 2.97 3.76 2.17 2.45 1.64 VDD 2.76 3.54 2.03 2.26 1.57 OPERA 3.09 3.86 2.20 2.43 1.66 Woodpecker 3.07 3.83 2.19 2.46 1.68 LRV 3.11 3.88 2.24 2.49 1.70 LURE 3.10 3.89 2.22 2.48 1.71 VDGD (ours) 3.21 4.11 2.57 2.78 1.79 CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22 Vanilla-sampling 2.44 3.37 1.68 1.84 1.23 VDD 2.22 3.15 1.53 1.68 1.17 OPERA 2.56 3.47 1.76 1.88 1.27 Woodpecker 2.57 3.47 1.78 1.90 1.26 LRV 2.58 3.49 1.80 1.92 1.25 LURE 2.54 3.42 1.75 1.85 1.22 VDGD (ours) 2.65 3.56 1.92 1.97 1.29 GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
+Table 10: Performance comparison of various LVLMs on MMMU benchmark (only questions tagged with open-ended generation). This is an extension of results shown in Figure 7.
+[…]
+!
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42 Vanilla-sampling 3.72 4.56 2.24 2.27 1.43 VDD 3.54 4.32 2.05 2.07 1.36 OPERA 3.80 4.63 2.29 2.35 1.50 Woodpecker 3.78 4.60 2.26 2.30 1.47 LRV 3.76 4.64 2.30 2.38 1.49 LURE 3.74 4.59 2.25 2.28 1.46 VDGD (ours) 3.92 4.78 2.52 2.42 1.54 LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32 Vanilla-sampling 3.68 4.39 2.38 2.04 1.36 VDD 3.38 4.09 2.35 1.87 1.10 OPERA 3.74 4.46 2.26 1.93 1.32 Woodpecker 3.46 4.53 2.34 1.97 1.38 LRV 3.53 4.42 2.39 1.96 1.27 LURE 3.33 4.41 2.37 1.91 1.29 VDGD (ours) 3.84 4.61 2.56 2.15 1.45 LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50 Vanilla-sampling 3.69 4.32 2.45 2.28 1.48 VDD 3.54 4.11 2.34 2.18 1.35 OPERA 3.80 4.47 2.59 2.46 1.57 Woodpecker 3.78 4.42 2.56 2.43 1.53 LRV 3.75 4.40 2.54 2.42 1.52 LURE 3.79 4.41 2.52 2.41 1.53 VDGD (ours) 3.90 4.53 2.79 2.51 1.58 mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33 Vanilla-sampling 3.80 4.45 2.51 2.12 1.34 VDD 3.62 4.27 2.35 1.98 1.24 OPERA 3.85 4.52 2.57 2.20 1.40 Woodpecker 3.84 4.56 2.60 2.24 1.42 LRV 3.82 4.54 2.59 2.22 1.41 LURE 3.86 4.58 2.61 2.25 1.45 VDGD (ours) 3.99 4.65 2.76 2.36 1.52 InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55 Vanilla-sampling 4.01 4.68 2.83 2.37 1.54 VDD 3.82 4.37 2.66 2.10 1.44 OPERA 4.13 4.76 2.92 2.42 1.59 Woodpecker 4.07 4.72 2.88 2.40 1.58 LRV 4.10 4.73 2.89 2.40 1.56 LURE 4.14 4.75 2.91 2.43 1.60 VDGD (ours) 4.28 4.89 3.12 2.54 1.66 CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44 Vanilla-sampling 3.73 4.38 2.74 2.61 1.42 VDD 3.58 4.27 2.64 2.50 1.37 OPERA 3.76 4.45 2.82 2.69 1.47 Woodpecker 3.75 4.44 2.79 2.64 1.45 LRV 3.80 4.51 2.86 2.70 1.49 LURE 3.82 4.53 2.88 2.73 1.52 VDGD (ours) 3.94 4.64 2.99 2.80 1.60 GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61
+Table 11: Performance comparison of various LVLMs on MathVista benchmark (only questions tagged with open-ended generation). This is an extension of results shown in Figure 7.
+[…]
+!
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23 Vanilla-sampling 3.14 4.42 2.45 2.20 1.26 VDD 2.98 4.10 2.20 2.07 1.16 OPERA 3.20 4.48 2.50 2.22 1.28 Woodpecker 3.21 4.49 2.46 2.21 1.25 LRV 3.25 4.51 2.51 2.23 1.29 LURE 3.20 4.47 2.49 2.20 1.28 VDGD (ours) 3.42 4.59 2.78 2.30 1.30 LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19 Vanilla-sampling 3.24 4.40 2.65 2.58 1.10 VDD 2.82 4.10 2.44 2.39 1.20 OPERA 3.22 4.40 2.49 2.58 1.18 Woodpecker 3.26 4.42 2.46 2.64 1.22 LRV 3.25 4.38 2.52 2.59 1.24 LURE 3.24 4.35 2.51 2.60 1.21 VDGD (ours) 3.47 4.56 2.84 2.51 1.80 LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35 Vanilla-sampling 3.40 4.38 3.10 3.27 1.30 VDD 3.26 4.19 2.96 3.13 1.20 OPERA 3.49 4.48 3.20 3.38 1.40 Woodpecker 3.45 4.45 3.21 3.35 1.42 LRV 3.42 4.40 3.18 3.34 1.39 LURE 3.38 4.36 3.18 3.28 1.36 VDGD (ours) 3.64 4.71 3.49 3.42 1.45 mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25 Vanilla-sampling 3.24 4.47 2.84 2.50 1.24 VDD 3.07 4.21 2.67 2.40 1.17 OPERA 3.29 4.52 2.92 2.58 1.27 Woodpecker 3.27 4.50 2.90 2.52 1.25 LRV 3.23 4.49 2.88 2.50 1.24 LURE 3.26 4.51 2.91 2.51 1.28 VDGD (ours) 3.34 4.64 3.15 2.70 1.32 InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37 Vanilla-sampling 3.27 4.53 3.33 2.57 1.35 VDD 3.08 4.24 3.10 2.44 1.30 OPERA 3.37 4.65 3.44 2.65 1.43 Woodpecker 3.38 4.63 3.45 2.66 1.44 LRV 3.40 4.69 3.47 2.70 1.46 LURE 3.36 4.64 3.42 2.68 1.42 VDGD (ours) 3.54 4.78 3.68 2.79 1.49 CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38 Vanilla-sampling 3.59 4.50 3.37 2.78 1.41 VDD 3.36 4.32 3.21 2.58 1.30 OPERA 3.68 4.60 3.42 2.84 1.45 Woodpecker 3.70 4.58 3.45 2.85 1.42 LRV 3.65 4.57 3.40 2.82 1.40 LURE 3.72 4.62 3.46 2.83 1.41 VDGD (ours) 3.89 4.82 3.59 2.89 1.49 GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
+Table 12: Performance comparison of various LVLMs on Oven benchmark. This is an extension of results shown in Figure 7.

### v3

@@
 Finally, we introduce VaLLu, a benchmark designed for comprehensive evaluation of the cognitive capabilities of LVLMs.
-Code is available here: https://github.com/Sreyan88/VDGD  [^1] Corresponding Author: sreyang@umd.edu.
-* denotes equal technical contribution.
 1 Introduction
@@
   1. 1.
-    We conduct extensive controlled experiments to investigate the causes of hallucinations in LVLMs, categorizing them into distinct types.
-We demonstrate that existing hallucination mitigation techniques are effective only for visual recognition (VR) prompts—specifically, those that involve object recognition and image descriptions—while proving ineffective for cognitive prompts that require reasoning.
-We further show that these techniques are biased towards reducing language hallucinations in real-world scenes but fail to address hallucinations in non-real-world contexts and other categories.
+    Through extensive experiments, we categorize hallucinations in LVLMs and show that existing mitigation techniques work well for visual recognition prompts but fail for reasoning-intensive cognitive prompts.
   2. 2.
-    Identifying the core issue, we reveal that LVLMs suffer from a critical visual perception gap.
-Although these models can recognize visual elements and possess the reasoning skills needed for factual response generation, they struggle to fully perceive the visual information in the context of the prompt, resulting in hallucinated outputs for cognitive reasoning tasks.
+    Inspired by this, we identify a critical visual perception gap in LVLMs: while they can recognize visual elements and reason, they struggle to contextualize visual information with the prompt, leading to hallucinations in cognitive reasoning tasks.
   3. 3.
     We introduce VDGD, a novel and training-free method that enhances LVLM reasoning by grounding response generation in visual descriptions.
-We evaluate VDGD on X benchmarks with prompts that require deliberate reasoning and how VDGD significantly outperforms existing techniques, improving performance by 2% - 33% by reducing hallucinations.
+We evaluate VDGD on 8 benchmarks with prompts that require deliberate reasoning and how VDGD significantly outperforms existing techniques, improving performance by 2% - 33% by reducing hallucinations.
   4. 4.
     Finally, we introduce VaLLu, a meticulously curated benchmark designed to evaluate the cognitive capabilities of LVLMs.
-Comprising 1,500 carefully selected instances, VaLLu is constructed by refining existing benchmarks and rigorously removing noisy and misleading examples identified through our thorough analysis.
-VaLLu emphasizes a balanced mix of diverse cognitive tasks, focusing exclusively on questions that require open-ended responses.
 2 Related Work
@@
 These developments have been driven by innovations in model architectures (Liu et al., 2023b; Ye et al., 2023; Wang et al., 2023b), training methods (Liu et al., 2023b), alignment techniques (Chen et al., 2024b), and data augmentation techniques (Liu et al., 2023b; Wang et al., 2024b).
-To assess the capabilities of these models, the research community has introduced a variety of benchmarks specifically designed to evaluate different aspects, including visual recognition and understanding in diverse scenarios (Liu et al., 2023a; Kim et al., 2022), reasoning (Lu et al., 2023; Yue et al., 2024), knowledge-based information retrieval (Hu et al., 2023), and other specialized tasks.
+To assess the capabilities of these models, the research community has introduced a variety of benchmarks specifically designed to evaluate different aspects, including visual recognition and understanding in diverse scenarios (Liu et al., 2023a; Kim et al., 2022), reasoning (Lu et al., 2023; Yue et al., 2024; Bai et al., 2024; Liu et al., 2024b), knowledge-based information retrieval (Hu et al., 2023), and other specialized tasks.
 Reducing Hallucinations to Improve Response Quality: Hallucination mitigation broadly aims to enhance the accuracy of model responses.
@@
 Moreover, despite the growing focus on evaluation and mitigation, the underlying causes of hallucinations remain largely unexplored.
+Recent methods like Compositional Chain-of-Thought Prompting (CCoT)(Mitra et al., 2024), Visual Table(Zhong et al., 2024), and Visual Evidence Prompting (Li et al., 2024) introduce structured intermediate representations (e.g., scene graphs or hierarchical descriptions) to improve reasoning.
+While these methods show strong performance on real-world scenes, they rely heavily on explicit object-centric representations and often struggle with abstract or non-real-world inputs (e.g., charts or mathematical graphs).
+Additionally, methods like Visual Table are not training-free and require substantial computational resources.
+In contrast, VDGD is lightweight, training-free, and generalizes effectively to diverse inputs, including abstract datasets like MMMU and MathVista.
 3 How Close Are We to Reliable LVLM Responses?
@@
 […]
-[width=]vllm.drawio5.pdf
+[width=]vllm.drawio5-compressed.pdf
 Figure 4: Types of Visual Recognition Hallucinations.
@@
 These are hallucinations caused when the model fails to accurately recognize visual elements in an input image or inaccurate reasoning or knowledge extraction, among other factors.
-(3) Style Hallucinations.
+(3) 1.3 Hallucinations.
 These are hallucinations caused by style imitation.
@@
 !
-Benchmark Baseline LLaVA-v1 LLaVA-1.5 LLaVA-1.6 mPLUG-Owl2 InternLM-X CogVLM CogVLM2 Qwen2-VL MMMU Vanilla-greedy 1.26 1.35 1.42 1.40 2.01 1.66 2.08 2.39 Vanilla-sampling 1.27 1.44 1.40 1.41 2.05 1.64 2.10 2.35 VDD 1.34 1.52 1.44 1.53 2.22 1.68 2.14 2.64 OPERA 1.30 1.43 1.57 1.64 2.25 1.62 2.21 2.44 Woodpecker 1.32 1.44 1.63 1.61 2.12 1.65 2.14 2.56 LRV 1.29 1.49 1.61 1.58 2.08 1.59 2.18 2.48 LURE 1.31 1.47 1.60 1.64 2.27 1.66 2.27 2.71 PAI 1.42 1.39 1.44 1.56 2.23 1.65 2.29 2.64 HALC 1.40 1.54 1.63 1.65 2.15 1.67 2.24 2.69 VDGD (ours) 1.49 (+18%) 1.62 (+20%) 1.75 (+23%) 1.72 (+23%) 2.39 (+19%) 1.71 (+3%) 2.47 (+19%) 2.91 (+22%) MathVista Vanilla-greedy 1.56 1.65 2.00 1.92 2.56 2.15 2.45 2.97 Vanilla-sampling 1.54 1.68 1.91 1.94 2.52 2.19 2.46 2.95 VDD 1.67 1.68 2.07 2.11 2.59 2.53 2.81 3.19 OPERA 1.64 1.83 2.04 2.04 2.62 2.30 2.53 3.22 Woodpecker 1.72 2.10 2.19 2.13 2.43 2.43 2.59 3.13 LRV 1.68 1.68 2.17 1.99 2.61 2.20 2.62 3.30 LURE 1.59 2.03 2.21 2.07 2.37 2.45 2.49 3.13 PAI 1.68 1.85 1.97 2.08 2.63 2.20 2.60 3.39 HALC 1.71 2.05 2.17 1.97 2.54 2.24 2.77 3.38 VDGD (ours) 1.84 (+18%) 2.19 (+33%) 2.44 (+22%) 2.24 (+17%) 2.88 (+13%) 2.62 (+22%) 2.93 (+20%) 3.59 (+21%) MMBench Vanilla-greedy 2.67 2.89 3.11 2.98 3.58 3.19 3.71 3.93 Vanilla-sampling 2.69 2.92 3.12 3.02 3.59 3.22 3.68 3.95 VDD 2.80 3.03 3.09 3.22 3.93 3.28 3.81 4.20 OPERA 2.81 2.94 3.17 3.10 3.95 3.22 3.93 4.05 Woodpecker 2.83 2.96 3.14 3.07 3.84 3.33 3.83 4.19 LRV 2.77 2.99 3.13 3.23 3.84 3.25 3.77 4.05 LURE 2.72 2.97 3.13 3.26 3.72 3.21 3.82 4.16 PAI 2.75 3.00 3.04 3.16 3.94 3.35 3.97 4.05 HALC 2.73 2.93 3.02 3.18 3.81 3.27 4.04 4.08 VDGD (ours) 2.93 (+10%) 3.12 (+8%) 3.26 (+5%) 3.32 (+11%) 4.08 (+14%) 3.42 (+7%) 4.15 (+12%) 4.31 (+10%) MME Vanilla-greedy 3.32 3.54 3.65 3.49 3.75 3.57 3.64 3.86 Vanilla-sampling 3.34 3.53 3.62 3.48 3.77 3.58 3.66 3.85 VDD 3.46 3.61 3.77 3.59 3.96 3.67 3.95 4.21 OPERA 3.42 3.59 3.82 3.54 3.93 3.60 3.88 4.09 Woodpecker 3.37 3.55 3.76 3.63 3.82 3.59 3.92 4.15 LRV 3.43 3.63 3.78 3.52 3.95 3.61 3.96 3.99 LURE 3.42 3.62 3.87 3.67 3.93 3.72 3.77 4.10 PAI 3.38 3.56 3.81 3.55 3.89 3.69 3.82 4.04 HALC 3.47 3.58 3.83 3.71 4.01 3.59 3.80 4.14 VDGD (ours) 3.59 (+8%) 3.70 (+5%) 3.99 (+9%) 3.82 (+9%) 4.12 (+10%) 3.79 (+6%) 4.09 (+12%) 4.34 (+12%) Oven Vanilla-greedy 2.44 2.66 3.13 2.86 3.35 3.35 3.48 3.64 Vanilla-sampling 2.45 2.65 3.10 2.84 3.33 3.37 3.45 3.61 VDD 2.20 2.44 2.96 2.67 3.10 3.21 3.39 3.59 OPERA 2.50 2.49 3.20 2.92 3.44 3.42 3.53 3.68 Woodpecker 2.46 2.46 3.21 2.90 3.45 3.45 3.52 3.70 LRV 2.51 2.52 3.18 2.88 3.47 3.40 3.57 3.74 LURE 2.49 2.51 3.18 2.91 3.42 3.46 3.55 3.72 PAI 2.54 2.60 3.26 2.95 3.48 3.48 3.60 3.78 HALC 2.52 2.57 3.24 2.94 3.50 3.45 3.61 3.80 VDGD (ours) 2.78 (+14%) 2.84 (+7%) 3.49 (+12%) 3.15 (+10%) 3.68 (+10%) 3.59 (+7%) 3.75 (+8%) 3.92 (+7%) LLaVA-Bench Vanilla-greedy 3.01 3.12 3.23 3.13 4.12 3.68 4.15 4.37 Vanilla-sampling 2.98 3.10 3.27 3.17 4.08 3.71 4.17 4.33 VDD 2.93 3.16 3.18 3.15 3.89 3.57 4.13 4.28 OPERA 3.08 3.18 3.30 3.22 4.16 3.75 4.22 4.40 Woodpecker 3.10 3.22 3.36 3.23 4.21 3.73 4.21 4.39 LRV 3.16 3.19 3.30 3.22 4.27 3.72 4.20 4.49 LURE 3.18 3.26 3.35 3.24 4.25 3.64 4.27 4.51 PAI 3.20 3.28 3.38 3.25 4.30 3.74 4.32 4.54 HALC 3.25 3.27 3.37 3.22 4.24 3.76 4.35 4.53 VDGD (ours) 3.36 (+12%) 3.47 (+11%) 3.52 (+9%) 3.38 (+8%) 4.45 (+8%) 3.96 (+8%) 4.50 (+8%) 4.78 (+9%) MMVET Vanilla-greedy 2.39 2.52 2.72 2.87 3.39 3.18 3.47 3.64 Vanilla-sampling 2.34 2.50 2.70 2.84 3.32 3.35 3.42 3.62 VDD 2.29 2.61 2.67 2.92 3.29 3.42 3.39 3.58 OPERA 2.48 2.67 2.77 3.01 3.37 3.39 3.52 3.69 Woodpecker 2.47 2.64 2.82 3.06 3.40 3.44 3.50 3.72 LRV 2.58 2.56 2.79 2.92 3.38 3.42 3.51 3.75 LURE 2.54 2.63 2.81 3.12 3.41 3.46 3.56 3.71 PAI 2.62 2.60 2.83 3.18 3.37 3.49 3.59 3.78 HALC 2.66 2.70 2.88 3.15 3.44 3.47 3.54 3.76 VDGD (ours) 2.85 (+19%) 3.01 (+19%) 3.23 (+19%) 3.45 (+20%) 3.78 (+12%) 3.57 (+12%) 3.74 (+8%) 3.99 (+10%) VaLLu Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82 3.23 3.45 Vanilla-sampling 1.86 2.01 1.64 2.18 2.70 2.83 3.21 3.40 VDD 1.47 1.55 1.80 1.80 2.32 2.63 3.19 3.42 OPERA 2.04 2.05 2.62 2.26 2.65 2.85 3.22 3.47 Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91 3.28 3.52 LRV 1.98 2.10 2.65 2.19 2.59 2.88 3.32 3.51 LURE 2.03 2.03 2.64 2.24 2.64 2.78 3.28 3.48 PAI 2.04 2.23 2.78 2.34 2.76 2.86 3.31 3.54 HALC 2.02 2.22 2.74 2.32 2.72 2.89 3.29 3.52 VDGD (ours) 2.16 (+11%) 2.64 (+30%) 3.16 (+20%) 2.72 (+24%) 3.45 (+30%) 3.01 (+7%) 3.48 (+8%) 3.67 (+6%)
+Benchmark Baseline LLaVA-v1 LLaVA-1.5 LLaVA-1.6 mPLUG-Owl2 InternLM-X CogVLM CogVLM2 Qwen2-VL MMMU Vanilla-greedy 1.26 1.35 1.42 1.40 2.01 1.66 2.08 2.39 Vanilla-sampling 1.27 1.44 1.40 1.41 2.05 1.64 2.10 2.35 VCD 1.34 1.52 1.44 1.53 2.22 1.68 2.14 2.64 OPERA 1.30 1.43 1.57 1.64 2.25 1.62 2.21 2.44 Woodpecker 1.32 1.44 1.63 1.61 2.12 1.65 2.14 2.56 LRV 1.29 1.49 1.61 1.58 2.08 1.59 2.18 2.48 LURE 1.31 1.47 1.60 1.64 2.27 1.66 2.27 2.71 PAI 1.42 1.39 1.44 1.56 2.23 1.65 2.29 2.64 HALC 1.40 1.54 1.63 1.65 2.15 1.67 2.24 2.69 VDGD (ours) 1.49 (+18%) 1.62 (+20%) 1.75 (+23%) 1.72 (+23%) 2.39 (+19%) 1.71 (+3%) 2.47 (+19%) 2.91 (+22%) MathVista Vanilla-greedy 1.56 1.65 2.00 1.92 2.56 2.15 2.45 2.97 Vanilla-sampling 1.54 1.68 1.91 1.94 2.52 2.19 2.46 2.95 VCD 1.67 1.68 2.07 2.11 2.59 2.53 2.81 3.19 OPERA 1.64 1.83 2.04 2.04 2.62 2.30 2.53 3.22 Woodpecker 1.72 2.10 2.19 2.13 2.43 2.43 2.59 3.13 LRV 1.68 1.68 2.17 1.99 2.61 2.20 2.62 3.30 LURE 1.59 2.03 2.21 2.07 2.37 2.45 2.49 3.13 PAI 1.68 1.85 1.97 2.08 2.63 2.20 2.60 3.39 HALC 1.71 2.05 2.17 1.97 2.54 2.24 2.77 3.38 VDGD (ours) 1.84 (+18%) 2.19 (+33%) 2.44 (+22%) 2.24 (+17%) 2.88 (+13%) 2.62 (+22%) 2.93 (+20%) 3.59 (+21%) MMBench Vanilla-greedy 2.67 2.89 3.11 2.98 3.58 3.19 3.71 3.93 Vanilla-sampling 2.69 2.92 3.12 3.02 3.59 3.22 3.68 3.95 VCD 2.80 3.03 3.09 3.22 3.93 3.28 3.81 4.20 OPERA 2.81 2.94 3.17 3.10 3.95 3.22 3.93 4.05 Woodpecker 2.83 2.96 3.14 3.07 3.84 3.33 3.83 4.19 LRV 2.77 2.99 3.13 3.23 3.84 3.25 3.77 4.05 LURE 2.72 2.97 3.13 3.26 3.72 3.21 3.82 4.16 PAI 2.75 3.00 3.04 3.16 3.94 3.35 3.97 4.05 HALC 2.73 2.93 3.02 3.18 3.81 3.27 4.04 4.08 VDGD (ours) 2.93 (+10%) 3.12 (+8%) 3.26 (+5%) 3.32 (+11%) 4.08 (+14%) 3.42 (+7%) 4.15 (+12%) 4.31 (+10%) MME Vanilla-greedy 3.32 3.54 3.65 3.49 3.75 3.57 3.64 3.86 Vanilla-sampling 3.34 3.53 3.62 3.48 3.77 3.58 3.66 3.85 VCD 3.46 3.61 3.77 3.59 3.96 3.67 3.95 4.21 OPERA 3.42 3.59 3.82 3.54 3.93 3.60 3.88 4.09 Woodpecker 3.37 3.55 3.76 3.63 3.82 3.59 3.92 4.15 LRV 3.43 3.63 3.78 3.52 3.95 3.61 3.96 3.99 LURE 3.42 3.62 3.87 3.67 3.93 3.72 3.77 4.10 PAI 3.38 3.56 3.81 3.55 3.89 3.69 3.82 4.04 HALC 3.47 3.58 3.83 3.71 4.01 3.59 3.80 4.14 VDGD (ours) 3.59 (+8%) 3.70 (+5%) 3.99 (+9%) 3.82 (+9%) 4.12 (+10%) 3.79 (+6%) 4.09 (+12%) 4.34 (+12%) Oven Vanilla-greedy 2.44 2.66 3.13 2.86 3.35 3.35 3.48 3.64 Vanilla-sampling 2.45 2.65 3.10 2.84 3.33 3.37 3.45 3.61 VCD 2.20 2.44 2.96 2.67 3.10 3.21 3.39 3.59 OPERA 2.50 2.49 3.20 2.92 3.44 3.42 3.53 3.68 Woodpecker 2.46 2.46 3.21 2.90 3.45 3.45 3.52 3.70 LRV 2.51 2.52 3.18 2.88 3.47 3.40 3.57 3.74 LURE 2.49 2.51 3.18 2.91 3.42 3.46 3.55 3.72 PAI 2.54 2.60 3.26 2.95 3.48 3.48 3.60 3.78 HALC 2.52 2.57 3.24 2.94 3.50 3.45 3.61 3.80 VDGD (ours) 2.78 (+14%) 2.84 (+7%) 3.49 (+12%) 3.15 (+10%) 3.68 (+10%) 3.59 (+7%) 3.75 (+8%) 3.92 (+7%) LLaVA-Bench Vanilla-greedy 3.01 3.12 3.23 3.13 4.12 3.68 4.15 4.37 Vanilla-sampling 2.98 3.10 3.27 3.17 4.08 3.71 4.17 4.33 VCD 2.93 3.16 3.18 3.15 3.89 3.57 4.13 4.28 OPERA 3.08 3.18 3.30 3.22 4.16 3.75 4.22 4.40 Woodpecker 3.10 3.22 3.36 3.23 4.21 3.73 4.21 4.39 LRV 3.16 3.19 3.30 3.22 4.27 3.72 4.20 4.49 LURE 3.18 3.26 3.35 3.24 4.25 3.64 4.27 4.51 PAI 3.20 3.28 3.38 3.25 4.30 3.74 4.32 4.54 HALC 3.25 3.27 3.37 3.22 4.24 3.76 4.35 4.53 VDGD (ours) 3.36 (+12%) 3.47 (+11%) 3.52 (+9%) 3.38 (+8%) 4.45 (+8%) 3.96 (+8%) 4.50 (+8%) 4.78 (+9%) MMVET Vanilla-greedy 2.39 2.52 2.72 2.87 3.39 3.18 3.47 3.64 Vanilla-sampling 2.34 2.50 2.70 2.84 3.32 3.35 3.42 3.62 VCD 2.29 2.61 2.67 2.92 3.29 3.42 3.39 3.58 OPERA 2.48 2.67 2.77 3.01 3.37 3.39 3.52 3.69 Woodpecker 2.47 2.64 2.82 3.06 3.40 3.44 3.50 3.72 LRV 2.58 2.56 2.79 2.92 3.38 3.42 3.51 3.75 LURE 2.54 2.63 2.81 3.12 3.41 3.46 3.56 3.71 PAI 2.62 2.60 2.83 3.18 3.37 3.49 3.59 3.78 HALC 2.66 2.70 2.88 3.15 3.44 3.47 3.54 3.76 VDGD (ours) 2.85 (+19%) 3.01 (+19%) 3.23 (+19%) 3.45 (+20%) 3.78 (+12%) 3.57 (+12%) 3.74 (+8%) 3.99 (+10%) VaLLu Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82 3.23 3.45 Vanilla-sampling 1.86 2.01 1.64 2.18 2.70 2.83 3.21 3.40 VCD 1.47 1.55 1.80 1.80 2.32 2.63 3.19 3.42 OPERA 2.04 2.05 2.62 2.26 2.65 2.85 3.22 3.47 Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91 3.28 3.52 LRV 1.98 2.10 2.65 2.19 2.59 2.88 3.32 3.51 LURE 2.03 2.03 2.64 2.24 2.64 2.78 3.28 3.48 PAI 2.04 2.23 2.78 2.34 2.76 2.86 3.31 3.54 HALC 2.02 2.22 2.74 2.32 2.72 2.89 3.29 3.52 VDGD (ours) 2.16 (+11%) 2.64 (+30%) 3.16 (+20%) 2.72 (+24%) 3.45 (+30%) 3.01 (+7%) 3.48 (+8%) 3.67 (+6%)
 Table 2: Performance comparison of VDGD with various baselines. VDGD outperforms by 2%-33%.
@@
 We employ greedy decoding for all methods as we find no difference in performance on sampling.
-We generate image descriptions for VDGD using the VCD mitigation technique.
 For LVLMs, we employ LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, CogVLM, Qwen2-VL (Wang et al., 2024b) and CogVLM2 (Hong et al., 2024).
@@
 Results on other full benchmarks are provided in Appendix L.6, together with scores for other aspects beyond factuality, where we show that VDGD does not compromise the depth and expressivity of responses.
-Ablation results in Appendix C demonstrate that removing KL divergence leads to drop in performance.
-Additionally, using only the description prefix without the image leads to performance drops, confirming that VDGD enhances reasoning through grounding rather than making the model rely completely on the description alone.
+Ablation results in Appendix C demonstrate that (i) removing KL divergence leads to drop in performance (ii) using only the description prefix without the image leads to performance drops confirming that VDGD enhances reasoning through grounding rather than making the model rely completely on the description alone, and (iii) more advanced captioning models can significantly enhance VDGD performance, while smaller captioning models provide competitive results with the added benefit of lower computational overhead.
 Qualitative examples.
-Fig 10 compares responses from vanilla greedy decoding with VDGD.
+Fig.
+10 compares responses from vanilla greedy decoding with VDGD.
 Grounding response generation results in more factual responses.
@@
 […]
-[width=]vllm_sk.pdf
+[width=]vllm_sk-compressed_2.pdf
 Figure 10: Illustration of responses generated with VDGD compared with greedy decoding. More in Sec. K.2.
@@
 Our results demonstrate that VDGD is the first method to effectively reduce hallucinations in responses that necessitate additional cognitive skills beyond just visual recognition.
-As part of future work, we would like to work on the current limitations of our work, which include (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
+As part of future work, we would like to work on (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix and (2) the requirement of prompting the LVLM twice.
 7 Reproducibility Statement
-We provide our code here: https://anonymous.4open.science/r/VDGD-1E04/.
-All codes will be open-sourced upon paper acceptance, including all checkpoints.
+We provide our code here: https://sreyan88.github.io/VDGD/.
 All experimental details, including training parameters and hyper-parameters, are provided in Section 5.2.
@@
     • K.4: We provide and in-depth explanation of all the details of Algorithm 1.
+    • K.5 Results for various values of $k$ for IT Hallucinations
     • K.6: We explain the elbow method used for logit analysis.
@@
     L: We provide additional results for all the experiments described in the paper.
+  12. 12.
+    M: We provide details about computation cost for VDGD.
+  13. 13.
+    N: We provide results for VDGD with beam search.
+  14. 14.
+    O: We provide some further disccusion about related works.
 Appendix B VDGD Failure Cases
@@
 […]
-[width=]failure_cases.pdf
+[width=]failure_cases_resized.pdf
 Figure 11: Illustration of failure cases of VDGD.
 Appendix C VDGD Ablation Results
-To demonstrate the effectiveness of VDGD, we conduct ablation studies on its key components: (i) VDGD (- VCD): Image descriptions are generated without using Visual Contrastive Decoding (VCD), (ii) VDGD (- KLD): KLD-based sampling is removed, and only the description is appended, (iii) VDGD (- Image): VDGD is applied without the original image input; only the image description is appended to the prompt.
+To demonstrate the effectiveness of VDGD, we conduct ablation studies on its key components: (i) VDGD (- VCD): Image descriptions are generated without using Visual Contrastive Decoding (VCD), (ii) VDGD (+ GPT-4 Captions): Image descriptions are generated using GPT-4-turbo-2024-04-09 instead of the LVLM itself.
+(iii) VDGD (+  Ramos et al.
+Captions): Image descriptions are generated using the small captioning model introduced by  Ramos et al.
+instead of the LVLM itself.
+(iv) VDGD (- KLD): KLD-based sampling is removed, and only the description is appended, (v) VDGD (- Image): VDGD is applied without the original image input; only the image description is appended to the prompt.
 Results are presented on the VDGD benchmark in Table 3.
@@
 This shows that the input image is still essential for the model.
+As we can clearly see, VDGD achieves a significant performance boost when using captions generated by GPT-4, highlighting the potential of higher-quality captions to enhance VDGD’s effectiveness.
+This is a case of implicit knowledge transfer from GPT-4 to VDGD through detailed descriptions.
+Additionally, VDGD delivers competitive results with a small-scale captioning model, illustrating an alternative, more computationally efficient approach to improving performance.
+These findings suggest that future advancements in both large and small captioning models could further enhance the performance and efficiency of VDGD.
 […]
@@
 We selected 500 samples or the size of the benchmark, whichever was lower.
+Scores presented in Table 4 are averaged across scores provided by the 5 students.
+Table 4 presents scores in the format of original benchmark metrics / GPT Scores / Expert Human Scores.
+Remarkably, we show that our proposed GPT score has a higher correlation to Expert Human Score.
+Nonetheless, our GPT Score also has a high correlation with the original benchmark metrics.
+[…]
+0.9!
+Dataset Baseline LLaVA-1.5 mPLUG-Owl2 InternLM-X CogVLM2 Avg Correlation MME Vanilla 1501.45 / 3.54 / 3.91 1450.19 / 3.49 / 3.82 1712.00 / 3.75 / 4.15 1512.45 / 3.66 / 4.02 0.95 / 0.96 / 0.92 Woodpecker 1503.84/3.55/3.94 1508.87/3.63/3.98 1746.93 / 3.82 / 4.21 1789.63 / 3.92 / 4.33 VDGD 1698.23 / 3.70 / 4.12 1724.27 / 3.82 /4.24 1852.48 / 4.12 / 4.56 1848.87 / 4.09 / 4.34 MMMU-Val Vanilla 30.30 / 1.35 / 2.02 32.72 / 1.40 / 2.14 43.35 / 2.01 / 2.70 44.37 / 2.08 / 2.83 0.99 / 0.96 / 0.96 Woodpecker 33.12 / 1.44 / 1.83 36.35 / 1.61 / 1.79 47.58 / 2.12 / 2.60 48.98 / 2.14 / 2.53 VDGD 36.89 / 1.62 / 2.37 38.42 / 1.72 / 2.55 53.86 / 2.39 / 2.80 56.43 / 2.47 / 3.16 LLaVA-Bench Vanilla 65.42 / 3.10 / 3.49 69.37 / 3.17 / 3.60 72.21 / 4.08 / 4.22 73.76 / 4.17 / 4.19 0.96 / 0.97 / 0.95 Woodpecker 70.48 / 3.22 / 3.52 70.68 / 3.23 / 3.74 74.87 / 4.21 / 4.48 74.52 / 4.21 / 4.17 VDGD 71.03 / 3.47 / 4.04 70.57 / 3.38 / 3.91 77.43 / 4.45 / 4.72 78.03 / 4.50 / 4.73 MMVET Vanilla 31.15 / 2.50 /2.41 36.3 / 2.84 / 2.99 51.27 / 3.32 / 3.70 60.42 / 3.42 / 3.55 Woodpecker 33.42 / 2.64 / 2.88 47.26 / 3.06 / 3.29 51.67 / 3.40 / 3.81 65.78 / 3.50 / 3.75 0.97 / 0.98 / 0.91 VDGD 45.79 / 3.01 / 3.27 62.41 / 3.45 / 3.90 74.57 / 3.78 / 4.16 71.98 / 3.74 / 4.10
+Table 4: Scores across 4 benchmarks in Original Metric / (Our proposed) GPT Score / (Our) Expert Human Score format.
+We also show the person correlation averaged across datasets.
+The correlation scores are in Original Metric-GPT Score / GPT Score-Expert Human Score / Original Metric-Expert Human Score format.
+GPT Scores show a higher correlation with Expert Human Scores.
 Appendix E VDGD Computational Complexity
@@
 From our in-depth analysis of the logit space, we see that language hallucinations in LVLMs have a unique characteristic – the probability of the most confident token in the vocabulary marginally decreased from the probability of the same token in the base LLM when prompted with the same prior context (the most confident token by the LVLM will also be the most confident token by the language-only base LLM as the token on a language hallucination).
-Thus, decoding-based mitigation methods like OPERA and VDD benefit from this unique property, and contrasting tokens help mitigate hallucinations.
+Thus, decoding-based mitigation methods like OPERA and VCD benefit from this unique property, and contrasting tokens help mitigate hallucinations.
 K.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
@@
 […]
-[width=]hallu_noisy_img_1.pdf
+[width=]hallu_noisy_img_1-compressed.pdf
 Figure 31: Noisy example 1 for HallusionBench benchmark.
@@
 !
-Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25 Vanilla-sampling 2.64 3.73 1.86 1.94 1.27 VDD 1.70 2.55 1.47 1.81 1.59 OPERA 2.60 3.75 2.04 1.89 1.30 Woodpecker 2.68 3.81 2.01 1.95 1.28 LRV 2.64 3.78 1.98 1.94 1.36 LURE 2.60 3.54 2.03 1.89 1.40 VDGD (ours) 2.38 3.39 2.16 2.46 1.34 LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25 Vanilla-sampling 2.57 3.65 2.01 1.92 1.22 VDD 1.94 2.99 1.55 1.65 1.05 OPERA 2.73 3.74 2.05 2.04 1.22 Woodpecker 2.74 3.87 2.05 2.03 1.29 LRV 2.76 3.82 2.10 2.03 1.28 LURE 2.72 3.76 2.03 2.05 1.27 VDGD (ours) 2.97 3.91 2.64 2.27 1.54 LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43 Vanilla-sampling 3.10 3.85 2.64 2.63 1.40 VDD 2.01 2.96 1.80 2.10 1.20 OPERA 3.10 4.01 2.62 2.60 1.42 Woodpecker 3.16 4.00 2.67 2.58 1.44 LRV 3.15 3.98 2.65 2.62 1.39 LURE 3.11 4.02 2.64 2.64 1.48 VDGD (ours) 3.51 4.04 3.16 2.87 1.58 mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28 Vanilla-sampling 2.72 3.60 2.18 1.91 1.24 VDD 2.16 2.95 1.80 1.85 1.26 OPERA 2.92 3.96 2.26 2.08 1.34 Woodpecker 2.89 3.90 2.23 2.02 1.24 LRV 2.90 3.94 2.19 2.05 1.28 LURE 2.92 3.89 2.24 2.06 1.32 VDGD (ours) 3.14 3.98 2.72 2.22 1.50 InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47 Vanilla-sampling 3.13 4.04 2.70 2.41 1.56 VDD 2.56 3.26 2.32 2.19 1.41 OPERA 3.14 4.20 2.65 2.32 1.46 Woodpecker 3.13 4.12 2.60 2.28 1.41 LRV 3.06 4.19 2.59 2.31 1.45 LURE 3.12 4.10 2.64 2.40 1.49 VDGD (ours) 3.57 4.37 3.45 2.65 1.68 CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55 Vanilla-sampling 3.26 4.21 2.83 2.40 1.55 VDD 2.80 3.62 2.63 2.10 1.58 OPERA 3.40 4.27 2.85 2.46 1.57 Woodpecker 3.44 4.30 2.91 2.45 1.59 LRV 3.42 4.26 2.88 2.42 1.54 LURE 3.39 4.20 2.78 2.41 1.52 VDGD (ours) 3.33 4.15 3.01 2.51 1.60 GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 2.61 3.69 1.95 1.92 1.25 Vanilla-sampling 2.64 3.73 1.86 1.94 1.27 VCD 1.70 2.55 1.47 1.81 1.59 OPERA 2.60 3.75 2.04 1.89 1.30 Woodpecker 2.68 3.81 2.01 1.95 1.28 LRV 2.64 3.78 1.98 1.94 1.36 LURE 2.60 3.54 2.03 1.89 1.40 VDGD (ours) 2.38 3.39 2.16 2.46 1.34 LLaVA-1.5 Vanilla-greedy 2.64 3.72 2.03 2.03 1.25 Vanilla-sampling 2.57 3.65 2.01 1.92 1.22 VCD 1.94 2.99 1.55 1.65 1.05 OPERA 2.73 3.74 2.05 2.04 1.22 Woodpecker 2.74 3.87 2.05 2.03 1.29 LRV 2.76 3.82 2.10 2.03 1.28 LURE 2.72 3.76 2.03 2.05 1.27 VDGD (ours) 2.97 3.91 2.64 2.27 1.54 LLaVA-1.6 Vanilla-greedy 3.08 3.81 2.63 2.67 1.43 Vanilla-sampling 3.10 3.85 2.64 2.63 1.40 VCD 2.01 2.96 1.80 2.10 1.20 OPERA 3.10 4.01 2.62 2.60 1.42 Woodpecker 3.16 4.00 2.67 2.58 1.44 LRV 3.15 3.98 2.65 2.62 1.39 LURE 3.11 4.02 2.64 2.64 1.48 VDGD (ours) 3.51 4.04 3.16 2.87 1.58 mPLUG-Owl2 Vanilla-greedy 2.73 3.64 2.20 1.96 1.28 Vanilla-sampling 2.72 3.60 2.18 1.91 1.24 VCD 2.16 2.95 1.80 1.85 1.26 OPERA 2.92 3.96 2.26 2.08 1.34 Woodpecker 2.89 3.90 2.23 2.02 1.24 LRV 2.90 3.94 2.19 2.05 1.28 LURE 2.92 3.89 2.24 2.06 1.32 VDGD (ours) 3.14 3.98 2.72 2.22 1.50 InternLM-X Vanilla-greedy 3.11 3.98 2.66 2.34 1.47 Vanilla-sampling 3.13 4.04 2.70 2.41 1.56 VCD 2.56 3.26 2.32 2.19 1.41 OPERA 3.14 4.20 2.65 2.32 1.46 Woodpecker 3.13 4.12 2.60 2.28 1.41 LRV 3.06 4.19 2.59 2.31 1.45 LURE 3.12 4.10 2.64 2.40 1.49 VDGD (ours) 3.57 4.37 3.45 2.65 1.68 CogVLM Vanilla-greedy 3.25 4.19 2.82 2.43 1.55 Vanilla-sampling 3.26 4.21 2.83 2.40 1.55 VCD 2.80 3.62 2.63 2.10 1.58 OPERA 3.40 4.27 2.85 2.46 1.57 Woodpecker 3.44 4.30 2.91 2.45 1.59 LRV 3.42 4.26 2.88 2.42 1.54 LURE 3.39 4.20 2.78 2.41 1.52 VDGD (ours) 3.33 4.15 3.01 2.51 1.60 GPT-4-Turbo - 4.02 4.44 3.63 2.94 1.59
 Table 10: Performance comparison of various LVLMs on VaLLu benchmark. This is an extension of results shown in Table 2
@@
 !
-Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00 Vanilla-sampling 1.82 2.82 1.38 1.56 1.06 VDD 1.70 2.67 1.22 1.40 0.95 OPERA 1.88 2.87 1.45 1.61 1.07 Woodpecker 1.89 2.88 1.48 1.60 1.06 LRV 1.83 2.85 1.40 1.57 1.08 LURE 1.91 2.90 1.52 1.63 1.09 VDGD (ours) 2.09 2.99 1.64 1.74 1.14 LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09 Vanilla-sampling 1.83 2.88 1.42 1.67 1.08 VDD 1.70 2.52 1.34 1.58 0.97 OPERA 2.18 3.08 1.62 1.75 1.22 Woodpecker 2.15 2.93 1.57 1.63 1.18 LRV 2.16 2.88 1.60 1.67 1.16 LURE 2.03 2.87 1.46 1.71 1.15 VDGD (ours) 2.30 3.17 1.85 1.86 1.20 LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97 Vanilla-sampling 1.80 2.63 1.50 1.78 0.96 VDD 1.68 2.57 1.36 1.68 0.94 OPERA 1.86 2.70 1.57 1.84 0.99 Woodpecker 1.90 2.74 1.63 1.89 1.04 LRV 1.88 2.72 1.61 1.85 1.02 LURE 1.92 2.73 1.65 1.86 1.03 VDGD (ours) 2.11 2.89 1.81 1.92 1.05 mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16 Vanilla-sampling 2.25 3.12 1.59 1.66 1.14 VDD 2.10 2.98 1.42 1.56 1.04 OPERA 2.35 3.23 1.67 1.74 1.18 Woodpecker 2.28 3.17 1.61 1.69 1.15 LRV 2.34 3.24 1.66 1.72 1.14 LURE 2.32 3.21 1.63 1.67 1.19 VDGD (ours) 2.48 3.36 1.82 1.78 1.23 InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63 Vanilla-sampling 2.97 3.76 2.17 2.45 1.64 VDD 2.76 3.54 2.03 2.26 1.57 OPERA 3.09 3.86 2.20 2.43 1.66 Woodpecker 3.07 3.83 2.19 2.46 1.68 LRV 3.11 3.88 2.24 2.49 1.70 LURE 3.10 3.89 2.22 2.48 1.71 VDGD (ours) 3.21 4.11 2.57 2.78 1.79 CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22 Vanilla-sampling 2.44 3.37 1.68 1.84 1.23 VDD 2.22 3.15 1.53 1.68 1.17 OPERA 2.56 3.47 1.76 1.88 1.27 Woodpecker 2.57 3.47 1.78 1.90 1.26 LRV 2.58 3.49 1.80 1.92 1.25 LURE 2.54 3.42 1.75 1.85 1.22 VDGD (ours) 2.65 3.56 1.92 1.97 1.29 GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 1.81 2.79 1.36 1.51 1.00 Vanilla-sampling 1.82 2.82 1.38 1.56 1.06 VCD 1.70 2.67 1.22 1.40 0.95 OPERA 1.88 2.87 1.45 1.61 1.07 Woodpecker 1.89 2.88 1.48 1.60 1.06 LRV 1.83 2.85 1.40 1.57 1.08 LURE 1.91 2.90 1.52 1.63 1.09 VDGD (ours) 2.09 2.99 1.64 1.74 1.14 LLaVA-1.5 Vanilla-greedy 1.86 2.90 1.47 1.70 1.09 Vanilla-sampling 1.83 2.88 1.42 1.67 1.08 VCD 1.70 2.52 1.34 1.58 0.97 OPERA 2.18 3.08 1.62 1.75 1.22 Woodpecker 2.15 2.93 1.57 1.63 1.18 LRV 2.16 2.88 1.60 1.67 1.16 LURE 2.03 2.87 1.46 1.71 1.15 VDGD (ours) 2.30 3.17 1.85 1.86 1.20 LLaVA-1.6 Vanilla-greedy 1.85 2.67 1.54 1.80 0.97 Vanilla-sampling 1.80 2.63 1.50 1.78 0.96 VCD 1.68 2.57 1.36 1.68 0.94 OPERA 1.86 2.70 1.57 1.84 0.99 Woodpecker 1.90 2.74 1.63 1.89 1.04 LRV 1.88 2.72 1.61 1.85 1.02 LURE 1.92 2.73 1.65 1.86 1.03 VDGD (ours) 2.11 2.89 1.81 1.92 1.05 mPLUG-Owl2 Vanilla-greedy 2.24 3.12 1.58 1.65 1.16 Vanilla-sampling 2.25 3.12 1.59 1.66 1.14 VCD 2.10 2.98 1.42 1.56 1.04 OPERA 2.35 3.23 1.67 1.74 1.18 Woodpecker 2.28 3.17 1.61 1.69 1.15 LRV 2.34 3.24 1.66 1.72 1.14 LURE 2.32 3.21 1.63 1.67 1.19 VDGD (ours) 2.48 3.36 1.82 1.78 1.23 InternLM-X Vanilla-greedy 2.95 3.71 2.14 2.39 1.63 Vanilla-sampling 2.97 3.76 2.17 2.45 1.64 VCD 2.76 3.54 2.03 2.26 1.57 OPERA 3.09 3.86 2.20 2.43 1.66 Woodpecker 3.07 3.83 2.19 2.46 1.68 LRV 3.11 3.88 2.24 2.49 1.70 LURE 3.10 3.89 2.22 2.48 1.71 VDGD (ours) 3.21 4.11 2.57 2.78 1.79 CogVLM Vanilla-greedy 2.42 3.34 1.69 1.83 1.22 Vanilla-sampling 2.44 3.37 1.68 1.84 1.23 VCD 2.22 3.15 1.53 1.68 1.17 OPERA 2.56 3.47 1.76 1.88 1.27 Woodpecker 2.57 3.47 1.78 1.90 1.26 LRV 2.58 3.49 1.80 1.92 1.25 LURE 2.54 3.42 1.75 1.85 1.22 VDGD (ours) 2.65 3.56 1.92 1.97 1.29 GPT-4-Turbo - 4.15 4.53 3.66 2.99 1.95
 Table 11: Performance comparison of various LVLMs on MMMU benchmark (only questions tagged with open-ended generation). This is an extension of results shown in Figure 7.
@@
 !
-Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42 Vanilla-sampling 3.72 4.56 2.24 2.27 1.43 VDD 3.54 4.32 2.05 2.07 1.36 OPERA 3.80 4.63 2.29 2.35 1.50 Woodpecker 3.78 4.60 2.26 2.30 1.47 LRV 3.76 4.64 2.30 2.38 1.49 LURE 3.74 4.59 2.25 2.28 1.46 VDGD (ours) 3.92 4.78 2.52 2.42 1.54 LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32 Vanilla-sampling 3.68 4.39 2.38 2.04 1.36 VDD 3.38 4.09 2.35 1.87 1.10 OPERA 3.74 4.46 2.26 1.93 1.32 Woodpecker 3.46 4.53 2.34 1.97 1.38 LRV 3.53 4.42 2.39 1.96 1.27 LURE 3.33 4.41 2.37 1.91 1.29 VDGD (ours) 3.84 4.61 2.56 2.15 1.45 LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50 Vanilla-sampling 3.69 4.32 2.45 2.28 1.48 VDD 3.54 4.11 2.34 2.18 1.35 OPERA 3.80 4.47 2.59 2.46 1.57 Woodpecker 3.78 4.42 2.56 2.43 1.53 LRV 3.75 4.40 2.54 2.42 1.52 LURE 3.79 4.41 2.52 2.41 1.53 VDGD (ours) 3.90 4.53 2.79 2.51 1.58 mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33 Vanilla-sampling 3.80 4.45 2.51 2.12 1.34 VDD 3.62 4.27 2.35 1.98 1.24 OPERA 3.85 4.52 2.57 2.20 1.40 Woodpecker 3.84 4.56 2.60 2.24 1.42 LRV 3.82 4.54 2.59 2.22 1.41 LURE 3.86 4.58 2.61 2.25 1.45 VDGD (ours) 3.99 4.65 2.76 2.36 1.52 InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55 Vanilla-sampling 4.01 4.68 2.83 2.37 1.54 VDD 3.82 4.37 2.66 2.10 1.44 OPERA 4.13 4.76 2.92 2.42 1.59 Woodpecker 4.07 4.72 2.88 2.40 1.58 LRV 4.10 4.73 2.89 2.40 1.56 LURE 4.14 4.75 2.91 2.43 1.60 VDGD (ours) 4.28 4.89 3.12 2.54 1.66 CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44 Vanilla-sampling 3.73 4.38 2.74 2.61 1.42 VDD 3.58 4.27 2.64 2.50 1.37 OPERA 3.76 4.45 2.82 2.69 1.47 Woodpecker 3.75 4.44 2.79 2.64 1.45 LRV 3.80 4.51 2.86 2.70 1.49 LURE 3.82 4.53 2.88 2.73 1.52 VDGD (ours) 3.94 4.64 2.99 2.80 1.60 GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.74 4.54 2.23 2.28 1.42 Vanilla-sampling 3.72 4.56 2.24 2.27 1.43 VCD 3.54 4.32 2.05 2.07 1.36 OPERA 3.80 4.63 2.29 2.35 1.50 Woodpecker 3.78 4.60 2.26 2.30 1.47 LRV 3.76 4.64 2.30 2.38 1.49 LURE 3.74 4.59 2.25 2.28 1.46 VDGD (ours) 3.92 4.78 2.52 2.42 1.54 LLaVA-1.5 Vanilla-greedy 3.65 4.38 2.36 2.01 1.32 Vanilla-sampling 3.68 4.39 2.38 2.04 1.36 VCD 3.38 4.09 2.35 1.87 1.10 OPERA 3.74 4.46 2.26 1.93 1.32 Woodpecker 3.46 4.53 2.34 1.97 1.38 LRV 3.53 4.42 2.39 1.96 1.27 LURE 3.33 4.41 2.37 1.91 1.29 VDGD (ours) 3.84 4.61 2.56 2.15 1.45 LLaVA-1.6 Vanilla-greedy 3.72 4.38 2.50 2.34 1.50 Vanilla-sampling 3.69 4.32 2.45 2.28 1.48 VCD 3.54 4.11 2.34 2.18 1.35 OPERA 3.80 4.47 2.59 2.46 1.57 Woodpecker 3.78 4.42 2.56 2.43 1.53 LRV 3.75 4.40 2.54 2.42 1.52 LURE 3.79 4.41 2.52 2.41 1.53 VDGD (ours) 3.90 4.53 2.79 2.51 1.58 mPLUG-Owl2 Vanilla-greedy 3.79 4.43 2.48 2.08 1.33 Vanilla-sampling 3.80 4.45 2.51 2.12 1.34 VCD 3.62 4.27 2.35 1.98 1.24 OPERA 3.85 4.52 2.57 2.20 1.40 Woodpecker 3.84 4.56 2.60 2.24 1.42 LRV 3.82 4.54 2.59 2.22 1.41 LURE 3.86 4.58 2.61 2.25 1.45 VDGD (ours) 3.99 4.65 2.76 2.36 1.52 InternLM-X Vanilla-greedy 4.02 4.69 2.84 2.35 1.55 Vanilla-sampling 4.01 4.68 2.83 2.37 1.54 VCD 3.82 4.37 2.66 2.10 1.44 OPERA 4.13 4.76 2.92 2.42 1.59 Woodpecker 4.07 4.72 2.88 2.40 1.58 LRV 4.10 4.73 2.89 2.40 1.56 LURE 4.14 4.75 2.91 2.43 1.60 VDGD (ours) 4.28 4.89 3.12 2.54 1.66 CogVLM Vanilla-greedy 3.72 4.39 2.76 2.63 1.44 Vanilla-sampling 3.73 4.38 2.74 2.61 1.42 VCD 3.58 4.27 2.64 2.50 1.37 OPERA 3.76 4.45 2.82 2.69 1.47 Woodpecker 3.75 4.44 2.79 2.64 1.45 LRV 3.80 4.51 2.86 2.70 1.49 LURE 3.82 4.53 2.88 2.73 1.52 VDGD (ours) 3.94 4.64 2.99 2.80 1.60 GPT-4-Turbo - 4.05 4.75 3.17 2.24 1.61
 Table 12: Performance comparison of various LVLMs on MathVista benchmark (only questions tagged with open-ended generation). This is an extension of results shown in Figure 7.
@@
 !
-Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23 Vanilla-sampling 3.14 4.42 2.45 2.20 1.26 VDD 2.98 4.10 2.20 2.07 1.16 OPERA 3.20 4.48 2.50 2.22 1.28 Woodpecker 3.21 4.49 2.46 2.21 1.25 LRV 3.25 4.51 2.51 2.23 1.29 LURE 3.20 4.47 2.49 2.20 1.28 VDGD (ours) 3.42 4.59 2.78 2.30 1.30 LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19 Vanilla-sampling 3.24 4.40 2.65 2.58 1.10 VDD 2.82 4.10 2.44 2.39 1.20 OPERA 3.22 4.40 2.49 2.58 1.18 Woodpecker 3.26 4.42 2.46 2.64 1.22 LRV 3.25 4.38 2.52 2.59 1.24 LURE 3.24 4.35 2.51 2.60 1.21 VDGD (ours) 3.47 4.56 2.84 2.51 1.80 LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35 Vanilla-sampling 3.40 4.38 3.10 3.27 1.30 VDD 3.26 4.19 2.96 3.13 1.20 OPERA 3.49 4.48 3.20 3.38 1.40 Woodpecker 3.45 4.45 3.21 3.35 1.42 LRV 3.42 4.40 3.18 3.34 1.39 LURE 3.38 4.36 3.18 3.28 1.36 VDGD (ours) 3.64 4.71 3.49 3.42 1.45 mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25 Vanilla-sampling 3.24 4.47 2.84 2.50 1.24 VDD 3.07 4.21 2.67 2.40 1.17 OPERA 3.29 4.52 2.92 2.58 1.27 Woodpecker 3.27 4.50 2.90 2.52 1.25 LRV 3.23 4.49 2.88 2.50 1.24 LURE 3.26 4.51 2.91 2.51 1.28 VDGD (ours) 3.34 4.64 3.15 2.70 1.32 InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37 Vanilla-sampling 3.27 4.53 3.33 2.57 1.35 VDD 3.08 4.24 3.10 2.44 1.30 OPERA 3.37 4.65 3.44 2.65 1.43 Woodpecker 3.38 4.63 3.45 2.66 1.44 LRV 3.40 4.69 3.47 2.70 1.46 LURE 3.36 4.64 3.42 2.68 1.42 VDGD (ours) 3.54 4.78 3.68 2.79 1.49 CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38 Vanilla-sampling 3.59 4.50 3.37 2.78 1.41 VDD 3.36 4.32 3.21 2.58 1.30 OPERA 3.68 4.60 3.42 2.84 1.45 Woodpecker 3.70 4.58 3.45 2.85 1.42 LRV 3.65 4.57 3.40 2.82 1.40 LURE 3.72 4.62 3.46 2.83 1.41 VDGD (ours) 3.89 4.82 3.59 2.89 1.49 GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
+Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23 Vanilla-sampling 3.14 4.42 2.45 2.20 1.26 VCD 2.98 4.10 2.20 2.07 1.16 OPERA 3.20 4.48 2.50 2.22 1.28 Woodpecker 3.21 4.49 2.46 2.21 1.25 LRV 3.25 4.51 2.51 2.23 1.29 LURE 3.20 4.47 2.49 2.20 1.28 VDGD (ours) 3.42 4.59 2.78 2.30 1.30 LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19 Vanilla-sampling 3.24 4.40 2.65 2.58 1.10 VCD 2.82 4.10 2.44 2.39 1.20 OPERA 3.22 4.40 2.49 2.58 1.18 Woodpecker 3.26 4.42 2.46 2.64 1.22 LRV 3.25 4.38 2.52 2.59 1.24 LURE 3.24 4.35 2.51 2.60 1.21 VDGD (ours) 3.47 4.56 2.84 2.51 1.80 LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35 Vanilla-sampling 3.40 4.38 3.10 3.27 1.30 VCD 3.26 4.19 2.96 3.13 1.20 OPERA 3.49 4.48 3.20 3.38 1.40 Woodpecker 3.45 4.45 3.21 3.35 1.42 LRV 3.42 4.40 3.18 3.34 1.39 LURE 3.38 4.36 3.18 3.28 1.36 VDGD (ours) 3.64 4.71 3.49 3.42 1.45 mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25 Vanilla-sampling 3.24 4.47 2.84 2.50 1.24 VCD 3.07 4.21 2.67 2.40 1.17 OPERA 3.29 4.52 2.92 2.58 1.27 Woodpecker 3.27 4.50 2.90 2.52 1.25 LRV 3.23 4.49 2.88 2.50 1.24 LURE 3.26 4.51 2.91 2.51 1.28 VDGD (ours) 3.34 4.64 3.15 2.70 1.32 InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37 Vanilla-sampling 3.27 4.53 3.33 2.57 1.35 VCD 3.08 4.24 3.10 2.44 1.30 OPERA 3.37 4.65 3.44 2.65 1.43 Woodpecker 3.38 4.63 3.45 2.66 1.44 LRV 3.40 4.69 3.47 2.70 1.46 LURE 3.36 4.64 3.42 2.68 1.42 VDGD (ours) 3.54 4.78 3.68 2.79 1.49 CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38 Vanilla-sampling 3.59 4.50 3.37 2.78 1.41 VCD 3.36 4.32 3.21 2.58 1.30 OPERA 3.68 4.60 3.42 2.84 1.45 Woodpecker 3.70 4.58 3.45 2.85 1.42 LRV 3.65 4.57 3.40 2.82 1.40 LURE 3.72 4.62 3.46 2.83 1.41 VDGD (ours) 3.89 4.82 3.59 2.89 1.49 GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
 Table 13: Performance comparison of various LVLMs on Oven benchmark. This is an extension of results shown in Figure 7.
+Appendix M Computational Efficiency
+Table 14 presents a computational analysis of various methods evaluated on LLaVA 1.5 using a 48GB GPU, comparing their average inference time and computational cost measured in teraFLOPs (T).
+While the VDGD method requires more computation than the baseline methods, the increase is modest and within a reasonable range when considering the significant performance gains it offers.
+Specifically, VDGD has an average inference time of 2.5 seconds and a computational cost of 11.9 teraFLOPs, which is a manageable increase compared to the Vanilla-greedy method’s 1.3 seconds and 9.3 teraFLOPs.
+Similarly, the combination of VDGD with a Small Captioning Model Ramos et al.
+(2023) demonstrates acceptable computational demands with an inference time of 2.1 seconds and 11.4 T.
+These results suggest that the enhanced performance of VDGD methods is achieved with computational requirements that are well within practical limits, making it a viable option.
+Method Average Inference Time (s) FLOPs (T)
+Vanilla-greedy 1.3 9.3
+VCD 1.9 10.2
+VDGD + Small Captioning Model 2.1 11.4
+VDGD 2.4 11.9
+Table 14: Comparison of average inference time and computational cost (in teraFLOPs) among VDGD and baseline methods evaluated on LLaVA 1.5 using a 48GB GPU.
+Appendix N VDGD with Beam Search
+VDGD can be seamlessly applied to beam search decoding without modifying its core methodology.
+Before a set of tokens is selected for each beam, VDGD reweights the logit space of the current logit based on the KL-Divergence between the current logit and the image description logits.
+This process is repeated at every decoding step.
+Another important point to note is that VDGD operates independently of prior response tokens, relying solely on the fixed image description tokens, which remain constant even during beam search.
+As a result, VDGD can be directly integrated into beam search decoding to achieve similar improvements in reducing hallucinations.
+Table 15 shows the result of VDGD with beam search on MMMU dataset.
+Benchmark LLaVA-v1 LLaVA-1.5
+Vanilla-greedy 1.26 1.35
+Vanilla-sampling 1.27 1.44
+VDD 1.34 1.52
+OPERA 1.30 1.43
+Woodpecker 1.32 1.44
+LRV 1.29 1.49
+LURE 1.31 1.47
+PAI 1.42 1.39
+HALC 1.40 1.54
+VDGD 1.42 1.62
+VDGD + Beam Search 1.39 1.58
+Table 15: Performance comparison of VDGD w/ beam search for LLaVA-v1 and LLaVA-1.5
+Appendix O Further Discussion on related works
+Recent advances in understanding hallucinations in Large Vision-Language Models (LVLMs) provide critical insights into addressing the persistent challenge of hallucination in multimodal systems.
+Bai et al.
+(2024) highlight the growing importance of mitigating hallucinations in LVLMs, particularly in tasks involving complex interactions between visual and textual data.
+Their comprehensive survey dissects hallucinations into granular categories, such as object attributes and relations, and identifies factors contributing to hallucinations across data, model architecture, and inference stages.
+Liu et al.
+(2023c) similarly focus on categorizing and analyzing hallucinations, emphasizing the importance of detailed evaluation metrics and benchmarks tailored to measure both generative and discriminative capabilities of LVLMs.
+They propose frameworks for systematically assessing hallucination symptoms and mitigating them through refined model alignment and enhanced multimodal representation.
+These studies underscore the necessity of bridging gaps in current LVLM evaluation methods and mitigation strategies by addressing unique challenges such as data quality, alignment module efficiency, and decoding mechanisms.
+Our proposed VDGD model aligns with these efforts by introducing a robust framework to reduce hallucinations, leveraging state-of-the-art insights while contributing novel mitigation pathways.
+This connection to the broader context enriches our understanding and opens pathways for integrating these findings into future iterations of VDGD and related LVLM frameworks.